### v1

# DESIGNER: Design-Logic-Guided Multidisciplinary Data Synthesis for LLM Reasoning

Abstract
Large language models (LLMs) have achieved remarkable success in many natural language tasks but still struggle with complex, multi-step reasoning, particularly across diverse disciplines. Existing reasoning datasets often either lack disciplinary breadth or the structural depth necessary to elicit robust reasoning behaviors. We propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages naturally available, extensive raw documents (book corpus and web corpus) to generate multidisciplinary challenging questions. A core innovation of our approach is the introduction of a Design Logic concept, which mimics the question-creation process of human educators. We use LLMs to reverse-engineer and abstract over 120,000 design logics from existing questions across various disciplines. By matching these design logics with disciplinary source materials, we are able to create reasoning questions that far surpass the difficulty and diversity of existing datasets. Based on this pipeline, we synthesized two large-scale reasoning datasets that span 75 disciplines: Design-Logic-Reasoning-Book (DLR-Book), containing 3.04 million challenging questions synthesized from the book corpus, and Design-Logic-Reasoning-Web (DLR-Web), with 1.66 million challenging questions from the web corpus. Our data analysis demonstrates that the questions synthesized by our method exhibit substantially greater difficulty and diversity than those in the baseline datasets. We validate the effectiveness of these datasets by conducting SFT experiments on the Qwen3-8B-Base and Qwen3-4B-Base models. The results show that our dataset significantly outperforms existing multidisciplinary datasets of the same volume. Training with the full datasets further enables the models to surpass the multidisciplinary reasoning performance of the official Qwen3-8B and Qwen3-4B models. [^1] Project page: https://attention-is-all-i-need.github.io/Design-Logic-Reasoning
footnotetext: * First two authors contributed equally.    Corresponding Authors: Yongchi Zhao, Jiaheng Liu.
1 Introduction
Large language models (LLMs) have demonstrated exceptional capabilities in various natural reasoning tasks (Brown et al., 2020; Kaplan et al., 2020; Hoffmann et al., 2022; Chowdhery et al., 2023; OpenAI, 2023), such as mathematics and coding, especially when utilizing long chain-of-thought (CoT) techniques (Guo et al., 2025; Moshkov et al., 2025; Cai et al., 2025). However, their performance in a wide range of university-level discipline-specific reasoning questions still falls short of human experts. A major bottleneck is the lack of a large-scale, high-quality, and diverse multidisciplinary reasoning data. Existing datasets primarily focus on math and programming, training models on questions collected from math and coding competition platforms (Moshkov et al., 2025; Cai et al., 2025). This approach heavily relies on the massive open question resources in these specific domains, but many other academic disciplines lack such abundant resources, which limits the development of LLMs’ multidisciplinary reasoning capabilities.
Data synthesis is an effective way to address data scarcity by generating a large number of reasoning questions from existing LLMs (Wang et al., 2023; Xu et al., 2023; Jung et al., 2025; Havrilla et al., 2025). Current question synthesis methods can be broadly categorized into query-centric and doc-centric approaches. Query-centric methods, centered on an initial “seed” question pool, iteratively evolve existing questions. These methods generate more complex and diverse new questions by rewriting existing ones, adding constraints (e.g., Evol-Instruct), or incorporating a chain-of-thought (Wang et al., 2023; Xu et al., 2023; Yu et al., 2025). However, the breadth and depth of the generated questions are severely limited by the quality and coverage of the initial seed pool. They are also susceptible to the biases of the generative model itself, making it difficult to transcend domain limitations and create high-quality questions across a wide range of disciplines. In contrast, doc-centric methods start with a vast collection of unstructured (e.g., web pages, books) or structured (e.g., knowledge graphs) documents (Yue et al., 2024; Yuan et al., 2025; Huang et al., 2025). By directly extracting or reasoning out question-answer pairs from these documents, this approach ensures that the synthesized data is closely tied to specific domain knowledge and facts, theoretically allowing it to cover a broader range of disciplines. The main challenge for these methods is to effectively control the difficulty and diversity of the generated questions and prevent them from degrading into simple surface-level factual recall tasks.
To address these issues, as shown in Figure 1, we propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages readily available, large-scale, multidisciplinary raw documents (e.g., book corpus and web corpus) to synthesize a massive number of multidisciplinary challenging questions. The core innovation of our method lies in the introduction of a novel concept: “Design Logic”. We observed that when human education experts design challenging and insightful questions, they don’t merely state facts. Instead, they follow a structured design process, such as: identifying key knowledge points $\rightarrow$ constructing a scenario $\rightarrow$ designing a reasoning path $\rightarrow$ pre-setting distractors. We contend that this process embodies a “Design Logic”: a sequence of deliberate steps that transforms fundamental knowledge points into complex, context-rich questions that demand multi-stage reasoning. This design logic itself is a valuable and transferable form of “meta-knowledge”. Therefore, our pipeline is not limited to directly extracting questions from documents. Instead, we first use a powerful LLM to reverse-engineer and abstract tens of thousands of “meta design logics” from existing high-quality question banks across various disciplines. During the synthesis phase, we precisely match the most suitable design logic to a source document and instruct the LLM to strictly follow the intrinsic reasoning steps of that logic to construct a brand-new question.
Specifically, our approach follows a refined pipeline. First, we perform comprehensive processing on the massive book corpus and web corpus through multi-dimensional labeling and filtering (e.g., by discipline, readability, and knowledge value) to build a high-quality source material library that covers a wide range of disciplines. Next, from a private question bank containing hundreds of millions of questions, we use clustering and sampling to select a subset of questions that are both difficult and diverse. We then instruct a powerful LLM to summarize and abstract over 120,000 structured “design logics” from this subset. In the question synthesis phase, we pioneered a two-stage “retrieve-and-generate” matching mechanism. First, we use vector similarity to retrieve a rough, coarse-grained set of the most relevant candidate logics for each source document from our design logic library. Then, an LLM performs a fine-grained evaluation to select the optimal logic and strictly applies this design logic’s steps to the source document content, generating a new and challenging reasoning question.
The main contributions of this paper can be summarized as follows:
  • We propose a data engineering pipeline for synthesizing challenging questions from raw text corpora. Based on this pipeline, we constructed two large-scale reasoning datasets: Design-Logic-Reasoning-Book (DLR-Book), containing 3.04 million challenging questions synthesized from the book corpus, and Design-Logic-Reasoning-Web (DLR-Web), with 1.66 million challenging questions from the web corpus. These datasets expand beyond common disciplines like math to cover 75 disciplines (including STEM, humanities, social sciences, arts, as well as applied and professional fields) to better train LLMs’ multidisciplinary reasoning capabilities.
  • We introduce a novel, scalable data synthesis paradigm based on “Design Logic”, which imitates the wisdom of human experts in question creation and significantly enhances the reasoning depth and diversity of doc-centric synthesized questions. Our data analysis demonstrates that the questions synthesized by our method exhibit substantially greater difficulty and diversity than those in the baseline datasets.
  • Our comprehensive comparative and ablation experiments on the Qwen3-8B-Base and Qwen3-4B-Base models (Yang et al., 2025a) show that training with our data not only significantly improves the model’s multidisciplinary reasoning capabilities but also outperforms existing multidisciplinary datasets with the same data volume. Training with our full dataset allows the resulting models to surpass the performance of the official Qwen3 models of the same size.
Figure 1: The Design-Logic-Guided Multidisciplinary Data Synthesis Pipeline.
2 Overview
We propose a data engineering pipeline for synthesizing challenging questions from raw text corpora (web corpus and book corpus). The complete pipeline is illustrated in Figure 1.
The first stage involves data curation, a meticulous process tailored to each of our three primary data sources and unified under a comprehensive 75-discipline taxonomy:
  • Web Corpus: We apply reasoning-oriented filtering and relabeling to a large web dataset, retaining only texts that are most useful for studying reasoning processes.
  • Book Corpus: This corpus undergoes preprocessing, discipline labeling, and a quality-prioritized sampling method to ensure the selection of high-readability and high-helpfulness text segments.
  • Question Bank: We annotate our question bank with multi-dimensional labels (discipline, difficulty, type) and use a clustering-based approach to select a diverse and high-quality subset.
The second stage focuses on question synthesis. We first extract underlying design logics from the sampled question subset. We then leverage a two-stage “retrieve-and-generate” mechanism to apply these reusable design logics to our curated text corpora. This process allows us to strictly follow the steps of each logic to construct millions of new reasoning questions, ensuring they are both challenging and diverse while mirroring the complexity of human-designed problems. Finally, we synthesize corresponding long CoT answers for each question, creating a valuable dataset for supervised fine-tuning.
Based on this pipeline, we constructed two large-scale reasoning datasets: Design-Logic-Reasoning-Book (DLR-Book), which contains 3.04 million challenging questions synthesized from the book corpus, and Design-Logic-Reasoning-Web (DLR-Web), with 1.66 million questions from the web corpus. These datasets expand beyond common disciplines like mathematics to cover 75 disciplines (including STEM and the humanities and social sciences), thereby better supporting the training of LLMs for multidisciplinary reasoning. Additionally, we have included concise reference answers in the datasets, generated from the text corpora concurrently with the questions, as well as long CoT responses generated by Qwen3-235B-A22B-Thinking-2507-FP8.
3 Data Curation
3.1 Data Collection
We curate three primary data sources and define a comprehensive discipline taxonomy:
  • Web Corpus: For the web corpus, we employ FineFineWeb [^2] https://huggingface.co/datasets/m-a-p/FineFineWeb, a filtered subset of the Common Crawl dataset.
  • Book Corpus: We utilize a proprietary library of books.
  • Question Bank: A proprietary repository of examination and practice items.
Discipline Taxonomy: We have established a classification system comprising 75 distinct disciplines, as shown in LABEL:tab:discipline_stats_en. This taxonomy provides comprehensive coverage across several major areas:
  • STEM: Science, Technology, Engineering, and Mathematics.
  • Humanities and Social Sciences: Including fields such as law, philosophy, and sociology.
  • Applied and Professional Fields: Encompassing domains like clinical medicine, education, and business administration.
  • Arts.
3.2 Data Processing and Filtering
3.2.1 Question Bank Processing
Multi-dimensional Labeling
Using Qwen3-30B-A3B (non-thinking), we annotate more than 150 million questions from the proprietary question bank with discipline, difficulty, and type labels. The specific prompts for these classifications are detailed in Figure 3, Figure 4, and Figure 5, respectively.
Clustering-Based Question Selection
To facilitate the extraction of design logic, we intend to filter a high-quality and diverse subset from the question bank. This subset is specifically curated to ensure a broad coverage of multiple disciplines and varying levels of difficulty. We compute embeddings with Qwen3-Embedding-4B and apply K-means clustering within each discipline (Ahmed et al., 2020). We then equally sample from each cluster with a difficulty ratio of Very Hard:Hard:Medium = 3:2:1, ensuring both diversity and sufficient difficulty. For each discipline, we determine the number of clusters via silhouette coefficient search and set per-discipline sample sizes according to the question bank’s discipline distribution. This process resulted in a curated set of 132,409 questions for design-logic extraction.
3.2.2 Book Corpus Processing
Text Preprocessing
We process each book at the chapter level. Chapters that exceed 5,000 words are further split into blocks up to 5,000 words. We then perform MinHash deduplication to obtain the final book segments.
Discipline Labeling and Quality Assessment
We assign discipline labels to book segments using a ModernBERT-large–based model that has been fine-tuned for this specific classification task (Warner et al., 2024). Subsequently, we introduce two key metrics for quality assessment: readability and helpfulness. Readability is predicted by a BERT-based classifier (Turc et al., 2019), which is designed to filter out incoherent or disorganized text. Helpfulness is scored by the fineweb-edu-classifier (Lozhkov et al., 2024), which provides a score on a scale of 0 to 5, allowing us to quantitatively assess the educational value of each document.
Quality-Prioritized Sampling
A total of over three million book segments were sampled from the corpus. For each discipline, we allocate quotas proportional to its frequency in both the book corpus and the question bank. To ensure the acquisition of high-quality data, we devised a quality-prioritized sampling policy. We first removed all samples for which the readability was classified as negative. Based on the helpfulness scores, the remaining candidate segments within each discipline are sorted in descending order, and samples are then selected from the top of this list until the predetermined quota is met. This process ensures that the sampled texts are of the highest quality available, with the vast majority of the final selected segments exhibiting a helpfulness score of $\geq 2$.
3.2.3 Web Corpus Processing
Reasoning-Oriented Filtering
Because not every FineFineWeb text contains sufficient reasoning content, we design a five-level scoring rubric (scores starting at 0; prompt in the Figure 6) to assess a text’s potential usefulness for studying reasoning processes. We utilized the Qwen3-30B-A3B (non-thinking) model to score 6.5 billion texts from the FineFineWeb dataset, retaining only those with a score of $\geq 3$.
Discipline Relabeling
Finally, we used the Qwen3-30B-A3B (non-thinking) model and the prompt specified in Figure 3 to relabel the retained texts, aligning the web corpus with our 75-discipline taxonomy.
4 Synthesizing Questions using Design Logic
4.1 Design Logic Extraction
Human educators typically construct exam questions by following a structured chain of design decisions that transforms simple knowledge concepts into complex and challenging questions rather than merely enumerating facts. A representative design process might involve: identifying the target objectives and knowledge points to be tested $\rightarrow$ constructing a contextual scenario $\rightarrow$ designing the logical reasoning process $\rightarrow$ formulating the correct answer $\rightarrow$ incorporating question traps and distracting information $\rightarrow$ validating the final question. Consequently, a solver must deconstruct the question through multi-step reasoning, which goes far beyond simple memorization (e.g., explaining a knowledge point).
Inspired by this practice, we propose a method for question synthesis based on “design logic”. Leveraging the prompt presented in Figure 7, we instruct an LLM (DeepSeek-R1-0528) to analyze the sampled authentic questions from our question bank. For each question, the LLM is tasked with three key steps: (i) deduce the question designer’s thought process, (ii) analyze how the question was constructed from the relevant knowledge points, and (iii) abstract and summarize the underlying design logic and principles. We specifically require the model to organize the resulting abstract design logic into the Mermaid format. This procedure yields a reusable pool of design logics, which the model can leverage these diverse logics to generate new questions based on provided source materials.
This approach eliminates the need for handcrafting a large set of disparate prompts, significantly increasing automation. It also substantially improves the difficulty and diversity of LLM-generated questions while aligning their structure more closely with questions authored by human teachers.
4.2 Design Logic Deduplication
To enhance the diversity of design principles provided to the LLMs, we further deduplicate the extracted design logic. This ensures a varied set of principles, which in turn facilitates the generation of diverse questions. Specifically, we use the Qwen3-Embedding-4B model to map each design logic into an embedding vector. For each discipline, we compute pairwise semantic similarities between all design logics to obtain a similarity matrix $S\in\mathbb{R}^{n\times n}$, where $S_{ij}$ denotes the semantic similarity between the $i$-th and $j$-th design logic.
Algorithm 1 Graph-based Deduplication via Centroid Selection
1:Input: A set of items $\mathcal{D}=\{d_{1},\dots,d_{n}\}$, a similarity matrix $S\in\mathbb{R}^{n\times n}$, a similarity threshold $\tau$.
2:Output: A deduplicated set of representative items $\mathcal{R}$.
3:
4:Initialize an undirected graph $G=(V,E)$ where $V=\{1,\dots,n\}$ and $E=\emptyset$.
5:Initialize the set of representatives $\mathcal{R}=\emptyset$.
6:
7:// Build a similarity graph where nodes are items and edges connect similar items.
8:for $i=1$ to $n$ do
9:  for $j=i+1$ to $n$ do
10:   if $S_{ij}>\tau$ then
11:     Add edge $(i,j)$ to $E$.
12:   end if
13:  end for
14:end for
15:
16:// Identify clusters of duplicates by finding connected components.
17:Let $\mathcal{C}\leftarrow\text{FindConnectedComponents}(G)$.
18:
19:// Select the most representative item (centroid) from each cluster.
20:for each connected component $C\in\mathcal{C}$ do
21:  Find centroid index $i^{*}=\underset{i\in C}{\arg\max}\sum_{j\in C,j\neq i}S_{ij}$.
22:  Add item $d_{i^{*}}$ to $\mathcal{R}$.
23:end for
24:
25:return $\mathcal{R}$.
We then apply graph-based connected-components deduplication. We model the $n$ design logics within a discipline as a graph $G=(V,E)$ with $|V|=n$. An undirected edge is added between nodes $i$ and $j$ if their semantic similarity $S_{ij}\geq\tau$. This construction forms connected components, where each component represents a group of mutually redundant items. From each connected component $C$, we select a unique representative — the item with the maximum sum of similarities to all other items in the component, and discard the rest. The complete procedure is detailed in Algorithm 1, with the similarity threshold set to $\tau=0.85$.
After deduplication, we retain a total of 125,328 unique design logics. The per-discipline counts are reported in Table LABEL:tab:discipline_stats_en.
4.3 Question Synthesis
To address the computational complexity of directly matching each design logic with every text segment, which would lead to a combinatorial explosion, we propose an approach similar to Retrieval-Augmented Generation (RAG). We compute the cosine similarity between the embeddings of each text segment and each design logic for each discipline-specific corpus (book or web) using the Qwen3-Embedding-4B model with a task-specific instruction (Figure 8). Specifically, for a text segment $t$ and a design logic $d$, we calculate $s(t,d)\,=\,\cos\!\bigl{(}\mathbf{e}(t),\,\mathbf{e}(d)\bigr{)}$. We then retain the top-5 design logics with the highest similarity scores as candidates for question synthesis for each text segment.
We then prompt an LLM (DeepSeek-R1-0528) to generate a challenging, graduate-level exam question (prompt detailed in Figure 9). Specifically, the LLM is instructed to (i) select the most suitable question-design logic from the top-5 candidates for the given source text, and (ii) strictly follow the corresponding logic and steps to construct a challenging question. This two-stage process acts as a coarse-to-fine ranking mechanism. The initial retrieval of the top-5 design logics using semantic similarity serves as a coarse-grained recall stage. The LLM’s subsequent selection of the most appropriate logic then functions as a fine-grained ranking stage. This strategy ensures a precise match between the design logic and the text, significantly enhancing the quality of the synthesized questions.
[…]
Question Deduplication and Decontamination
We implemented a two-stage filtering pipeline to manage redundancy and avoid evaluation leakage. First, we used MinHash-based detection to remove highly similar questions. Second, to address potential data contamination, we decontaminated our synthetic questions by comparing them against all evaluation benchmarks employed in this study. This was achieved through the standard 13-gram decontamination method, during which non-semantic punctuation was ignored. Any question flagged as similar is discarded.
Using the filtered book and web corpora from Section 3.2, we generate one corresponding reasoning question per text segment with DeepSeek-R1-0528. After deduplication and decontamination, our final reasoning dataset comprises 3,040,620 challenging questions synthesized from the book corpus and 1,658,541 challenging questions synthesized from the web corpus.
4.4 Response Synthesis
To demonstrate that our synthesized questions can effectively elicit and transfer the long CoT capabilities of a reasoning model and improve the performance of models trained on this data, we employ Qwen3-235B-A22B-Thinking-2507-FP8 to generate a corresponding long CoT response for each synthesized question. These question-response pairs are then used for supervised fine-tuning (SFT).
5 Data Analysis
To begin our quality assessment, we performed a quantitative analysis comparing the complexity and difficulty of our synthesized datasets (DLR-Book and DLR-Web) with the baseline datasets. This analysis focused on the distribution of question types, the lengths of questions and responses, the overall difficulty of the questions, and their diversity.
5.1 Question Type Analysis
As shown in Table 1, problem-solving questions form the majority of our synthesized datasets, constituting 64.92% of the book corpus and 63.91% of the web corpus. Multiple-choice questions are the next most common type, comprising 29.94% and 32.43%, respectively. This distribution analysis indicates that the composition of our synthesized questions is skewed towards reasoning-based types, as opposed to simple recall-oriented ones.
Table 1: Distribution of question types in our synthesized datasets.
Type DLR-Book DLR-Web
Problem-solving question 64.92% 63.91%
Multiple-choice question 29.94% 32.43%
Proof question 4.39% 2.67%
Other question types 0.75% 0.99%
5.2 Data Complexity
In Table 2, we present a comparison of the average lengths of questions and their corresponding responses for our two synthesized datasets against WebInstruct (Full) and NaturalReasoning. The responses were all generated by the Qwen3-235B-A22B-Thinking-2507-FP8 model. It is evident that the questions generated by our method, as well as the responses required to answer them, are significantly longer than those in WebInstruct (Full) (Yue et al., 2024) and NaturalReasoning (Yuan et al., 2025). This observation indirectly substantiates the high degree of difficulty and complexity inherent in our synthesized questions. This substantial increase in question length indicates that our synthesis method generates questions with more detailed context and more intricate premises, requiring a deeper understanding of the input. We likewise observe longer response lengths in our datasets, a direct consequence of the increased question complexity. This observation suggests that solving our synthesized questions requires a deeper chain of reasoning compared to the baselines.
Table 2: Average length (in characters) of questions and corresponding responses. The responses were all generated by the Qwen3-235B-A22B-Thinking-2507-FP8 model. The average lengths for WebInstruct (Full) and NaturalReasoning were calculated from a random sample of 304,181 instances.
Dataset Avg. Question Length Avg. Response Length
WebInstruct (Full) 180.43 11133.88
NaturalReasoning 332.08 17162.85
DLR-Book 1284.79 18090.06
DLR-Web 1205.96 17528.59
5.3 Difficulty Analysis
To provide a more direct measure of difficulty, we used the Qwen3-30B-A3B-Instruct-2507 model to assign difficulty labels to questions from our book and web corpora, as well as from WebInstruct (Full) and NaturalReasoning. Based on the complexity and length of the required reasoning, we categorized questions into four difficulty levels: Easy, Medium, Hard, and Very Hard.
As shown in Table 3, our datasets are significantly more difficult. The combined percentage of “Hard” and “Very Hard” questions is notably high: 89.84% for DLR-Book and 83.15% for DLR-Web. This contrasts sharply with the baseline datasets. DLR-Book contains 54.66% “Very Hard” questions, substantially more than the 31.11% in NaturalReasoning and 3.59% in WebInstruct (Full). Conversely, the proportion of “Easy” questions in our datasets is negligible (0.27% and 0.72%), especially when compared to WebInstruct (Full)’s 39.02%.
Collectively, these quantitative analyses of question length and difficulty distribution consistently demonstrate that our data synthesis method produces questions of significantly greater complexity and difficulty than the baseline datasets.
Table 3: Difficulty distribution of questions across datasets.
Dataset Easy Medium Hard Very Hard
WebInstruct (Full) 39.02% 39.58% 17.84% 3.59%
NaturalReasoning 2.10% 26.25% 40.54% 31.11%
DLR-Book 0.27% 9.88% 35.18% 54.66%
DLR-Web 0.72% 16.13% 36.24% 46.91%
5.4 Diversity Analysis
To evaluate the semantic diversity of the synthesized questions, we conducted a quantitative analysis using distance-based metrics. This approach assesses diversity by examining the distribution of question embeddings in a high-dimensional semantic space.
Diversity Metrics and Experimental Setup
We first generated high-dimensional vector representations for each question using the Qwen3-Embedding-4B model. For a given question set $X=\{x_{1},x_{2},\dots,x_{N}\}$, this process yields a corresponding set of embedding vectors $E=\{e_{1},e_{2},\dots,e_{N}\}$, where $e_{i}\in\mathbb{R}^{d}$ is the $d$-dimensional embedding for question $x_{i}$. We uniformly sample N = 300,000 questions from each dataset and compute the following five distance-based metrics in the embedding space.
  1. 1.
    Mean Cosine Distance (DistMean Cosine) (Yang et al., 2025b): The average cosine distance between all unique pairs of embeddings, calculated as $M_{\text{cosine}}=\frac{2}{N(N-1)}\sum_{i<j}(1-\cos(e_{i},e_{j}))$. A higher value indicates greater semantic dissimilarity.
  2. 2.
    Mean L2 Distance (DistMean L2) (Yang et al., 2025b): The average Euclidean distance between all unique pairs of embeddings, calculated as $M_{\text{L2}}=\frac{2}{N(N-1)}\sum_{i<j}\|e_{i}-e_{j}\|_{2}$. This measures the average separation of questions in the embedding space.
  3. 3.
    1-Nearest Neighbor Distance (1-NN Distance) (Stasaski & Hearst, 2022): The average cosine distance from each embedding to its single nearest neighbor, given by $M_{\text{1-NN}}=\frac{1}{N}\sum_{i}d_{1}(e_{i})$, where $d_{1}(e_{i})$ is the cosine distance from $e_{i}$ to its closest neighbor. This metric highlights the presence of tightly clustered, near-identical questions.
  4. 4.
    Cluster Inertia (Du & Black, 2019): The total squared distance of samples to their closest cluster center after applying the K-means clustering algorithm. It is calculated as $M_{\text{inertia}}=\sum_{i=1}^{N}\min_{j}\|e_{i}-c_{j}\|_{2}^{2}$, where $c_{j}$ are the cluster centroids. This measures the overall spread and density of the data clusters.
  5. 5.
    Radius (Lai et al., 2020): The geometric mean of the standard deviations of the embedding dimensions, modeling the data as a multi-dimensional Gaussian distribution: $M_{\text{Radius}}=(\prod_{j=1}^{d}\sigma_{j})^{1/d}$, where $\sigma_{j}$ is the standard deviation along the $j$-th dimension. It directly quantifies the spread of the data in the semantic space.
As detailed in Table 4, our datasets, DLR-Book and DLR-Web, consistently demonstrate a greater diversity of questions compared to the baseline datasets across all five semantic diversity metrics. The higher Mean Cosine Distance and Mean L2 Distance values confirm that our synthesized questions are, on average, more semantically distinct than those in WebInstruct (Full) and NaturalReasoning, indicating a broader conceptual scope. The most notable difference is observed in the 1-NN Distance, where our datasets score approximately twice as high as the baselines. This suggests our method generates far fewer semantically redundant questions. Furthermore, the Cluster Inertia and Radius scores for our datasets indicate that the generated questions occupy a larger and more varied volume within the semantic embedding space. This quantitative evidence confirms that our synthesis pipeline produces not only more complex and difficult questions but also a significantly more diverse set of questions.
Table 4: Semantic diversity metrics for our synthesized datasets and the baseline datasets, calculated on a uniform sample of 300,000 questions from each. Higher is better for all metrics shown.
Dataset DistMean Cosine DistMean L2 1-NN Distance Cluster Inertia Radius
WebInstruct (Full) 0.7762 1.2436 0.1830 205,590.02 0.0169
NaturalReasoning 0.8233 1.2818 0.1915 226,288.91 0.0173
DLR-Book 0.8471 1.3008 0.3726 238,100.2656 0.0176
DLR-Web 0.8494 1.3026 0.3897 238,039.5000 0.0177
6 Experiments
In this study, we evaluate the effectiveness of our synthesized data by conducting supervised fine-tuning (SFT) on base models. The resulting models are then assessed on a diverse range of widely-used benchmarks. Our extensive experiments demonstrate that the data synthesized by our method can effectively elicit the long CoT reasoning capabilities of LLMs for complex, multidisciplinary questions. Furthermore, with the same amount of training data, our method yields superior model performance compared to existing datasets like NaturalReasoning and WebInstruct (Full).
6.1 Experimental Setup
SFT Settings
In the majority of our experiments, we perform SFT on the Qwen3-8B-Base model. Subsequently, we also validate the effectiveness of our synthesized data across other model scales. For all SFT experiments, we adhere to the hyperparameters detailed in Table 5.
Table 5: Hyperparameters for Supervised Fine-Tuning (SFT).
Parameter Value
Epoch 6
Batch Size 64
Learning Rate 1e-5
Learning Rate Schedule Cosine decay to 0
Evaluation Settings
To ensure a fair comparison, we employ a zero-shot evaluation setting for all trained models, using the consistent generation configuration specified in Table 6. For each benchmark, we conduct $N$ independent sampling rollouts for each test instance. Our evaluation involves the following three metrics:
  • Pass@1: The mean accuracy (%) across the $N$ rollouts. We report both the Pass@1 score and its standard deviation to show the performance and stability of our method.
  • CoT-SC: The accuracy (%) determined by a majority vote over the $N$ generated samples. This metric is equivalent to Self-Consistency with Chain-of-Thought (CoT-SC) method and is reported for all cases where $N>1$.
  • Pass@N: The proportion (%) of questions for which at least one correct solution is found among $N$ rollouts.
Table 6: Generation configuration for evaluation.
Parameter Value
Temperature 0.6
Top-K 20
Top-P 0.95
Max Context Length 32768
Benchmarks. The benchmarks used in our evaluation, along with their respective disciplines and the number of rollouts ($N$), are detailed in Table 7.
Table 7: Evaluation benchmarks, their disciplines, and the number of rollouts (NN).
Benchmark Disciplines Rollout ($N$)
AIME 2024 Mathematics 64
AIME 2025 Mathematics 64
MATH-500 (Lightman et al., 2024) Mathematics 4
GPQA-Diamond (Rein et al., 2024) Physics, Chemistry, Biology 10
GPQA-Main (Rein et al., 2024) Physics, Chemistry, Biology 10
SuperGPQA (Du et al., 2025) 285 graduate-level disciplines across 13 fields (e.g., Engineering, Management, Economics, Education, History, Science, Medicine, Law, Sociology, Philosophy, Agriculture, Literature) 1
MMLU (Hendrycks et al., 2020) 57 disciplines including STEM, humanities, social sciences, and professional fields 1
MMLU-Pro (Wang et al., 2024) Math, Physics, Chemistry, Law, Engineering, Economics, Health, Psychology, Business, Biology, Philosophy, Computer Science, History, Other 1
6.2 Supervised Fine-Tuning Experiments
We performed supervised fine-tuning (SFT) on the Qwen3-4B-Base [^3] https://huggingface.co/Qwen/Qwen3-4B-Base and Qwen3-8B-Base [^4] https://huggingface.co/Qwen/Qwen3-8B-Base models using our synthetic datasets, DLR-Book and DLR-Web. The resulting SFT models were then compared against the official Qwen3-4B [^5] https://huggingface.co/Qwen/Qwen3-4B and Qwen3-8B [^6] https://huggingface.co/Qwen/Qwen3-8B models (thinking mode) under our evaluation settings.
Table 8: Performance of models fine-tuned on our datasets. The models are trained on DLR-Web, DLR-Book, or the combined DLR-Web+Book dataset. Baselines include the official Qwen3-4B and Qwen3-8B models (thinking mode), evaluated under our settings. The best and second-best scores in each column are shown in bold and underlined, respectively.
Model MMLU MMLU-Pro GPQA-Diamond GPQA-Main SuperGPQA
Pass@1 Pass@1 Pass@1 CoT-SC Pass@1 CoT-SC Pass@1
[…]
As shown in Table 8, supervised fine-tuning with our synthetic datasets significantly improves model performance. For both the 4B and 8B model scales, our fine-tuned models consistently outperform the official Qwen3-4B and Qwen3-8B (thinking mode) baselines across all reasoning benchmarks. Notably, models trained on the combined “DLR-Web+Book” dataset achieve the best results across all benchmarks, while those trained on the “DLR-Book” dataset achieve the second-best. This improvement is particularly pronounced on highly complex reasoning tasks like GPQA-Diamond. These results affirm the efficacy of our data synthesis strategy for enhancing the reasoning capabilities of language models.
6.3 Performance on Advanced Mathematical Reasoning
For challenging mathematical competition questions, such as those from AIME, we additionally report Pass@N results. Prior work has shown that methods like Reinforcement Learning with Verifiable Rewards (RLVR) primarily enhance Pass@1 accuracy by better selecting among pre-existing reasoning paths, without significantly improving Pass@N rates (Yue et al., 2025). While our models’ comparatively lower Pass@1 scores are expected due to our exclusive use of SFT without a subsequent RL phase, the significant improvement in Pass@N for our fine-tuned models is noteworthy. This result suggests that our data successfully introduces novel reasoning patterns, thereby expanding the model’s solution space and increasing its problem-solving potential. This expanded capability provides a strong foundation for future work to apply RL and convert these reasoning skills into higher Pass@1 performance.
Table 9: Performance on mathematical reasoning benchmarks, including Pass@N results. The best results in each column are in bold, and the second-best are underlined. The models trained on DLR-Web+Book show marked improvements in Pass@N, indicating an expanded reasoning capability.
Model MATH-500 AIME 2024 AIME 2025
Pass@1 CoT-SC Pass@N Pass@1 CoT-SC Pass@N Pass@1 CoT-SC Pass@N
[…]
6.4 Data Scaling Effects
To validate the scalability of our data synthesis methodology, we conducted a series of experiments to evaluate the reasoning capabilities of models trained on progressively larger datasets synthesized by our approach. The results, illustrated in Figure 2, demonstrate a clear and positive correlation between the volume of synthetic data and model performance across multiple challenging reasoning benchmarks.
The consistent improvement across this diverse set of benchmarks validates that our synthesis process is not overfit to a specific domain or benchmark but rather enhances a general, robust reasoning capability. These strong scaling laws confirm that our method provides a reliable pathway to achieving superior model performance through data augmentation, and future work can leverage our proposed design logics to synthesize even larger datasets for continued improvement.
Figure 2: The model’s performance on four key benchmarks (AIME 2025, MMLU-Pro, GPQA-Diamond, and SuperGPQA) as a function of the training data quantity. The performance metric is Pass@N for AIME 2025, while all others are Pass@1.
6.5 Comparison with Baseline Datasets
We conducted a comparative analysis between our synthetically generated data and other prominent open-source synthetic datasets. To ensure a fair comparison, given the varying sizes of the original datasets, we randomly sampled an equal number of 304,181 instances from each dataset for the fine-tuning experiments. For the purpose of equitable comparison, we regenerated long CoT responses for all baseline methods (including WebInstruct (Full) and NaturalReasoning) using the same Qwen3-235B-A22B-Thinking-2507-FP8 model.
As shown in Table 10, our Design-Logic-Reasoning datasets consistently outperform WebInstruct (Full) and NaturalReasoning across all benchmarks. Specifically, DLR-Book achieves the best results on MMLU, MMLU-Pro, and GPQA-Diamond, while DLR-Web attains the highest performance on SuperGPQA and the second-best on other benchmarks. These results validate the superior quality and effectiveness of our data generation approach compared to existing methods.
Table 10: Comparison with other doc-centric synthetic datasets on the Qwen3-8B-Base model. All experiments were conducted using a subsample of 304,181 instances from each dataset. The best results in each column are in bold, and the second-best are underlined.
Dataset MMLU MMLU-Pro GPQA-Diamond SuperGPQA
Pass@1 Pass@1 Pass@1 CoT-SC Pass@1
[…]
6.6 Ablation Studies
We conducted an ablation study on our question-synthesis pipeline to quantify the contributions of two key components: coarse-to-fine matching between design logics and corpus segments, and the use of explicit design logics. To ensure a fair comparison, all ablations synthesize questions from the same book corpus by uniformly sampling 304,181 text segments. Where applicable, we used identical retrieval with Qwen3-Embedding-4B and generated model responses with the same model, Qwen3-235B-A22B-Thinking-2507-FP8.
  • w/o Coarse Ranking: This setting bypasses the semantic similarity-based retrieval of the top-5 relevant design logics. Instead, the LLM is prompted to select the most suitable logic from a set of 5 randomly chosen logics.
  • w/o Fine Ranking: This setting removes the LLM-based re-ranking stage. The single most similar design logic (the top-1 result from the coarse retrieval) is used directly to generate the question.
  • w/ Example Questions: This setting replaces our use of abstract design logics. Instead of being guided by a logic, the LLM is prompted to generate a new question by imitating the style and structure of the most suitable exemplar. This exemplar is selected from a set of 5 relevant example questions that are retrieved from the question bank.
  • DESIGNER: This is our complete method, which includes coarse retrieval of the top-5 design logics by similarity, followed by LLM-based fine selection and synthesis based on “design logic”.
Table 11 summarizes the results of our ablation studies. Our complete “DESIGNER” approach consistently outperforms ablated configurations across most benchmarks. The removal of either coarse-grained retrieval (w/o Coarse Ranking) or fine-grained LLM re-ranking (w/o Fine Ranking) causes a noticeable performance drop, confirming the value of each stage in our two-step matching process. The “w/o Fine Ranking” configuration, in particular, performs well on the MATH benchmark, suggesting the initial retrieval is highly effective at identifying relevant logics. However, the consistent improvements across other benchmarks validate the full coarse-to-fine pipeline’s benefit for handling complex, multidisciplinary reasoning. Finally, the superior performance of design logics over concrete “Example Questions” confirms our hypothesis that explicit logical structures are more accurate and robust guides for high-quality question synthesis.
Table 11: Ablation study of our data synthesis pipeline. All experiments were conducted on the Qwen3-8B-Base model using data synthesized from the same book corpus. The best results in each column are in bold, and the second-best are underlined.
Method MATH-500 MMLU GPQA-Diamond SuperGPQA
Pass@1 CoT-SC Pass@1 Pass@1 CoT-SC Pass@1
[…]
6.7 Effect of Source Corpus Quality on Synthesized Data
Table 12: Effect of source-corpus quality on SFT outcomes with equal per-discipline and total size. All experiments were conducted on the Qwen3-8B-Base model. The best results in each column are in bold.
Source Corpus MATH-500 MMLU MMLU-Pro GPQA-Diamond SuperGPQA
Pass@1 CoT-SC Pass@1 Pass@1 Pass@1 CoT-SC Pass@1
[…]
It is widely accepted that book corpora are of higher quality than web corpora. To test how source corpus quality affects our data synthesis method, we conducted a controlled SFT experiment. We compared data synthesized from a high-quality book corpus to that from a web corpus. For a fair comparison, we ensured the disciplinary data distribution was identical across both datasets. For disciplines with limited data in the FineFineWeb corpus, we used all available instances. For other disciplines, we used random sampling. This process resulted in two final datasets of equal size (282,857 instances each) and identical per-discipline distribution.
As shown in Table 12, the results validate our hypothesis. The model fine-tuned on data from the book corpus consistently outperforms the one trained on data from the web corpus across almost all benchmarks. This performance gain is most notable in complex reasoning tasks like GPQA-Diamond and SuperGPQA, confirming that our synthesis method benefits from higher-quality source material. However, our method also proves robust to variations in source quality, as the performance gap between the two corpora is not substantial. This indicates that our approach effectively synthesizes high-quality questions to improve model performance, regardless of the initial corpus quality.
7 Related Work
Query-Centric Data Synthesis
The core idea of query-centric data synthesis is to generate new training data by iteratively expanding or evolving an existing query pool. Self-Instruct (Wang et al., 2023) samples questions from an existing query pool and leverages an LLM to generate new question-answer pairs, which are then used to fine-tune the model. Wizard LM (Xu et al., 2023) (Luo et al., 2025) (Luo et al., 2024) and Auto Evol-Instruct (Zeng et al., 2024) employ instruction evolution to produce more complex and diverse data. CoT-Self-Instruct (Yu et al., 2025) integrates the CoT reasoning mechanism into the instruction generation process, enhancing the quality of synthetic data through a “think-first, then-generate” approach. Prismatic Synthesis (Jung et al., 2025) pushes the definition of data diversity from superficial textual features to a more fundamental level of model behavior, aiming to maximize the diversity of model-induced gradients. The SPARQ method (Havrilla et al., 2025) introduces a framework for evaluating diversity and difficulty in data synthesis, arguing that high-difficulty data improves in-distribution performance while diversity enhances out-of-distribution generalization. Although these query-centric methods attempt to optimize the difficulty and diversity of synthetic data, they are still limited by the initial query pool and the distribution biases of the model used for synthesis, making it difficult to cover questions across various disciplines.
Doc-Centric Data Synthesis
Doc-centric data synthesis methods start with a large amount of unstructured (e.g., CC, books) or structured (e.g., knowledge graphs) textual data. They generate question-answer pairs by directly extracting or inferring them from documents, thereby ensuring the synthetic data is closely tied to specific domain knowledge and facts. UltraChat  (Ding et al., 2023) generates questions about world knowledge from knowledge sources like Wikidata and meta topics pre-defined by human experts (e.g., technology, health). Humpback (Li et al., 2024) also uses a large web corpus as a data source, training a “backward model” to infer an instruction that would lead to a given document. The KPDDS method (Huang et al., 2025) constructs a Topic-level Co-occurrence Probability Matrix (TCPM) based on extracted topics and knowledge points, guiding data synthesis by sampling from this matrix. MAmmoTH2 (Yue et al., 2024) mines high-quality questions directly from existing web content using a three-step “Recall-Extract-Refine” pipeline, but the difficulty of its extracted questions is low. NaturalReasoning (Yuan et al., 2025) also synthesizes data from a large-scale pre-training corpus, but its process is more focused on generating high-difficulty reasoning questions, though the diversity of its synthetic questions is limited by the single prompt. Our work uniquely introduces the concept of design logic, analogizing the doc-centric data synthesis process to a human teacher creating questions. By matching documents with appropriate design logic, we significantly increase the difficulty and diversity of the synthetic questions.
Reasoning Data Synthesis
Using synthetic long CoT data for data distillation is an effective way to improve the reasoning abilities of smaller models. DeepSeek (Guo et al., 2025) distills a series of smaller open-source models using high-quality data samples generated by DeepSeek-R1, significantly enhancing the reasoning capabilities of these smaller models. OpenMathReasoning (Moshkov et al., 2025) extracts high-quality math problems from an online math forum (AoPS) and leverages powerful existing models to generate long CoT responses for these problems, significantly improving the models’ mathematical abilities. OmniThought (Cai et al., 2025) collects questions from the fields of mathematics, code, and science, and generates a long CoT dataset based on multiple teacher models. However, these methods primarily focus on generating high-quality reasoning processes for existing questions, and the open-source community still lacks high-quality, multidisciplinary original questions. Our method, in contrast, focuses on leveraging naturally available multidisciplinary documents to synthesize difficult and diverse multidisciplinary questions by introducing design logic.
8 Conclusion
This paper introduces DESIGNER, a novel design-logic-guided data synthesis pipeline designed to address the scarcity of high-quality, multidisciplinary reasoning data for LLMs. Our core innovation is the concept of “Design Logic”, which abstracts the strategic process human experts use to create challenging questions. By leveraging these logics, we generated two large-scale datasets, DLR-Book and DLR-Web, comprising millions of complex questions across 75 disciplines from diverse raw text sources. Our synthesized questions are shown to be significantly more difficult and diverse than those from existing methods, moving beyond simple factual recall to require deep, multi-step reasoning. We validated the effectiveness of our approach through extensive experiments, demonstrating that models fine-tuned with our data achieve substantial performance gains in multidisciplinary reasoning. Specifically, models trained on our datasets not only outperform those fine-tuned with existing multidisciplinary datasets of the same volume but can also surpass the multidisciplinary reasoning performance of the official Qwen3 models of the same size, validating the effectiveness of our method. This work provides a scalable paradigm for creating diverse and challenging reasoning data, paving a new path for advancing LLMs’ reasoning capabilities beyond domain-specific tasks.
Appendix A Design Logic and Question Quantity
Table 13: Number of Design Logics, Book Questions, and Web Questions by Discipline
Discipline Design Logic Book Question Web Question
Aerospace Science and Technology 980 15025 1815
Agricultural Engineering 966 14950 11312
Agricultural Resources and Environment 409 14978 1443
Animal Husbandry 121 15023 318
Archaeology 776 15061 2572
Architecture 801 14931 4235
Art and Design 998 15164 12720
Astronomy 1600 40010 14915
Atmospheric Sciences 973 15029 3416
Basic Medicine 2289 49930 7118
Bioengineering 1041 11216 1059
Biology 4654 200078 111988
Biomedical Engineering 790 15120 5439
Business Administration 3873 99739 46337
Chemical Engineering and Technology 1393 20014 5235
Chemistry 6430 199839 68211
Chinese History 784 14876 14564
Chinese Language and Literature 998 14984 22233
Civil Engineering 950 14899 7773
Clinical Medicine 2844 99735 77182
Computer Science and Technology 4674 99796 219474
Control Science and Engineering 1338 20013 3254
Ecology 968 15113 4285
Economics 3864 99906 20064
Education 1287 20018 9824
Electrical Engineering 2904 59937 36949
Electronic Science and Technology 1445 19880 18864
English and Foreign Languages 985 29908 1344
Environmental Science and Engineering 1478 19954 28226
Ethnology 210 14856 51
Food Science and Engineering 768 15073 2470
Forensic Medicine 295 15010 34
Geography 2812 59987 12295
Geological Resources and Geological Engineering 173 15058 374
Geology 987 14958 5447
Geophysics 982 20050 3736
History of Science and Technology 494 14991 111
Hydraulic Engineering 664 14973 649
Information Resources Management 356 14968 69
Information and Communication Engineering 2367 39825 4777
Instrument Science and Technology 687 15017 378
Journalism and Communication 932 14847 4326
Law 2706 80110 62699
Management Science and Engineering 1317 20020 10440
Marine Sciences 852 14968 1228
Materials Science and Engineering 1949 40121 12263
Mathematics 9884 299464 181537
Mechanical Engineering 2955 59998 55356
Mechanics 1752 40046 4012
Mining Engineering 282 14986 438
Naval Architecture and Ocean Engineering 571 15034 515
Nuclear Science and Technology 1448 19988 1981
Nursing 767 15078 2228
Optical Engineering 927 14995 3272
Petroleum and Natural Gas Engineering 571 15040 988
Pharmacy 1829 39921 12756
Philosophy 4363 100029 128004
Physical Education 890 14912 10712
Physics 5768 199771 104982
Political Science 3268 59686 52586
Power Engineering and Engineering Thermophysics 704 14918 1199
Psychology 4336 99502 95585
Public Administration 960 14970 26669
Public Health and Preventive Medicine 1949 39820 19649
Remote Sensing Science and Technology 344 8081 271
Safety Science and Engineering 835 14986 409
Sociology 2093 39941 32361
Statistics 3388 59950 24806
Stomatology 592 15002 456
Surveying and Mapping Science and Technology 509 14964 358
Textile Science and Engineering 187 14978 325
Transportation Engineering 882 14929 4956
Urban and Rural Planning 576 15052 89
Veterinary Medicine 547 14918 3022
World History 987 39703 5503
Table 13: (Continued) Number of Design Logics, Book Questions, and Web Questions by Discipline
Appendix B Prompts
Prompt for Discipline Classification You are a professional multidisciplinary data labeling expert specializing in the classification of multidisciplinary academic questions. Please select the ONE most relevant label from the given list of discipline labels for the input question data. For question data that you cannot determine, use the ‘‘Unknown Discipline’’ label. Please directly output ‘‘labels’’: ‘‘(the label you selected)’’.
# List of Discipline Labels:
[‘Mathematics’, ‘Biology’, ‘Chemistry’, ‘Physics’, ‘Computer Science and Technology’, ‘Philosophy’, ‘Psychology’, ‘Business Administration’, ‘Clinical Medicine’, ‘Economics’, ‘Law’, ‘Political Science’, ‘Statistics’, ‘Electrical Engineering’, ‘Geography’, ‘Mechanical Engineering’, ‘Basic Medicine’, ‘Information and Communication Engineering’, ‘Sociology’, ‘Materials Science and Engineering’, ‘Pharmacy’, ‘Public Health and Preventive Medicine’, ‘Mechanics’, ‘Astronomy’, ‘World History’, ‘Bioengineering’, ‘English and Foreign Languages’, ‘Chemical Engineering and Technology’, ‘Electronic Science and Technology’, ‘Environmental Science and Engineering’, ‘Nuclear Science and Technology’, ‘Control Science and Engineering’, ‘Management Science and Engineering’, ‘Education’, ‘Geophysics’, ‘Art and Design’, ‘Agricultural Engineering’, ‘Aerospace Science and Technology’, ‘Atmospheric Sciences’, ‘Chinese Language and Literature’, ‘Civil Engineering’, ‘Ecology’, ‘Geology’, ‘Nursing’, ‘Optical Engineering’, ‘Public Administration’, ‘Journalism and Communication’, ‘Physical Education’, ‘Marine Sciences’, ‘Safety Science and Engineering’, ‘Architecture’, ‘Transportation Engineering’, ‘Power Engineering and Engineering Thermophysics’, ‘Food Science and Engineering’, ‘Archaeology’, ‘Biomedical Engineering’, ‘Chinese History’, ‘Veterinary Medicine’, ‘Instrument Science and Technology’, ’Hydraulic Engineering’, ‘Stomatology’, ‘Urban and Rural Planning’, ‘Petroleum and Natural Gas Engineering’, ‘Naval Architecture and Ocean Engineering’, ‘Surveying and Mapping Science and Technology’, ‘History of Science and Technology’, ‘Agricultural Resources and Environment’, ‘Remote Sensing Science and Technology’, ‘Information Resources Management’, ‘Mining Engineering’, ‘Forensic Medicine’, ‘Ethnology’, ‘Textile Science and Engineering’, ‘Geological Resources and Geological Engineering’, ‘Animal Husbandry’, ‘Other’, ‘Non-disciplinary’, ‘Unknown Discipline’]
# Example 1
Input: ‘‘Consider a photon traveling at the speed of light. How does the photon experience space, and what are the implications of relativistic beaming on its perception of spatial dimensions? Provide a detailed explanation, including any relevant mathematical derivations and physical principles.’’
Output: ‘‘labels’’: ‘‘Physics’’
# Example 2
Input: ‘‘A heavy pole, of mass M and length L, is freely hinged to a wall at the point O. A rope connects the other end of the pole, B, to a fixed point A on the wall above O. The system is in equilibrium, with the pole making an angle of $\theta$ with the horizontal, and the rope making an angle of $\alpha$ with the horizontal. Explore how the system’s parameters (M, L, $\theta$, $\alpha$) affect its equilibrium and stability.’’
Output: ‘‘labels’’: ‘‘Mechanics’’
# Example 3
Input: ‘‘If John rented a car for \$150 and had to buy 8 gallons of gas at \$3.50 per gallon to fill it up, and the final expense is \$0.50 per mile, how much did it cost him to drive 320 miles?’’
Output: ‘‘labels’’: ‘‘Mathematics’’
# Input Question Data
Input: ‘‘{text}’’
Output:
Figure 3: The few-shot prompt used for discipline classification. The model is presented with a list of disciplines and three examples, and is then asked to classify a given question, which replaces the {text} placeholder.
Prompt for Difficulty Classification You are an expert in education and examination, specializing in classifying the difficulty levels of multidisciplinary questions. For the given question, please evaluate its difficulty based on the complexity and length of the reasoning required to answer it. Label it as one of the following: **Easy**, **Medium**, **Hard**, or **Very Hard**. Please directly output ‘‘Difficulty: (Your chosen label)’’.
# Example 1
Input: ‘‘Consider a photon traveling at the speed of light. How does the photon experience space, and what are the implications of relativistic beaming on its perception of spatial dimensions? Provide a detailed explanation, including any relevant mathematical derivations and physical principles.’’
Output: ‘‘Difficulty: Very Hard’’
# Example 2
Input: ‘‘A heavy pole, of mass M and length L, is freely hinged to a wall at the point O. A rope connects the other end of the pole, B, to a fixed point A on the wall above O. The system is in equilibrium, with the pole making an angle of $\theta$ with the horizontal, and the rope making an angle of $\alpha$ with the horizontal. Explore how the system’s parameters (M, L, $\theta$, $\alpha$) affect its equilibrium and stability.’’
Output: ‘‘Difficulty: Hard’’
# Example 3
Input: ‘‘If John rented a car for \$150 and had to buy 8 gallons of gas at \$3.50 per gallon to fill it up, and the final expense is \$0.50 per mile, how much did it cost him to drive 320 miles?’’
Output: ‘‘Difficulty: Easy’’
# Given Question
Input: ‘‘{text}’’
Output:
Figure 4: The few-shot prompt used for difficulty classification. The model is presented with three examples and is then asked to classify the difficulty of a given question, which replaces the {text} placeholder.
Prompt for Question Type Classification You are an expert in education and examination, specializing in classifying question types. For the given question, please evaluate its question type and label it as one of the following: **Problem-solving question**, **Multiple-choice question**, **Proof question**, or **Other question types**. For any question that you cannot determine, use the ‘‘Other question types’’ label. Please directly output ‘‘Question type: (Your chosen label)’’.
# Example 1
Input: ‘‘Determine the number of $k$-letter sequences composed of the letters AA and $B$ such that the sequence contains at least two consecutive AA’s.’’
Output: ‘‘Question type: Problem-solving question’’
# Example 2
Input: ‘‘Consider the function $f(x)=\frac{e^{x}}{x}$. The value of the integral $I=\int_{1}^{\infty}\left(\frac{e^{x}}{x}-\frac{e^{-x}}{x}\right)dx$ is ___.’’
Output: ‘‘Question type: Other question types’’
# Example 3
Input: ‘‘Given that $a\in\{-1,2,\frac{1}{2},3,\frac{1}{3}\}$, if $f(x)=x^{a}$ is an odd function and is monotonically increasing on $(0,+\infty)$, then the possible values of the real number $a$ are ( ).
 A: 1,3-1,3
 B: $\frac{1}{3},3$
 C: $1,\frac{1}{3},3$
 D: $\frac{1}{3},\frac{1}{2},3$’’
Output: ‘‘Question type: Multiple-choice question’’
# Given Question
Input: ‘‘{text}’’
Output:
Figure 5: The few-shot prompt used for question type classification. The model is presented with three examples and is then asked to classify the type of a given question, which replaces the {text} placeholder.
Prompt for Reasoning-Oriented Filtering You will be provided with text from the internet.
Evaluate the following text extract for its potential usefulness for studying reasoning process. Use the following 5-point scoring system described below. Start from 0, points are accumulated based on the satisfaction of each criterion:
(1) Add 1 point if the extract contains any reasoning or thinking process.
(2) Add 1 point if the extract contains any explicit subgoal setting, where the writer breaks down the problem into smaller, intermediate goals. Subgoal setting might look like:
 • ‘‘First, we need to find ..., then we can determine ...’’
 • ‘‘To solve ..., let’s first ..., then ...’’
 • ‘‘Let’s tackle ... in three parts: (1) ..., (2) ..., and (3) ...’’
 • ‘‘To ..., I’$l$ first ..., then ...’’
(3) Add 1 point if the extract contains any verification steps. We want to mark instances where the writer explicitly checks their own work, such as by comparing the result to a known value or by checking the result of a calculation. Verification steps might look like:
 • ‘‘Let’s check ...’’
 • ‘‘To verify this is correct, I’$l$ ...’’
 • ‘‘Let’s test ... with a simple case: ...’’
 • ‘‘To ensure this solution is valid, I’$l$ check if ...’’
(4) Add 1 point if the text contains any backtracking behavior, where the writer realizes a path won’t work and explicitly goes back to try a different approach. An example of backtracking is: ‘‘Let me try again’’, ‘‘Wait’’, ‘‘I made a mistake’’, or ‘‘we need to try a different sequence of operations’’. We want to mark instances where the writer abandons a thought and backtracks to a previous computation.
(5) Add 1 point if the text contains any backward-chaining behavior, where the writer is working towards a goal but starts from the goal and works backward. It might like:
 • ‘‘To solve ..., let’s start with what we want to prove: ...Let’s verify this.’’
 • ‘‘If we want to find ..., let’s start with the desired result and work backward.’’
 • ‘‘To determine ..., I know the result ... Working backward from this final state using
# Task Format
Format your response in markdown as follows:
## Thoughts
[Brief description describing what behavior was noticed and where subgoal setting may have occurred, less than 100 words]
## Final score
[total points]
# Text to evaluate for reasoning degree
{text}
# Response
Figure 6: The prompt used for the reasoning-oriented filtering task. It defines a five-level scoring rubric to assess the usefulness of a text (which replaces the {text} placeholder) for studying reasoning.
Prompt for Design Logic Extraction You are an expert educator and a specialist in exam question design. Below, I have provided an exam question. Your task is to deduce the thought process of the question designer. Analyze how they constructed this question based on the relevant knowledge points. You need to go beyond the specific details of the question and its knowledge points to abstract and summarize the underlying design logic and principles behind the question.
The goal is for me to be able to use this abstracted design logic to create other high-quality, challenging questions that require complex logical reasoning for different knowledge points and source materials.
**Finally, you must organize the abstracted question-design logic you have summarized into English Mermaid format.**
--- Analyze the Question Design Logic from the Following Question ---
**Question:**
{text}
Figure 7: The prompt used for design logic extraction. The model is instructed to reverse-engineer the thought process behind a given question (which replaces the {text} placeholder) and to structure the abstracted logic in Mermaid format.
Instruction for Design Logic Retrieval Given a book snippet, retrieve the most suitable question-design logic in Mermaid format for creating a challenging exam question from the book snippet.
Figure 8: The task-specific instruction used for retrieving the most suitable design logic for a given text segment. Embeddings for both text segments and design logics are computed under this instruction using the Qwen3-Embedding-4B model, enabling similarity-based retrieval.
Prompt for Question Synthesis You are an expert in the field of education and examination design, and you are writing exam questions. Your task is to use the provided text to generate a high-quality exam question. Please follow the steps below to generate an English exam question and a reference answer:
**1. Create an Exam Question:**
- Based on the provided source text, write a challenging exam question at the graduate-level or above.
- Below are five question-design logics provided in Mermaid format. You need to select the most suitable question-design logic for creating a challenging question from the source text, and then strictly follow the corresponding question-design logic and steps to create a challenging question. Please record which design logic you used (by number) and output the corresponding numeric ID in the ‘‘id’’ field of the JSON below.
- The question should require critical thinking and test deep understanding and problem-solving skills, not just simple fact recall.
- The question must be self-contained and answerable without using the source text. If the question you write requires an answer based on the content of the source text, you must include the corresponding content and information from the source text within the question itself to make it self-contained.
- Ensure the question is self-contained, clear, without missing information or ambiguity, and has a correct answer.
- For multiple-choice questions, you should first analyze and determine the answer, then design the options to ensure that one specific option is the correct answer. The questions you design need to include as many options as possible (four or more). Do not be limited to only four options (A, B, C, D).
**2. Provide the Reference Answer:**
- Use the information in the source text to write a concise and accurate reference answer to the question you just created.
[…]
**At the end of your response, please organize your results into the following JSON format:**
‘‘exam_question’’: ‘‘*(Your question goes here)*’’,
‘‘reference_answer’’: ‘‘*(Your reference answer goes here)*’’,
‘‘id’’: ‘‘*(The ID of the logic you selected goes here)*’’
**--- Question-Design Logic 1 ---**
‘‘‘Mermaid
{logic1}
**--- Question-Design Logic 2 ---**
‘‘‘Mermaid
{logic2}
**--- Question-Design Logic 3 ---**
‘‘‘Mermaid
{logic3}
**--- Question-Design Logic 4 ---**
‘‘‘Mermaid
{logic4}
**--- Question-Design Logic 5 ---**
‘‘‘Mermaid
{logic5}
**--- Source Text for Question Creation ---**
{text}
Figure 9: The prompt used for question synthesis. The LLM is provided with a source text and five candidate design logics retrieved via semantic similarity. It is instructed to select the most suitable logic and strictly follow it to generate a graduate-level question and a corresponding reference answer, structured in a JSON format. The placeholders {logic1} through {logic5} and {text} are replaced with specific design logics and the source text, respectively.

### v2

@@
 Large language models (LLMs) have achieved remarkable success in many natural language tasks but still struggle with complex, multi-step reasoning, particularly across diverse disciplines.
-Existing reasoning datasets often either lack disciplinary breadth or the structural depth necessary to elicit robust reasoning behaviors.
-We propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages naturally available, extensive raw documents (book corpus and web corpus) to generate multidisciplinary challenging questions.
-A core innovation of our approach is the introduction of a Design Logic concept, which mimics the question-creation process of human educators.
+Existing reasoning datasets often lack disciplinary breadth, reasoning depth, and diversity, and lack guiding principles for question synthesis.
+We propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages naturally available, extensive raw documents (e.g., book corpus and web corpus) to generate multidisciplinary challenging questions.
+We introduce the concept of “design logic” and instruct LLMs to mimic human educators’ question-creation process, enabling automated synthesis of large-scale, high-difficulty questions.
 We use LLMs to reverse-engineer and abstract over 120,000 design logics from existing questions across various disciplines.
-By matching these design logics with disciplinary source materials, we are able to create reasoning questions that far surpass the difficulty and diversity of existing datasets.
-Based on this pipeline, we synthesized two large-scale reasoning datasets that span 75 disciplines: Design-Logic-Reasoning-Book (DLR-Book), containing 3.04 million challenging questions synthesized from the book corpus, and Design-Logic-Reasoning-Web (DLR-Web), with 1.66 million challenging questions from the web corpus.
-Our data analysis demonstrates that the questions synthesized by our method exhibit substantially greater difficulty and diversity than those in the baseline datasets.
-We validate the effectiveness of these datasets by conducting SFT experiments on the Qwen3-8B-Base and Qwen3-4B-Base models.
-The results show that our dataset significantly outperforms existing multidisciplinary datasets of the same volume.
-Training with the full datasets further enables the models to surpass the multidisciplinary reasoning performance of the official Qwen3-8B and Qwen3-4B models.
+By matching these design logics with source documents, we are able to create reasoning questions that far surpass the difficulty and diversity of existing datasets.
+Using this pipeline, we synthesized two large-scale reasoning datasets that span 75 disciplines: DLR-Book (3.04 million questions from the book corpus) and DLR-Web (1.66 million questions from the web corpus).
+Data analysis indicates that the questions synthesized by our method exhibit greater difficulty and diversity compared to those in the baseline datasets.
+We validate our synthesized data through supervised fine-tuning (SFT) on the Qwen3 and Llama3 model families.
+Our data substantially enhances their multidisciplinary reasoning capabilities, outperforming existing datasets.
+Notably, after SFT on our datasets, the base versions of these models even surpass their official instruction-tuned counterparts.
 [^1] Project page: https://attention-is-all-i-need.github.io/Design-Logic-Reasoning
@@
 1 Introduction
-Large language models (LLMs) have demonstrated exceptional capabilities in various natural reasoning tasks (Brown et al., 2020; Kaplan et al., 2020; Hoffmann et al., 2022; Chowdhery et al., 2023; OpenAI, 2023), such as mathematics and coding, especially when utilizing long chain-of-thought (CoT) techniques (Guo et al., 2025; Moshkov et al., 2025; Cai et al., 2025).
-However, their performance in a wide range of university-level discipline-specific reasoning questions still falls short of human experts.
-A major bottleneck is the lack of a large-scale, high-quality, and diverse multidisciplinary reasoning data.
-Existing datasets primarily focus on math and programming, training models on questions collected from math and coding competition platforms (Moshkov et al., 2025; Cai et al., 2025).
-This approach heavily relies on the massive open question resources in these specific domains, but many other academic disciplines lack such abundant resources, which limits the development of LLMs’ multidisciplinary reasoning capabilities.
-Data synthesis is an effective way to address data scarcity by generating a large number of reasoning questions from existing LLMs (Wang et al., 2023; Xu et al., 2023; Jung et al., 2025; Havrilla et al., 2025).
-Current question synthesis methods can be broadly categorized into query-centric and doc-centric approaches.
-Query-centric methods, centered on an initial “seed” question pool, iteratively evolve existing questions.
-These methods generate more complex and diverse new questions by rewriting existing ones, adding constraints (e.g., Evol-Instruct), or incorporating a chain-of-thought (Wang et al., 2023; Xu et al., 2023; Yu et al., 2025).
-However, the breadth and depth of the generated questions are severely limited by the quality and coverage of the initial seed pool.
-They are also susceptible to the biases of the generative model itself, making it difficult to transcend domain limitations and create high-quality questions across a wide range of disciplines.
-In contrast, doc-centric methods start with a vast collection of unstructured (e.g., web pages, books) or structured (e.g., knowledge graphs) documents (Yue et al., 2024; Yuan et al., 2025; Huang et al., 2025).
-By directly extracting or reasoning out question-answer pairs from these documents, this approach ensures that the synthesized data is closely tied to specific domain knowledge and facts, theoretically allowing it to cover a broader range of disciplines.
-The main challenge for these methods is to effectively control the difficulty and diversity of the generated questions and prevent them from degrading into simple surface-level factual recall tasks.
-To address these issues, as shown in Figure 1, we propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages readily available, large-scale, multidisciplinary raw documents (e.g., book corpus and web corpus) to synthesize a massive number of multidisciplinary challenging questions.
-The core innovation of our method lies in the introduction of a novel concept: “Design Logic”.
+Large language models (LLMs) have demonstrated exceptional capabilities in various natural reasoning tasks (Chowdhery et al., 2023; OpenAI, 2023), such as mathematics and coding (Trinh et al., 2024; Zhu et al., 2024), especially when utilizing long chain-of-thought (CoT) techniques (Wei et al., 2022; Jaech et al., 2024; Guo et al., 2025).
+However, they still lag behind human experts on university-level, discipline-specific reasoning (Phan et al., 2025), largely due to the scarcity of large-scale, high-quality, and diverse training data.
+Existing datasets focus mainly on math and programming, drawing on competition platforms rich in open-ended questions (Moshkov et al., 2025; Cai et al., 2025), while most other disciplines lack comparable resources.
+This scarcity limits the development of LLMs’ multidisciplinary reasoning capabilities.
+Data synthesis with LLMs is an effective solution to data scarcity (Wang et al., 2024a).
+Existing question synthesis methods fall into two categories: query-centric and doc-centric.
+Query-centric approaches expand seed questions through rewriting, added constraints (e.g., Evol-Instruct (Xu et al., 2023)), or incorporating chain-of-thought reasoning (Wang et al., 2023; Yu et al., 2025), but they are limited by the coverage of the seed pool and inherent model biases.
+Doc-centric methods instead generate questions from unstructured (e.g., web, books) or structured (e.g., knowledge graphs) sources (Yue et al., 2024; Yuan et al., 2025; Huang et al., 2025), ensuring broad disciplinary coverage grounded in authentic knowledge.
+However, they struggle to control difficulty and diversity, often degenerating into factual recall.
+Figure 1: Left: The procedure by which human experts construct questions reflects a “Design Logic”: a systematic sequence of deliberate steps that transforms fundamental knowledge points into complex, context-rich questions requiring multi-stage reasoning.
+Right: DESIGNER emulates this process by matching logics to raw corpus and synthesizing diverse, multidisciplinary questions.
+To address these issues, we propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages large-scale, multidisciplinary raw documents (e.g., book corpus and web corpus) to synthesize challenging questions across diverse disciplines.
+The central insight of our approach is the notion of “Design Logic”, which encapsulates how human experts transform knowledge into complex, reasoning-intensive exam questions.
 We observed that when human education experts design challenging and insightful questions, they don’t merely state facts.
-Instead, they follow a structured design process, such as: identifying key knowledge points $\rightarrow$ constructing a scenario $\rightarrow$ designing a reasoning path $\rightarrow$ pre-setting distractors.
-We contend that this process embodies a “Design Logic”: a sequence of deliberate steps that transforms fundamental knowledge points into complex, context-rich questions that demand multi-stage reasoning.
+Instead, they follow a structured design process, as illustrated in Figure 1.
 This design logic itself is a valuable and transferable form of “meta-knowledge”.
-Therefore, our pipeline is not limited to directly extracting questions from documents.
-Instead, we first use a powerful LLM to reverse-engineer and abstract tens of thousands of “meta design logics” from existing high-quality question banks across various disciplines.
-During the synthesis phase, we precisely match the most suitable design logic to a source document and instruct the LLM to strictly follow the intrinsic reasoning steps of that logic to construct a brand-new question.
-Specifically, our approach follows a refined pipeline.
-First, we perform comprehensive processing on the massive book corpus and web corpus through multi-dimensional labeling and filtering (e.g., by discipline, readability, and knowledge value) to build a high-quality source material library that covers a wide range of disciplines.
-Next, from a private question bank containing hundreds of millions of questions, we use clustering and sampling to select a subset of questions that are both difficult and diverse.
-We then instruct a powerful LLM to summarize and abstract over 120,000 structured “design logics” from this subset.
-In the question synthesis phase, we pioneered a two-stage “retrieve-and-generate” matching mechanism.
-First, we use vector similarity to retrieve a rough, coarse-grained set of the most relevant candidate logics for each source document from our design logic library.
-Then, an LLM performs a fine-grained evaluation to select the optimal logic and strictly applies this design logic’s steps to the source document content, generating a new and challenging reasoning question.
+Specifically, our pipeline is illustrated in Figure 2.
+First, we process large-scale book and web corpora with multi-dimensional labeling and filtering (discipline, readability, knowledge value) to construct a high-quality source material library.
+From a question bank of hundreds of millions, we cluster and sample a diverse set of difficult questions, from which an LLM reverse-engineers and abstracts over 120K structured “design logics” to construct a reusable design logic library.
+In question synthesis, we adopt a two-stage retrieve-and-generate mechanism: (1) vector similarity retrieves coarse candidate logics for each source document, and (2) an LLM performs a fine-grained evaluation to select the optimal logic and generates a reasoning question from the source document by strictly following its steps.
+This approach addresses the absence of guiding principles in prior data synthesis methods, enabling the automated generation of a large number of diverse and high-difficulty exam questions while reducing reliance on expensive manual creation.
 The main contributions of this paper can be summarized as follows:
   • We propose a data engineering pipeline for synthesizing challenging questions from raw text corpora.
-Based on this pipeline, we constructed two large-scale reasoning datasets: Design-Logic-Reasoning-Book (DLR-Book), containing 3.04 million challenging questions synthesized from the book corpus, and Design-Logic-Reasoning-Web (DLR-Web), with 1.66 million challenging questions from the web corpus.
-These datasets expand beyond common disciplines like math to cover 75 disciplines (including STEM, humanities, social sciences, arts, as well as applied and professional fields) to better train LLMs’ multidisciplinary reasoning capabilities.
-  • We introduce a novel, scalable data synthesis paradigm based on “Design Logic”, which imitates the wisdom of human experts in question creation and significantly enhances the reasoning depth and diversity of doc-centric synthesized questions.
-Our data analysis demonstrates that the questions synthesized by our method exhibit substantially greater difficulty and diversity than those in the baseline datasets.
-  • Our comprehensive comparative and ablation experiments on the Qwen3-8B-Base and Qwen3-4B-Base models (Yang et al., 2025a) show that training with our data not only significantly improves the model’s multidisciplinary reasoning capabilities but also outperforms existing multidisciplinary datasets with the same data volume.
-Training with our full dataset allows the resulting models to surpass the performance of the official Qwen3 models of the same size.
+Using this pipeline, we constructed two large-scale reasoning datasets: DLR-Book (3.04 million questions from the book corpus) and DLR-Web (1.66 million questions from the web corpus).
+These datasets span 75 disciplines, including STEM, humanities, social sciences, applied and professional fields, and arts, extending beyond common disciplines.
+  • Inspired by the question-creation process of human teachers, we propose a question synthesis method guided by “Design Logic” and construct a corresponding library of design logics.
+Our data analysis indicates that the questions synthesized by our method exhibit greater difficulty and diversity compared to those in the baseline datasets.
+  • We validate the effectiveness of our synthesized data through comprehensive comparative and ablation experiments on the Qwen3 (Yang et al., 2025a) and Llama3 (Grattafiori et al., 2024) model families.
+The results demonstrate that the data synthesized by our method significantly enhances the multidisciplinary reasoning capabilities of LLMs.
+2 Data Curation
 Figure 2: The Design-Logic-Guided Multidisciplinary Data Synthesis Pipeline.
-2 Overview
-We propose a data engineering pipeline for synthesizing challenging questions from raw text corpora (web corpus and book corpus). The complete pipeline is illustrated in Figure 1.
-The first stage involves data curation, a meticulous process tailored to each of our three primary data sources and unified under a comprehensive 75-discipline taxonomy:
-  • Web Corpus: We apply reasoning-oriented filtering and relabeling to a large web dataset, retaining only texts that are most useful for studying reasoning processes.
-  • Book Corpus: This corpus undergoes preprocessing, discipline labeling, and a quality-prioritized sampling method to ensure the selection of high-readability and high-helpfulness text segments.
-  • Question Bank: We annotate our question bank with multi-dimensional labels (discipline, difficulty, type) and use a clustering-based approach to select a diverse and high-quality subset.
-The second stage focuses on question synthesis.
-We first extract underlying design logics from the sampled question subset.
-We then leverage a two-stage “retrieve-and-generate” mechanism to apply these reusable design logics to our curated text corpora.
-This process allows us to strictly follow the steps of each logic to construct millions of new reasoning questions, ensuring they are both challenging and diverse while mirroring the complexity of human-designed problems.
-Finally, we synthesize corresponding long CoT answers for each question, creating a valuable dataset for supervised fine-tuning.
-Based on this pipeline, we constructed two large-scale reasoning datasets: Design-Logic-Reasoning-Book (DLR-Book), which contains 3.04 million challenging questions synthesized from the book corpus, and Design-Logic-Reasoning-Web (DLR-Web), with 1.66 million questions from the web corpus.
-These datasets expand beyond common disciplines like mathematics to cover 75 disciplines (including STEM and the humanities and social sciences), thereby better supporting the training of LLMs for multidisciplinary reasoning.
-Additionally, we have included concise reference answers in the datasets, generated from the text corpora concurrently with the questions, as well as long CoT responses generated by Qwen3-235B-A22B-Thinking-2507-FP8.
-3 Data Curation
-3.1 Data Collection
+We curate three data sources for question synthesis: a proprietary question bank, a book corpus, and a web corpus, all aligned to a unified 75-discipline taxonomy (see Appendix A).
+Figure 2 (Phase 1) illustrates the overall data processing pipeline.
+2.1 Question Bank Processing
+We annotate over 150 million questions from the proprietary bank with discipline, difficulty, and type using Qwen3-30B-A3B (non-thinking mode), following prompts in Figure 7, Figure 8, and Figure 9, respectively.
+To obtain a high-quality and diverse subset, we compute embeddings with Qwen3-Embedding-4B and apply K-means clustering within each discipline (Ahmed et al., 2020), with cluster numbers determined by silhouette search.
+From each cluster, we draw an equal number of questions following a fixed difficulty ratio of 3:2:1 (Very Hard:Hard:Medium), and align per-discipline sizes with the overall bank distribution.
+This process resulted in a curated set of 132,409 questions for design-logic extraction.
+2.2 Book Corpus Processing
+This corpus is processed at the chapter level, with chapters over 5,000 words split into smaller blocks and deduplicated via MinHash.
+Discipline labels are assigned by a ModernBERT-large classifier fine-tuned for disciplinary classification (Warner et al., 2024).
+Readability is predicted with a BERT-based model (Turc et al., 2019) to filter incoherent or disorganized text, while helpfulness (0–5) is scored by the fineweb-edu-classifier (Lozhkov et al., 2024) to quantify educational value.
+Based on quotas proportional to discipline frequencies in both the book corpus and the question bank, segments with negative readability are removed, remaining candidates are ranked by helpfulness, and top items are selected until quotas are met.
+This procedure yields three million high-quality segments, with most scoring helpfulness $\geq 2$.
+2.3 Web Corpus Processing
+We apply reasoning-oriented filtering and discipline relabeling to the FineFineWeb corpus, scoring 6.5 billion texts with Qwen3-30B-A3B (non-thinking mode) using a five-level rubric (prompt in Figure 10) and retaining those with scores $\geq$3.
+The selected texts are then relabeled with the same model (prompt in Figure 7) to align with our 75-discipline taxonomy.
+3 Design-Logic-Guided Data Synthesis
+Figure 2 (Phase 2 and Phase 3) illustrates the overall data synthesis pipeline.
+3.1 Design Logic Extraction
+Human educators design exam questions through structured steps that transform knowledge points into complex challenges, rather than simple fact recall.
+A typical process involves identifying objectives, constructing contexts, designing reasoning paths, formulating answers, adding distractors, and validating the question.
+Solvers must engage in multi-step reasoning beyond memorization.
+Inspired by this, we propose a design-logic-based question synthesis method.
+Using the prompt in Figure 11, we instruct an LLM (DeepSeek-R1-0528) to analyze authentic questions and (i) infer the designer’s thought process, (ii) trace construction from knowledge points, and (iii) abstract underlying design principles, expressed in Mermaid format.
+This produces a reusable pool of design logics that guides new question generation from source materials.
+3.2 Design Logic Deduplication
+To enhance the diversity of design principles, we deduplicate extracted logics using semantic similarity.
+Each logic is embedded with Qwen3-Embedding-4B, and pairwise similarities yield a matrix $S\in\mathbb{R}^{n\times n}$.
+Within each discipline, we construct a graph $G=(V,E)$ where nodes represent logics and edges connect pairs with $S_{ij}\geq\tau$.
+Connected components in the graph correspond to redundant design logic groups.
+From each group, we retain the item with the highest similarity sum.
+With $\tau=0.85$, this graph-based deduplication procedure (Algorithm 1) yields 125,328 unique design logics, with per-discipline counts in Table C.
+We provide several examples of our synthesized design logic in Appendix J.
+3.3 Question Synthesis
+To avoid the combinatorial explosion resulting from exhaustively matching design logics with text segments, we adopt a retrieval-augmented approach.
+For each discipline-specific corpus, we compute cosine similarity between embeddings of a text segment $t$ and a design logic $d$ using Qwen3-Embedding-4B with task-specific instructions (Figure 12): $s(t,d)\,=\,\cos\!\bigl(\mathbf{e}(t),\,\mathbf{e}(d)\bigr)$.
+The top-5 logics with the highest similarity are retained as candidates.
+We then prompt DeepSeek-R1-0528 (Figure 13) to: (i) select the most suitable logic from the top-5 candidates, and (ii) synthesize a graduate-level exam question strictly following its steps.
+This two-stage process forms a coarse-to-fine ranking: similarity-based retrieval provides coarse recall, while the LLM refines the match to ensure precise alignment between text and logic, thereby improving question quality.
+For each question, the LLM also generates a concise reference answer.
+Question Deduplication and Decontamination
+We employ a two-stage filtering pipeline: (i) MinHash-based deduplication to remove near-duplicates, and (ii) 13-gram decontamination against all evaluation benchmarks to prevent leakage.
+Using the curated book and web corpora (Section 2), DeepSeek-R1-0528 generates one reasoning question per text segment.
+After filtering, the final dataset contains 3,040,620 questions from books and 1,658,541 from web sources.
+3.4 Response Synthesis
+To demonstrate that our synthesized questions can effectively elicit and transfer the long CoT capabilities of a reasoning model and improve the performance of models trained on this data, we employ Qwen3-235B-A22B-Thinking-2507-FP8 to generate a corresponding long CoT response for each synthesized question.
+These question-response pairs are then used for supervised fine-tuning (SFT).
+4 Data Analysis
+To assess the quality of our synthesized datasets (DLR-Book and DLR-Web), we conduct a quantitative analysis comparing their difficulty, diversity, and disciplinary distribution against the baseline datasets.
+The baseline datasets and benchmarks we used are detailed in Table 5 and Table 10, respectively.
+In all tables, we highlight the best value in boldface and the second-best with an underline.
+The distribution of question types for our synthesized datasets is detailed in Appendix E.
+4.1 Difficulty Analysis
+We utilized the Qwen3-30B-A3B-Instruct-2507 model with the prompt shown in Figure 8 to assign difficulty labels to both our datasets and the baseline datasets.
+To facilitate an intuitive comparison of difficulty levels, we further applied the same labeling procedure to several commonly used benchmarks.
+As shown in Figure 3, our datasets are significantly more difficult.
+Notably, the proportion of “Very Hard” questions in our datasets is substantially higher than that in all baseline datasets and benchmarks.
+In contrast, the proportion of “Easy” questions is negligible (0.72% in DLR-Web and 0.27% in DLR-Book).
+The detailed distribution of difficulty levels is reported in Table 7.
+Figure 3: Difficulty distributions of questions across different datasets and benchmarks.
+4.2 Diversity Analysis
+To quantify dataset diversity, we generate high-dimensional vector representations for 300,000 uniformly sampled questions from each dataset using the Qwen3-Embedding-4B model.
+We then compute the following five distance-based diversity metrics in the embedding space: Mean Cosine Distance (Yang et al., 2025b), Mean L2 Distance (Yang et al., 2025b), 1-Nearest Neighbor (1-NN) Distance (Stasaski & Hearst, 2022), Cluster Inertia (Du & Black, 2019), and Radius (Lai et al., 2020).
+Detailed definitions and formulas for these metrics are provided in Appendix G.
+Table 1: Results of semantic diversity metrics for our synthesized datasets and the baseline datasets. Higher values indicate better performance across all metrics.
+Dataset Mean Cosine Distance Mean L2 Distance 1-NN Distance Cluster Inertia Radius
+OpenThoughts3 0.8037 1.2656 0.0051 206,369.22 0.0172
+Nemotron-Post-Training-v1 0.8243 1.2827 0.1290 226,211.00 0.0174
+WebInstruct (Full) 0.7762 1.2436 0.1830 205,590.02 0.0169
+NaturalReasoning 0.8233 1.2818 0.1915 226,288.91 0.0173
+DLR-Web 0.8494 1.3026 0.3897 238,039.50 0.0177
+DLR-Book 0.8471 1.3008 0.3726 238,100.26 0.0176
+As detailed in Table 1, our datasets, DLR-Book and DLR-Web, consistently demonstrate a greater diversity of questions compared to the baseline datasets across all five semantic diversity metrics.
+The higher Mean Cosine Distance and Mean L2 Distance values confirm that our synthesized questions are, on average, more semantically distinct than those in the baseline datasets, indicating a broader conceptual scope.
+The most notable difference is observed in the 1-NN Distance, where our datasets score approximately twice as high as the baselines.
+This suggests our method generates far fewer semantically redundant questions.
+Furthermore, the Cluster Inertia and Radius scores for our datasets indicate that the generated questions occupy a larger and more varied volume within the semantic embedding space.
+This quantitative evidence confirms that our synthesis pipeline produces not only more complex and difficult questions but also a significantly more diverse set of questions.
+Additionally, we observe that questions synthesized from the web corpus (DLR-Web) exhibit greater diversity on most metrics than those synthesized from the book corpus (DLR-Book).
+4.3 Disciplinary Distribution
+We also used Qwen3-30B-A3B and the prompt in Figure 7 to assign disciplinary labels to the baseline datasets.
+We present a comparison of discipline distributions between our dataset and baseline datasets in Figure 4.
+For visualization, we highlight the most dominant disciplines within each dataset and those most representative of broader academic categories, while aggregating the remaining disciplines into the gray category “Other.” It is evident that many existing multidisciplinary datasets are heavily skewed toward a few disciplines, such as mathematics, leading to a highly imbalanced distribution across disciplines.
+Among them, only the Nemotron-Post-Training-v1 dataset exhibits a distribution comparable to ours, but its question difficulty and reasoning depth are substantially lower.
+The detailed number of questions per discipline in our dataset is provided in Table C.
+Figure 4: Discipline distribution across different datasets.
+5 Experiments
+5.1 Experimental Setup
+SFT Settings
+For all SFT experiments, we adhere to the hyperparameters detailed in Table 9.
+Evaluation Settings
+To ensure a fair comparison, we employ a zero-shot evaluation setting for all trained models, using the consistent generation configuration specified in Table 9.
+Depending on the characteristics of each benchmark, we perform $N$ independent sampling rollouts for each test instance.
+For $N=1$, we report accuracy.
+For $N>1$, we report mean accuracy (%) together with the standard deviation, and we also report the accuracy (%) under the Self-Consistency with Chain-of-Thought (CoT-SC) method, defined as the majority vote over the $N$ generated samples.
+Benchmarks
+We evaluate the model’s multidisciplinary reasoning capability using the most widely adopted benchmarks.
+The benchmarks, along with their respective disciplines and the number of rollouts ($N$), are detailed in Table 10.
+5.2 Supervised Fine-Tuning (SFT) Experiments
+We performed SFT on base models from the Qwen3 and Llama3 series using our synthetic datasets, DLR-Book and DLR-Web.
+The resulting models were subsequently compared against their official instruction-tuned counterparts under our evaluation framework.
+Table 2: SFT experiment results. The models are trained on DLR-Web, DLR-Book, or the combined DLR-Web + DLR-Book dataset.
+Model MMLU MMLU-Pro GPQA-Diamond GPQA-Main SuperGPQA
+Accuracy Accuracy Accuracy CoT-SC Accuracy CoT-SC Accuracy
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+As shown in Table 2, SFT with our synthetic datasets significantly improves model performance.
+The consistent improvements across diverse benchmarks demonstrate that our synthetic data method does not overfit to any specific domain or benchmark, but instead enhances the model’s general and robust reasoning capability.
+Notably, the multidisciplinary reasoning performance of the base models fine-tuned on our datasets even surpasses that of the official instruction-tuned models across all benchmarks.
+This improvement is particularly pronounced on highly complex reasoning tasks like GPQA-Diamond.
+These results affirm the efficacy of our data synthesis strategy for enhancing the multidisciplinary reasoning capabilities of language models.
+Table 3: Comparison with other baseline datasets on the Qwen3-8B-Base model.
+Dataset MMLU MMLU-Pro GPQA-Diamond GPQA-Diamond (CoT-SC) SuperGPQA
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+5.3 Comparison with Baseline Datasets
+We conducted a comparative analysis between our synthetically generated data and other prominent open-source synthetic datasets.
+To ensure a fair comparison, given the varying sizes of the original datasets, we randomly sampled an equal number of 304,181 instances from each dataset for the SFT experiments.
+For OpenThoughts3 and Nemotron-Post-Training-v1, the synthetic datasets already contain elaborate long CoT responses, so we retain their native versions.
+In contrast, the native responses in WebInstruct (Full) and NaturalReasoning are relatively simple.
+To enable a fairer comparison, we regenerate long CoT responses for them using the same model and settings as in our dataset, ensuring that performance differences arise solely from the questions.
+As shown in Table 3, the models trained on our datasets consistently outperform those trained on other baseline datasets across all benchmarks.
+Specifically, DLR-Book achieves the best performance on MMLU, MMLU-Pro, and GPQA-Diamond, while DLR-Web attains the highest score on SuperGPQA and ranks second on most other benchmarks.
+These results demonstrate the superior quality and effectiveness of our data synthesis approach compared to existing methods.
+5.4 Effect of Data Scaling
+To validate the scalability of our data synthesis methodology, we conducted experiments using the Qwen3-8B-Base model, training it on progressively larger subsets sampled from the DLR-Book dataset.
+The results, shown in Figure 6, demonstrate a consistent positive correlation between the scale of the synthetic data and model performance across all benchmarks.
+The observed scaling laws confirm that our method provides a reliable pathway to achieving superior model performance by synthesizing more data.
+Future work can leverage our proposed design logics to synthesize even larger datasets for continued improvement.
+Figure 5: Model performance as a function of data scale.
+Figure 6: The impact of source corpus quality on the synthesized data of our method.
+5.5 Effect of Source Corpus Quality
+It is widely acknowledged that book corpora are of higher quality than web corpora.
+To evaluate the impact of source quality on our data synthesis method, we conducted an SFT experiment on the Qwen3-8B-Base model, comparing data synthesized from the high-quality book corpus with that from the web corpus.
+To ensure a fair comparison, we matched the disciplinary distribution across both sources.
+For disciplines with limited FineFineWeb data, we used all available instances; otherwise, we applied random sampling.
+This procedure yielded two datasets of identical size (282,857 instances) and per-discipline distribution.
+As shown in Figure 6, the model fine-tuned on data from the book corpus consistently outperforms the one trained on data from the web corpus across most benchmarks, with the largest gains on complex reasoning tasks such as GPQA-Diamond and SuperGPQA.
+Nevertheless, the performance gap between the two corpora is modest, demonstrating our method’s robustness to source quality variations and its effectiveness at generating high-quality questions even from lower-quality corpora.
+5.6 Ablation Studies
+We conducted an ablation study on our question-synthesis pipeline using the Qwen3-8B-Base to assess the contributions of three key components: guidance from design logics, the coarse ranking, and the fine ranking of design logics.
+To ensure fair comparison, all ablation variants synthesized questions from the same book corpus by uniformly sampling 304,181 instances.
+Where applicable, we used identical retrieval with Qwen3-Embedding-4B and generated responses using the same model, Qwen3-235B-A22B-Thinking-2507-FP8.
+Specifically, we evaluate the following four settings:
+  • DESIGNER: This is our complete method, which includes coarse retrieval of the top-5 design logics by similarity, followed by LLM-based fine selection and synthesis based on “design logic”.
+  • w/o Design Logic : This setting replaces our use of explicit design logics.
+Instead of being guided by a logic, the LLM is prompted to generate a new question by imitating the style and structure of the most suitable exemplar.
+This exemplar is selected from a set of 5 relevant example questions that are retrieved from the question bank.
+  • w/o Coarse Ranking: This setting bypasses the semantic similarity-based retrieval of the top-5 relevant design logics.
+Instead, the LLM is prompted to select the most suitable logic from a set of 5 randomly chosen design logics.
+  • w/o Fine Ranking: This setting removes the LLM-based re-ranking stage. The single most similar design logic (the top-1 result from the coarse retrieval) is used directly to generate the question.
+Table 4 summarizes the results of our ablation studies.
+Our complete “DESIGNER” approach consistently outperforms ablated configurations across most benchmarks.
+The better performance of using design logics relative to directly using questions as examples confirmed our hypothesis that explicit logical structures are more accurate and robust guides for high-quality question synthesis.
+The removal of either coarse-grained retrieval (w/o Coarse Ranking) or fine-grained LLM re-ranking (w/o Fine Ranking) causes a noticeable performance drop, confirming the value of each stage in our two-step matching process.
+Table 4: Ablation study of our data synthesis pipeline.
+Method MMLU MMLU-Pro GPQA-Diamond GPQA-Diamond (CoT-SC) SuperGPQA
+[…]
+[…]
+[…]
+[…]
+6 Related Work
+Data Synthesis Paradigms
+Existing data synthesis methods can be broadly grouped into query-centric and doc-centric approaches.
+Query-centric approaches iteratively expand a seed query pool: Self-Instruct (Wang et al., 2023) samples from an initial pool to generate new QA pairs, while Wizard LM (Luo et al., 2024; Xu et al., 2023; Luo et al., 2025) and Auto Evol-Instruct (Zeng et al., 2024) evolve instructions for greater diversity and complexity.
+CoT-Self-Instruct (Yu et al., 2025) integrates reasoning into generation, Prismatic Synthesis (Jung et al., 2025) emphasizes gradient-level diversity, and SPARQ (Havrilla et al., 2025) evaluates difficulty and diversity, linking them to in- and out-of-distribution performance.
+However, these methods remain constrained by the coverage of the seed pool and the inherent model biases.
+Doc-centric approaches instead derive questions directly from raw corpora, ensuring stronger factual grounding: UltraChat (Ding et al., 2023) generates world-knowledge questions from sources like Wikidata, Humpback (Li et al., 2024) infers instructions backward from documents, MAmmoTH2 (Yue et al., 2024) mines web content via a Recall-Extract-Refine pipeline.
+NaturalReasoning (Yuan et al., 2025) focuses on generating high-difficulty reasoning questions, but its diversity is inherently limited by a single prompt template.
+Furthermore, several studies have explored generating new questions through external guidance, such as KPDDS (Huang et al., 2025), and the structurally guided approaches of Xu et al.
+(2025), Wang et al.
+(2025), and Bu et al.
+(2025).
+However, these approaches have two primary limitations compared to our method.
+First, they are query-centric methods that synthesize new questions from seed questions, which inherently constrains them to the scope of the initial seed pool and prevents generation from richer raw corpora.
+Second, their methodologies are tailored for disciplines with strong logical and structural reasoning, such as mathematics and coding, and cannot be readily extended to arbitrary multidisciplinary fields, such as the humanities, social sciences, and medicine.
+In contrast, our method synthesizes challenging and diverse questions by introducing multidisciplinary design logics to align with the corresponding documents.
+This approach is discipline-agnostic, making it applicable to any discipline.
+Reasoning Data Synthesis
+Another line of work leverages synthetic long CoT data to distill reasoning abilities into smaller models.
+DeepSeek (Guo et al., 2025) transfers reasoning skills from DeepSeek-R1 to open-source models, while OpenMathReasoning (Moshkov et al., 2025) augments math ability by generating long CoT traces for AoPS problems.
+OmniThought (Cai et al., 2025) collects multidisciplinary questions (math, code, science) to build a large CoT distillation dataset, and OpenThoughts (Guha et al., 2025) systematically explores recipes for generating long reasoning traces across these domains.
+While these approaches primarily focus on generating high-quality reasoning processes for existing questions, they lack the capability to produce original, diverse, and multidisciplinary questions, which is the central focus of our method.
+7 Conclusion
+In this paper, we introduced DESIGNER, a novel design-logic-guided data synthesis pipeline designed to address the scarcity of high-quality, multidisciplinary reasoning data for LLMs.
+Our core innovation is the concept of “Design Logic”, which abstracts the strategic process that human experts use to create challenging questions.
+By leveraging these logics, we generated two large-scale datasets, DLR-Book and DLR-Web, comprising millions of complex questions across 75 disciplines from diverse raw text sources.
+We validated the effectiveness of our approach through extensive experiments, demonstrating that models fine-tuned with our data achieve substantial performance gains in multidisciplinary reasoning.
+Ethics Statement
+All materials used in this study were subjected to filtering to minimize the inclusion of potentially harmful content.
+We encourage researchers who utilize our datasets to implement stricter filtering and processing measures to further ensure that the data conforms to the principles of ethical use.
+Reproducibility Statement
+We provide all complete and detailed data processing steps used for dataset synthesis in the main paper and the appendix.
+To facilitate reproducibility and independent verification, we will release a subset of our synthesized data that does not involve legal compliance concerns or potential conflicts of interest.
+The released data will include our synthesized questions, reference answers, long CoT responses, the corresponding source corpora and design logic, as well as associated labels (e.g., discipline, difficulty, type), enabling reproduction and validation of our method’s effectiveness and the quality of the synthesized data.
+The Use of Large Language Models (LLMs)
+We employed Large Language Models (LLMs) to assist us in polishing our paper and writing code.
+Appendix A Data Collection Details
 We curate three primary data sources and define a comprehensive discipline taxonomy:
@@
   • Question Bank: A proprietary repository of examination and practice items.
-Discipline Taxonomy: We have established a classification system comprising 75 distinct disciplines, as shown in LABEL:tab:discipline_stats_en.
-This taxonomy provides comprehensive coverage across several major areas:
+Discipline Taxonomy: We have established a disciplinary classification system comprising 75 distinct disciplines, as shown in Table C.
+This discipline taxonomy provides comprehensive coverage across several major areas:
   • STEM: Science, Technology, Engineering, and Mathematics.
@@
   • Arts.
-3.2 Data Processing and Filtering
-3.2.1 Question Bank Processing
-Multi-dimensional Labeling
-Using Qwen3-30B-A3B (non-thinking), we annotate more than 150 million questions from the proprietary question bank with discipline, difficulty, and type labels.
-The specific prompts for these classifications are detailed in Figure 3, Figure 4, and Figure 5, respectively.
-Clustering-Based Question Selection
-To facilitate the extraction of design logic, we intend to filter a high-quality and diverse subset from the question bank.
-This subset is specifically curated to ensure a broad coverage of multiple disciplines and varying levels of difficulty.
-We compute embeddings with Qwen3-Embedding-4B and apply K-means clustering within each discipline (Ahmed et al., 2020).
-We then equally sample from each cluster with a difficulty ratio of Very Hard:Hard:Medium = 3:2:1, ensuring both diversity and sufficient difficulty.
-For each discipline, we determine the number of clusters via silhouette coefficient search and set per-discipline sample sizes according to the question bank’s discipline distribution.
-This process resulted in a curated set of 132,409 questions for design-logic extraction.
-3.2.2 Book Corpus Processing
-Text Preprocessing
-We process each book at the chapter level. Chapters that exceed 5,000 words are further split into blocks up to 5,000 words. We then perform MinHash deduplication to obtain the final book segments.
-Discipline Labeling and Quality Assessment
-We assign discipline labels to book segments using a ModernBERT-large–based model that has been fine-tuned for this specific classification task (Warner et al., 2024).
-Subsequently, we introduce two key metrics for quality assessment: readability and helpfulness.
-Readability is predicted by a BERT-based classifier (Turc et al., 2019), which is designed to filter out incoherent or disorganized text.
-Helpfulness is scored by the fineweb-edu-classifier (Lozhkov et al., 2024), which provides a score on a scale of 0 to 5, allowing us to quantitatively assess the educational value of each document.
-Quality-Prioritized Sampling
-A total of over three million book segments were sampled from the corpus.
-For each discipline, we allocate quotas proportional to its frequency in both the book corpus and the question bank.
-To ensure the acquisition of high-quality data, we devised a quality-prioritized sampling policy.
-We first removed all samples for which the readability was classified as negative.
-Based on the helpfulness scores, the remaining candidate segments within each discipline are sorted in descending order, and samples are then selected from the top of this list until the predetermined quota is met.
-This process ensures that the sampled texts are of the highest quality available, with the vast majority of the final selected segments exhibiting a helpfulness score of $\geq 2$.
-3.2.3 Web Corpus Processing
-Reasoning-Oriented Filtering
-Because not every FineFineWeb text contains sufficient reasoning content, we design a five-level scoring rubric (scores starting at 0; prompt in the Figure 6) to assess a text’s potential usefulness for studying reasoning processes.
-We utilized the Qwen3-30B-A3B (non-thinking) model to score 6.5 billion texts from the FineFineWeb dataset, retaining only those with a score of $\geq 3$.
-Discipline Relabeling
-Finally, we used the Qwen3-30B-A3B (non-thinking) model and the prompt specified in Figure 3 to relabel the retained texts, aligning the web corpus with our 75-discipline taxonomy.
-4 Synthesizing Questions using Design Logic
-4.1 Design Logic Extraction
-Human educators typically construct exam questions by following a structured chain of design decisions that transforms simple knowledge concepts into complex and challenging questions rather than merely enumerating facts.
-A representative design process might involve: identifying the target objectives and knowledge points to be tested $\rightarrow$ constructing a contextual scenario $\rightarrow$ designing the logical reasoning process $\rightarrow$ formulating the correct answer $\rightarrow$ incorporating question traps and distracting information $\rightarrow$ validating the final question.
-Consequently, a solver must deconstruct the question through multi-step reasoning, which goes far beyond simple memorization (e.g., explaining a knowledge point).
-Inspired by this practice, we propose a method for question synthesis based on “design logic”.
-Leveraging the prompt presented in Figure 7, we instruct an LLM (DeepSeek-R1-0528) to analyze the sampled authentic questions from our question bank.
-For each question, the LLM is tasked with three key steps: (i) deduce the question designer’s thought process, (ii) analyze how the question was constructed from the relevant knowledge points, and (iii) abstract and summarize the underlying design logic and principles.
-We specifically require the model to organize the resulting abstract design logic into the Mermaid format.
-This procedure yields a reusable pool of design logics, which the model can leverage these diverse logics to generate new questions based on provided source materials.
-This approach eliminates the need for handcrafting a large set of disparate prompts, significantly increasing automation.
-It also substantially improves the difficulty and diversity of LLM-generated questions while aligning their structure more closely with questions authored by human teachers.
-4.2 Design Logic Deduplication
-To enhance the diversity of design principles provided to the LLMs, we further deduplicate the extracted design logic.
-This ensures a varied set of principles, which in turn facilitates the generation of diverse questions.
-Specifically, we use the Qwen3-Embedding-4B model to map each design logic into an embedding vector.
-For each discipline, we compute pairwise semantic similarities between all design logics to obtain a similarity matrix $S\in\mathbb{R}^{n\times n}$, where $S_{ij}$ denotes the semantic similarity between the $i$-th and $j$-th design logic.
+Appendix B Design Logic Deduplication Algorithm
 Algorithm 1 Graph-based Deduplication via Centroid Selection
@@
 25:return $\mathcal{R}$.
-We then apply graph-based connected-components deduplication.
-We model the $n$ design logics within a discipline as a graph $G=(V,E)$ with $|V|=n$.
-An undirected edge is added between nodes $i$ and $j$ if their semantic similarity $S_{ij}\geq\tau$.
-This construction forms connected components, where each component represents a group of mutually redundant items.
-From each connected component $C$, we select a unique representative — the item with the maximum sum of similarities to all other items in the component, and discard the rest.
-The complete procedure is detailed in Algorithm 1, with the similarity threshold set to $\tau=0.85$.
-After deduplication, we retain a total of 125,328 unique design logics. The per-discipline counts are reported in Table LABEL:tab:discipline_stats_en.
-4.3 Question Synthesis
-To address the computational complexity of directly matching each design logic with every text segment, which would lead to a combinatorial explosion, we propose an approach similar to Retrieval-Augmented Generation (RAG).
-We compute the cosine similarity between the embeddings of each text segment and each design logic for each discipline-specific corpus (book or web) using the Qwen3-Embedding-4B model with a task-specific instruction (Figure 8).
-Specifically, for a text segment $t$ and a design logic $d$, we calculate $s(t,d)\,=\,\cos\!\bigl{(}\mathbf{e}(t),\,\mathbf{e}(d)\bigr{)}$.
-We then retain the top-5 design logics with the highest similarity scores as candidates for question synthesis for each text segment.
-We then prompt an LLM (DeepSeek-R1-0528) to generate a challenging, graduate-level exam question (prompt detailed in Figure 9).
-Specifically, the LLM is instructed to (i) select the most suitable question-design logic from the top-5 candidates for the given source text, and (ii) strictly follow the corresponding logic and steps to construct a challenging question.
-This two-stage process acts as a coarse-to-fine ranking mechanism.
-The initial retrieval of the top-5 design logics using semantic similarity serves as a coarse-grained recall stage.
-The LLM’s subsequent selection of the most appropriate logic then functions as a fine-grained ranking stage.
-This strategy ensures a precise match between the design logic and the text, significantly enhancing the quality of the synthesized questions.
+Appendix C Design Logic and Question Quantity
 […]
-Question Deduplication and Decontamination
-We implemented a two-stage filtering pipeline to manage redundancy and avoid evaluation leakage.
-First, we used MinHash-based detection to remove highly similar questions.
-Second, to address potential data contamination, we decontaminated our synthetic questions by comparing them against all evaluation benchmarks employed in this study.
-This was achieved through the standard 13-gram decontamination method, during which non-semantic punctuation was ignored.
-Any question flagged as similar is discarded.
-Using the filtered book and web corpora from Section 3.2, we generate one corresponding reasoning question per text segment with DeepSeek-R1-0528.
-After deduplication and decontamination, our final reasoning dataset comprises 3,040,620 challenging questions synthesized from the book corpus and 1,658,541 challenging questions synthesized from the web corpus.
-4.4 Response Synthesis
-To demonstrate that our synthesized questions can effectively elicit and transfer the long CoT capabilities of a reasoning model and improve the performance of models trained on this data, we employ Qwen3-235B-A22B-Thinking-2507-FP8 to generate a corresponding long CoT response for each synthesized question.
-These question-response pairs are then used for supervised fine-tuning (SFT).
-5 Data Analysis
-To begin our quality assessment, we performed a quantitative analysis comparing the complexity and difficulty of our synthesized datasets (DLR-Book and DLR-Web) with the baseline datasets.
-This analysis focused on the distribution of question types, the lengths of questions and responses, the overall difficulty of the questions, and their diversity.
-5.1 Question Type Analysis
-As shown in Table 1, problem-solving questions form the majority of our synthesized datasets, constituting 64.92% of the book corpus and 63.91% of the web corpus.
-Multiple-choice questions are the next most common type, comprising 29.94% and 32.43%, respectively.
-This distribution analysis indicates that the composition of our synthesized questions is skewed towards reasoning-based types, as opposed to simple recall-oriented ones.
-Table 1: Distribution of question types in our synthesized datasets.
-Type DLR-Book DLR-Web
-Problem-solving question 64.92% 63.91%
-Multiple-choice question 29.94% 32.43%
-Proof question 4.39% 2.67%
-Other question types 0.75% 0.99%
-5.2 Data Complexity
-In Table 2, we present a comparison of the average lengths of questions and their corresponding responses for our two synthesized datasets against WebInstruct (Full) and NaturalReasoning.
-The responses were all generated by the Qwen3-235B-A22B-Thinking-2507-FP8 model.
-It is evident that the questions generated by our method, as well as the responses required to answer them, are significantly longer than those in WebInstruct (Full) (Yue et al., 2024) and NaturalReasoning (Yuan et al., 2025).
-This observation indirectly substantiates the high degree of difficulty and complexity inherent in our synthesized questions.
-This substantial increase in question length indicates that our synthesis method generates questions with more detailed context and more intricate premises, requiring a deeper understanding of the input.
-We likewise observe longer response lengths in our datasets, a direct consequence of the increased question complexity.
-This observation suggests that solving our synthesized questions requires a deeper chain of reasoning compared to the baselines.
-Table 2: Average length (in characters) of questions and corresponding responses.
-The responses were all generated by the Qwen3-235B-A22B-Thinking-2507-FP8 model.
-The average lengths for WebInstruct (Full) and NaturalReasoning were calculated from a random sample of 304,181 instances.
-Dataset Avg. Question Length Avg. Response Length
-WebInstruct (Full) 180.43 11133.88
-NaturalReasoning 332.08 17162.85
-DLR-Book 1284.79 18090.06
-DLR-Web 1205.96 17528.59
-5.3 Difficulty Analysis
-To provide a more direct measure of difficulty, we used the Qwen3-30B-A3B-Instruct-2507 model to assign difficulty labels to questions from our book and web corpora, as well as from WebInstruct (Full) and NaturalReasoning.
-Based on the complexity and length of the required reasoning, we categorized questions into four difficulty levels: Easy, Medium, Hard, and Very Hard.
-As shown in Table 3, our datasets are significantly more difficult.
-The combined percentage of “Hard” and “Very Hard” questions is notably high: 89.84% for DLR-Book and 83.15% for DLR-Web.
-This contrasts sharply with the baseline datasets.
-DLR-Book contains 54.66% “Very Hard” questions, substantially more than the 31.11% in NaturalReasoning and 3.59% in WebInstruct (Full).
-Conversely, the proportion of “Easy” questions in our datasets is negligible (0.27% and 0.72%), especially when compared to WebInstruct (Full)’s 39.02%.
-Collectively, these quantitative analyses of question length and difficulty distribution consistently demonstrate that our data synthesis method produces questions of significantly greater complexity and difficulty than the baseline datasets.
-Table 3: Difficulty distribution of questions across datasets.
-Dataset Easy Medium Hard Very Hard
-WebInstruct (Full) 39.02% 39.58% 17.84% 3.59%
-NaturalReasoning 2.10% 26.25% 40.54% 31.11%
-DLR-Book 0.27% 9.88% 35.18% 54.66%
-DLR-Web 0.72% 16.13% 36.24% 46.91%
-5.4 Diversity Analysis
-To evaluate the semantic diversity of the synthesized questions, we conducted a quantitative analysis using distance-based metrics.
-This approach assesses diversity by examining the distribution of question embeddings in a high-dimensional semantic space.
-Diversity Metrics and Experimental Setup
-We first generated high-dimensional vector representations for each question using the Qwen3-Embedding-4B model.
-For a given question set $X=\{x_{1},x_{2},\dots,x_{N}\}$, this process yields a corresponding set of embedding vectors $E=\{e_{1},e_{2},\dots,e_{N}\}$, where $e_{i}\in\mathbb{R}^{d}$ is the $d$-dimensional embedding for question $x_{i}$.
-We uniformly sample N = 300,000 questions from each dataset and compute the following five distance-based metrics in the embedding space.
-  1. 1.
-    Mean Cosine Distance (DistMean Cosine) (Yang et al., 2025b): The average cosine distance between all unique pairs of embeddings, calculated as $M_{\text{cosine}}=\frac{2}{N(N-1)}\sum_{i<j}(1-\cos(e_{i},e_{j}))$.
-A higher value indicates greater semantic dissimilarity.
-  2. 2.
-    Mean L2 Distance (DistMean L2) (Yang et al., 2025b): The average Euclidean distance between all unique pairs of embeddings, calculated as $M_{\text{L2}}=\frac{2}{N(N-1)}\sum_{i<j}\|e_{i}-e_{j}\|_{2}$.
-This measures the average separation of questions in the embedding space.
-  3. 3.
-    1-Nearest Neighbor Distance (1-NN Distance) (Stasaski & Hearst, 2022): The average cosine distance from each embedding to its single nearest neighbor, given by $M_{\text{1-NN}}=\frac{1}{N}\sum_{i}d_{1}(e_{i})$, where $d_{1}(e_{i})$ is the cosine distance from $e_{i}$ to its closest neighbor.
-This metric highlights the presence of tightly clustered, near-identical questions.
-  4. 4.
-    Cluster Inertia (Du & Black, 2019): The total squared distance of samples to their closest cluster center after applying the K-means clustering algorithm.
-It is calculated as $M_{\text{inertia}}=\sum_{i=1}^{N}\min_{j}\|e_{i}-c_{j}\|_{2}^{2}$, where $c_{j}$ are the cluster centroids.
-This measures the overall spread and density of the data clusters.
-  5. 5.
-    Radius (Lai et al., 2020): The geometric mean of the standard deviations of the embedding dimensions, modeling the data as a multi-dimensional Gaussian distribution: $M_{\text{Radius}}=(\prod_{j=1}^{d}\sigma_{j})^{1/d}$, where $\sigma_{j}$ is the standard deviation along the $j$-th dimension.
-It directly quantifies the spread of the data in the semantic space.
-As detailed in Table 4, our datasets, DLR-Book and DLR-Web, consistently demonstrate a greater diversity of questions compared to the baseline datasets across all five semantic diversity metrics.
-The higher Mean Cosine Distance and Mean L2 Distance values confirm that our synthesized questions are, on average, more semantically distinct than those in WebInstruct (Full) and NaturalReasoning, indicating a broader conceptual scope.
-The most notable difference is observed in the 1-NN Distance, where our datasets score approximately twice as high as the baselines.
-This suggests our method generates far fewer semantically redundant questions.
-Furthermore, the Cluster Inertia and Radius scores for our datasets indicate that the generated questions occupy a larger and more varied volume within the semantic embedding space.
-This quantitative evidence confirms that our synthesis pipeline produces not only more complex and difficult questions but also a significantly more diverse set of questions.
-Table 4: Semantic diversity metrics for our synthesized datasets and the baseline datasets, calculated on a uniform sample of 300,000 questions from each. Higher is better for all metrics shown.
-Dataset DistMean Cosine DistMean L2 1-NN Distance Cluster Inertia Radius
-WebInstruct (Full) 0.7762 1.2436 0.1830 205,590.02 0.0169
-NaturalReasoning 0.8233 1.2818 0.1915 226,288.91 0.0173
-DLR-Book 0.8471 1.3008 0.3726 238,100.2656 0.0176
-DLR-Web 0.8494 1.3026 0.3897 238,039.5000 0.0177
-6 Experiments
-In this study, we evaluate the effectiveness of our synthesized data by conducting supervised fine-tuning (SFT) on base models.
-The resulting models are then assessed on a diverse range of widely-used benchmarks.
-Our extensive experiments demonstrate that the data synthesized by our method can effectively elicit the long CoT reasoning capabilities of LLMs for complex, multidisciplinary questions.
-Furthermore, with the same amount of training data, our method yields superior model performance compared to existing datasets like NaturalReasoning and WebInstruct (Full).
-6.1 Experimental Setup
-SFT Settings
-In the majority of our experiments, we perform SFT on the Qwen3-8B-Base model.
-Subsequently, we also validate the effectiveness of our synthesized data across other model scales.
-For all SFT experiments, we adhere to the hyperparameters detailed in Table 5.
-Table 5: Hyperparameters for Supervised Fine-Tuning (SFT).
-Parameter Value
-Epoch 6
-Batch Size 64
-Learning Rate 1e-5
-Learning Rate Schedule Cosine decay to 0
-Evaluation Settings
-To ensure a fair comparison, we employ a zero-shot evaluation setting for all trained models, using the consistent generation configuration specified in Table 6.
-For each benchmark, we conduct $N$ independent sampling rollouts for each test instance.
-Our evaluation involves the following three metrics:
-  • Pass@1: The mean accuracy (%) across the $N$ rollouts. We report both the Pass@1 score and its standard deviation to show the performance and stability of our method.
-  • CoT-SC: The accuracy (%) determined by a majority vote over the $N$ generated samples.
-This metric is equivalent to Self-Consistency with Chain-of-Thought (CoT-SC) method and is reported for all cases where $N>1$.
-  • Pass@N: The proportion (%) of questions for which at least one correct solution is found among $N$ rollouts.
-Table 6: Generation configuration for evaluation.
-Parameter Value
-Temperature 0.6
-Top-K 20
-Top-P 0.95
-Max Context Length 32768
-Benchmarks. The benchmarks used in our evaluation, along with their respective disciplines and the number of rollouts ($N$), are detailed in Table 7.
-Table 7: Evaluation benchmarks, their disciplines, and the number of rollouts (NN).
-Benchmark Disciplines Rollout ($N$)
-AIME 2024 Mathematics 64
-AIME 2025 Mathematics 64
-MATH-500 (Lightman et al., 2024) Mathematics 4
-GPQA-Diamond (Rein et al., 2024) Physics, Chemistry, Biology 10
-GPQA-Main (Rein et al., 2024) Physics, Chemistry, Biology 10
-SuperGPQA (Du et al., 2025) 285 graduate-level disciplines across 13 fields (e.g., Engineering, Management, Economics, Education, History, Science, Medicine, Law, Sociology, Philosophy, Agriculture, Literature) 1
-MMLU (Hendrycks et al., 2020) 57 disciplines including STEM, humanities, social sciences, and professional fields 1
-MMLU-Pro (Wang et al., 2024) Math, Physics, Chemistry, Law, Engineering, Economics, Health, Psychology, Business, Biology, Philosophy, Computer Science, History, Other 1
-6.2 Supervised Fine-Tuning Experiments
-We performed supervised fine-tuning (SFT) on the Qwen3-4B-Base [^3] https://huggingface.co/Qwen/Qwen3-4B-Base and Qwen3-8B-Base [^4] https://huggingface.co/Qwen/Qwen3-8B-Base models using our synthetic datasets, DLR-Book and DLR-Web.
-The resulting SFT models were then compared against the official Qwen3-4B [^5] https://huggingface.co/Qwen/Qwen3-4B and Qwen3-8B [^6] https://huggingface.co/Qwen/Qwen3-8B models (thinking mode) under our evaluation settings.
-Table 8: Performance of models fine-tuned on our datasets.
-The models are trained on DLR-Web, DLR-Book, or the combined DLR-Web+Book dataset.
-Baselines include the official Qwen3-4B and Qwen3-8B models (thinking mode), evaluated under our settings.
-The best and second-best scores in each column are shown in bold and underlined, respectively.
-Model MMLU MMLU-Pro GPQA-Diamond GPQA-Main SuperGPQA
-Pass@1 Pass@1 Pass@1 CoT-SC Pass@1 CoT-SC Pass@1
+Number of Design Logics, Book Questions, and Web Questions by Discipline
+Discipline Design Logic Book Question Web Question
+[…]
+Discipline Design Logic Book Question Web Question
 […]
 […]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-As shown in Table 8, supervised fine-tuning with our synthetic datasets significantly improves model performance.
-For both the 4B and 8B model scales, our fine-tuned models consistently outperform the official Qwen3-4B and Qwen3-8B (thinking mode) baselines across all reasoning benchmarks.
-Notably, models trained on the combined “DLR-Web+Book” dataset achieve the best results across all benchmarks, while those trained on the “DLR-Book” dataset achieve the second-best.
-This improvement is particularly pronounced on highly complex reasoning tasks like GPQA-Diamond.
-These results affirm the efficacy of our data synthesis strategy for enhancing the reasoning capabilities of language models.
-6.3 Performance on Advanced Mathematical Reasoning
-For challenging mathematical competition questions, such as those from AIME, we additionally report Pass@N results.
-Prior work has shown that methods like Reinforcement Learning with Verifiable Rewards (RLVR) primarily enhance Pass@1 accuracy by better selecting among pre-existing reasoning paths, without significantly improving Pass@N rates (Yue et al., 2025).
-While our models’ comparatively lower Pass@1 scores are expected due to our exclusive use of SFT without a subsequent RL phase, the significant improvement in Pass@N for our fine-tuned models is noteworthy.
-This result suggests that our data successfully introduces novel reasoning patterns, thereby expanding the model’s solution space and increasing its problem-solving potential.
-This expanded capability provides a strong foundation for future work to apply RL and convert these reasoning skills into higher Pass@1 performance.
-Table 9: Performance on mathematical reasoning benchmarks, including Pass@N results.
-The best results in each column are in bold, and the second-best are underlined.
-The models trained on DLR-Web+Book show marked improvements in Pass@N, indicating an expanded reasoning capability.
-Model MATH-500 AIME 2024 AIME 2025
-Pass@1 CoT-SC Pass@N Pass@1 CoT-SC Pass@N Pass@1 CoT-SC Pass@N
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-6.4 Data Scaling Effects
-To validate the scalability of our data synthesis methodology, we conducted a series of experiments to evaluate the reasoning capabilities of models trained on progressively larger datasets synthesized by our approach.
-The results, illustrated in Figure 2, demonstrate a clear and positive correlation between the volume of synthetic data and model performance across multiple challenging reasoning benchmarks.
-The consistent improvement across this diverse set of benchmarks validates that our synthesis process is not overfit to a specific domain or benchmark but rather enhances a general, robust reasoning capability.
-These strong scaling laws confirm that our method provides a reliable pathway to achieving superior model performance through data augmentation, and future work can leverage our proposed design logics to synthesize even larger datasets for continued improvement.
-Figure 2: The model’s performance on four key benchmarks (AIME 2025, MMLU-Pro, GPQA-Diamond, and SuperGPQA) as a function of the training data quantity.
-The performance metric is Pass@N for AIME 2025, while all others are Pass@1.
-6.5 Comparison with Baseline Datasets
-We conducted a comparative analysis between our synthetically generated data and other prominent open-source synthetic datasets.
-To ensure a fair comparison, given the varying sizes of the original datasets, we randomly sampled an equal number of 304,181 instances from each dataset for the fine-tuning experiments.
-For the purpose of equitable comparison, we regenerated long CoT responses for all baseline methods (including WebInstruct (Full) and NaturalReasoning) using the same Qwen3-235B-A22B-Thinking-2507-FP8 model.
-As shown in Table 10, our Design-Logic-Reasoning datasets consistently outperform WebInstruct (Full) and NaturalReasoning across all benchmarks.
-Specifically, DLR-Book achieves the best results on MMLU, MMLU-Pro, and GPQA-Diamond, while DLR-Web attains the highest performance on SuperGPQA and the second-best on other benchmarks.
-These results validate the superior quality and effectiveness of our data generation approach compared to existing methods.
-Table 10: Comparison with other doc-centric synthetic datasets on the Qwen3-8B-Base model.
-All experiments were conducted using a subsample of 304,181 instances from each dataset.
-The best results in each column are in bold, and the second-best are underlined.
-Dataset MMLU MMLU-Pro GPQA-Diamond SuperGPQA
-Pass@1 Pass@1 Pass@1 CoT-SC Pass@1
-[…]
-[…]
-[…]
-[…]
-6.6 Ablation Studies
-We conducted an ablation study on our question-synthesis pipeline to quantify the contributions of two key components: coarse-to-fine matching between design logics and corpus segments, and the use of explicit design logics.
-To ensure a fair comparison, all ablations synthesize questions from the same book corpus by uniformly sampling 304,181 text segments.
-Where applicable, we used identical retrieval with Qwen3-Embedding-4B and generated model responses with the same model, Qwen3-235B-A22B-Thinking-2507-FP8.
-  • w/o Coarse Ranking: This setting bypasses the semantic similarity-based retrieval of the top-5 relevant design logics.
-Instead, the LLM is prompted to select the most suitable logic from a set of 5 randomly chosen logics.
-  • w/o Fine Ranking: This setting removes the LLM-based re-ranking stage. The single most similar design logic (the top-1 result from the coarse retrieval) is used directly to generate the question.
-  • w/ Example Questions: This setting replaces our use of abstract design logics.
-Instead of being guided by a logic, the LLM is prompted to generate a new question by imitating the style and structure of the most suitable exemplar.
-This exemplar is selected from a set of 5 relevant example questions that are retrieved from the question bank.
-  • DESIGNER: This is our complete method, which includes coarse retrieval of the top-5 design logics by similarity, followed by LLM-based fine selection and synthesis based on “design logic”.
-Table 11 summarizes the results of our ablation studies.
-Our complete “DESIGNER” approach consistently outperforms ablated configurations across most benchmarks.
-The removal of either coarse-grained retrieval (w/o Coarse Ranking) or fine-grained LLM re-ranking (w/o Fine Ranking) causes a noticeable performance drop, confirming the value of each stage in our two-step matching process.
-The “w/o Fine Ranking” configuration, in particular, performs well on the MATH benchmark, suggesting the initial retrieval is highly effective at identifying relevant logics.
-However, the consistent improvements across other benchmarks validate the full coarse-to-fine pipeline’s benefit for handling complex, multidisciplinary reasoning.
-Finally, the superior performance of design logics over concrete “Example Questions” confirms our hypothesis that explicit logical structures are more accurate and robust guides for high-quality question synthesis.
-Table 11: Ablation study of our data synthesis pipeline.
-All experiments were conducted on the Qwen3-8B-Base model using data synthesized from the same book corpus.
-The best results in each column are in bold, and the second-best are underlined.
-Method MATH-500 MMLU GPQA-Diamond SuperGPQA
-Pass@1 CoT-SC Pass@1 Pass@1 CoT-SC Pass@1
-[…]
-[…]
-[…]
-[…]
-6.7 Effect of Source Corpus Quality on Synthesized Data
-Table 12: Effect of source-corpus quality on SFT outcomes with equal per-discipline and total size.
-All experiments were conducted on the Qwen3-8B-Base model.
-The best results in each column are in bold.
-Source Corpus MATH-500 MMLU MMLU-Pro GPQA-Diamond SuperGPQA
-Pass@1 CoT-SC Pass@1 Pass@1 Pass@1 CoT-SC Pass@1
-[…]
-[…]
-It is widely accepted that book corpora are of higher quality than web corpora.
-To test how source corpus quality affects our data synthesis method, we conducted a controlled SFT experiment.
-We compared data synthesized from a high-quality book corpus to that from a web corpus.
-For a fair comparison, we ensured the disciplinary data distribution was identical across both datasets.
-For disciplines with limited data in the FineFineWeb corpus, we used all available instances.
-For other disciplines, we used random sampling.
-This process resulted in two final datasets of equal size (282,857 instances each) and identical per-discipline distribution.
-As shown in Table 12, the results validate our hypothesis.
-The model fine-tuned on data from the book corpus consistently outperforms the one trained on data from the web corpus across almost all benchmarks.
-This performance gain is most notable in complex reasoning tasks like GPQA-Diamond and SuperGPQA, confirming that our synthesis method benefits from higher-quality source material.
-However, our method also proves robust to variations in source quality, as the performance gap between the two corpora is not substantial.
-This indicates that our approach effectively synthesizes high-quality questions to improve model performance, regardless of the initial corpus quality.
-7 Related Work
-Query-Centric Data Synthesis
-The core idea of query-centric data synthesis is to generate new training data by iteratively expanding or evolving an existing query pool.
-Self-Instruct (Wang et al., 2023) samples questions from an existing query pool and leverages an LLM to generate new question-answer pairs, which are then used to fine-tune the model.
-Wizard LM (Xu et al., 2023) (Luo et al., 2025) (Luo et al., 2024) and Auto Evol-Instruct (Zeng et al., 2024) employ instruction evolution to produce more complex and diverse data.
-CoT-Self-Instruct (Yu et al., 2025) integrates the CoT reasoning mechanism into the instruction generation process, enhancing the quality of synthetic data through a “think-first, then-generate” approach.
-Prismatic Synthesis (Jung et al., 2025) pushes the definition of data diversity from superficial textual features to a more fundamental level of model behavior, aiming to maximize the diversity of model-induced gradients.
-The SPARQ method (Havrilla et al., 2025) introduces a framework for evaluating diversity and difficulty in data synthesis, arguing that high-difficulty data improves in-distribution performance while diversity enhances out-of-distribution generalization.
-Although these query-centric methods attempt to optimize the difficulty and diversity of synthetic data, they are still limited by the initial query pool and the distribution biases of the model used for synthesis, making it difficult to cover questions across various disciplines.
-Doc-Centric Data Synthesis
-Doc-centric data synthesis methods start with a large amount of unstructured (e.g., CC, books) or structured (e.g., knowledge graphs) textual data.
-They generate question-answer pairs by directly extracting or inferring them from documents, thereby ensuring the synthetic data is closely tied to specific domain knowledge and facts.
-UltraChat  (Ding et al., 2023) generates questions about world knowledge from knowledge sources like Wikidata and meta topics pre-defined by human experts (e.g., technology, health).
-Humpback (Li et al., 2024) also uses a large web corpus as a data source, training a “backward model” to infer an instruction that would lead to a given document.
-The KPDDS method (Huang et al., 2025) constructs a Topic-level Co-occurrence Probability Matrix (TCPM) based on extracted topics and knowledge points, guiding data synthesis by sampling from this matrix.
-MAmmoTH2 (Yue et al., 2024) mines high-quality questions directly from existing web content using a three-step “Recall-Extract-Refine” pipeline, but the difficulty of its extracted questions is low.
-NaturalReasoning (Yuan et al., 2025) also synthesizes data from a large-scale pre-training corpus, but its process is more focused on generating high-difficulty reasoning questions, though the diversity of its synthetic questions is limited by the single prompt.
-Our work uniquely introduces the concept of design logic, analogizing the doc-centric data synthesis process to a human teacher creating questions.
-By matching documents with appropriate design logic, we significantly increase the difficulty and diversity of the synthetic questions.
-Reasoning Data Synthesis
-Using synthetic long CoT data for data distillation is an effective way to improve the reasoning abilities of smaller models.
-DeepSeek (Guo et al., 2025) distills a series of smaller open-source models using high-quality data samples generated by DeepSeek-R1, significantly enhancing the reasoning capabilities of these smaller models.
-OpenMathReasoning (Moshkov et al., 2025) extracts high-quality math problems from an online math forum (AoPS) and leverages powerful existing models to generate long CoT responses for these problems, significantly improving the models’ mathematical abilities.
-OmniThought (Cai et al., 2025) collects questions from the fields of mathematics, code, and science, and generates a long CoT dataset based on multiple teacher models.
-However, these methods primarily focus on generating high-quality reasoning processes for existing questions, and the open-source community still lacks high-quality, multidisciplinary original questions.
-Our method, in contrast, focuses on leveraging naturally available multidisciplinary documents to synthesize difficult and diverse multidisciplinary questions by introducing design logic.
-8 Conclusion
-This paper introduces DESIGNER, a novel design-logic-guided data synthesis pipeline designed to address the scarcity of high-quality, multidisciplinary reasoning data for LLMs.
-Our core innovation is the concept of “Design Logic”, which abstracts the strategic process human experts use to create challenging questions.
-By leveraging these logics, we generated two large-scale datasets, DLR-Book and DLR-Web, comprising millions of complex questions across 75 disciplines from diverse raw text sources.
-Our synthesized questions are shown to be significantly more difficult and diverse than those from existing methods, moving beyond simple factual recall to require deep, multi-step reasoning.
-We validated the effectiveness of our approach through extensive experiments, demonstrating that models fine-tuned with our data achieve substantial performance gains in multidisciplinary reasoning.
-Specifically, models trained on our datasets not only outperform those fine-tuned with existing multidisciplinary datasets of the same volume but can also surpass the multidisciplinary reasoning performance of the official Qwen3 models of the same size, validating the effectiveness of our method.
-This work provides a scalable paradigm for creating diverse and challenging reasoning data, paving a new path for advancing LLMs’ reasoning capabilities beyond domain-specific tasks.
-Appendix A Design Logic and Question Quantity
-Table 13: Number of Design Logics, Book Questions, and Web Questions by Discipline
-Discipline Design Logic Book Question Web Question
-Aerospace Science and Technology 980 15025 1815
 Agricultural Engineering 966 14950 11312
@@
 World History 987 39703 5503
-Table 13: (Continued) Number of Design Logics, Book Questions, and Web Questions by Discipline
-Appendix B Prompts
+Appendix D Baseline Datasets
+Table 5: The baseline datasets utilized in our experiments. For the Nemotron-Post-Training-v1 dataset, only the data from the math and STEM categories was used.
+Dataset URL
+OpenThoughts3
+(Guha et al., 2025)
+https://huggingface.co/datasets/open-thoughts/OpenThoughts3-1.2M
+Nemotron-Post-Training-v1
+(Bercovich et al., 2025)
+https://huggingface.co/datasets/nvidia/Nemotron-Post-Training-Dataset-v1
+WebInstruct (Full)
+(Yue et al., 2024)
+https://huggingface.co/datasets/TIGER-Lab/WebInstructFull
+NaturalReasoning
+(Yuan et al., 2025)
+https://huggingface.co/datasets/facebook/natural_reasoning
+Appendix E Question Type Analysis
+As shown in Table 6, problem-solving questions form the majority of our synthesized datasets, constituting 64.92% of the book corpus and 63.91% of the web corpus.
+Multiple-choice questions are the next most common type, comprising 29.94% and 32.43%, respectively.
+This distribution analysis indicates that the composition of our synthesized questions is skewed towards reasoning-based types, as opposed to simple recall-oriented ones.
+Table 6: Distribution of question types in our synthesized datasets.
+Type DLR-Book DLR-Web
+Problem-solving question 64.92% 63.91%
+Multiple-choice question 29.94% 32.43%
+Proof question 4.39% 2.67%
+Other question types 0.75% 0.99%
+Appendix F Detailed Difficulty Analysis
+Table 7: Difficulty distributions of questions across different datasets and benchmarks.
+Dataset Easy Medium Hard Very Hard
+GSM8K 78.09% 21.91% 0.0% 0.0%
+MMLU 60.51% 34.09% 5.29% 0.11%
+MMLU-Pro 34.48% 44.04% 20.86% 0.62%
+GPQA-Diamond 0.51% 19.7% 71.72% 8.08%
+GPQA-Main 0.22% 26.56% 67.41% 5.8%
+SuperGPQA 35.97% 32.18% 29.88% 1.98%
+OpenThoughts3 1.09% 19.32% 70.51% 9.07%
+Nemotron-Post-Training-v1 32.27% 48.61% 16.09% 3.04%
+WebInstruct (Full) 39.02% 39.58% 17.84% 3.57%
+NaturalReasoning 2.10% 26.25% 40.54% 31.11%
+DLR-Web 0.72% 16.13% 36.24% 46.91%
+DLR-Book 0.27% 9.88% 35.18% 54.66%
+Appendix G Detailed Diversity Metrics
+We first generated high-dimensional vector representations for each question using the Qwen3-Embedding-4B model.
+For a given question set $X=\{x_{1},x_{2},\dots,x_{N}\}$, this process yields a corresponding set of embedding vectors $E=\{e_{1},e_{2},\dots,e_{N}\}$, where $e_{i}\in\mathbb{R}^{d}$ is the $d$-dimensional embedding for question $x_{i}$.
+We uniformly sample N = 300,000 questions from each dataset and compute the following five distance-based metrics in the embedding space.
+  1. 1.
+    Mean Cosine Distance: The average cosine distance between all unique pairs of embeddings, calculated as $M_{\text{cosine}}=\frac{2}{N(N-1)}\sum_{i<j}(1-\cos(e_{i},e_{j}))$.
+A higher value indicates greater semantic dissimilarity.
+  2. 2.
+    Mean L2 Distance: The average Euclidean distance between all unique pairs of embeddings, calculated as $M_{\text{L2}}=\frac{2}{N(N-1)}\sum_{i<j}\|e_{i}-e_{j}\|_{2}$.
+This measures the average separation of questions in the embedding space.
+  3. 3.
+    1-Nearest Neighbor Distance (1-NN Distance): The average cosine distance from each embedding to its single nearest neighbor, given by $M_{\text{1-NN}}=\frac{1}{N}\sum_{i}d_{1}(e_{i})$, where $d_{1}(e_{i})$ is the cosine distance from $e_{i}$ to its closest neighbor.
+This metric highlights the presence of tightly clustered, near-identical questions.
+  4. 4.
+    Cluster Inertia: The total squared distance of samples to their closest cluster center after applying the K-means clustering algorithm.
+It is calculated as $M_{\text{inertia}}=\sum_{i=1}^{N}\min_{j}\|e_{i}-c_{j}\|_{2}^{2}$, where $c_{j}$ are the cluster centroids.
+This measures the overall spread and density of the data clusters.
+  5. 5.
+    Radius: The geometric mean of the standard deviations of the embedding dimensions, modeling the data as a multi-dimensional Gaussian distribution: $M_{\text{Radius}}=(\prod_{j=1}^{d}\sigma_{j})^{1/d}$, where $\sigma_{j}$ is the standard deviation along the $j$-th dimension.
+It directly quantifies the spread of the data in the semantic space.
+Appendix H Experimental Details
+Table 8: Hyperparameters for Supervised Fine-Tuning (SFT).
+Parameter Value
+Epoch 6
+Batch Size 64
+Learning Rate 1e-5
+Learning Rate Schedule Cosine decay to 0
+Table 9: Generation configuration for evaluation.
+Parameter Value
+Temperature 0.6
+Top-K 20
+Top-P 0.95
+Max Context Length 32768
+Table 10: Evaluation benchmarks, their disciplines, and the number of rollouts (NN).
+Benchmark Disciplines Rollout ($N$)
+MMLU (Hendrycks et al., 2020) 57 disciplines including STEM, humanities, social sciences, and professional fields 1
+MMLU-Pro (Wang et al., 2024b) Math, Physics, Chemistry, Law, Engineering, Economics, Health, Psychology, Business, Biology, Philosophy, Computer Science, History, Other 1
+GPQA-Diamond (Rein et al., 2024) Physics, Chemistry, Biology 10
+GPQA-Main (Rein et al., 2024) Physics, Chemistry, Biology 10
+SuperGPQA (Du et al., 2025) 285 graduate-level disciplines across 13 fields (e.g., Engineering, Management, Economics, Education, History, Science, Medicine, Law, Sociology, Philosophy, Agriculture, Literature) 1
+Appendix I Prompts
 Prompt for Discipline Classification You are a professional multidisciplinary data labeling expert specializing in the classification of multidisciplinary academic questions.
@@
 The placeholders {logic1} through {logic5} and {text} are replaced with specific design logics and the source text, respectively.
+Appendix J Design Logic Examples
+In this section, we provide design logic examples for several representative disciplines, such as Computer Science and Technology, Clinical Medicine, Mathematics, Law, Psychology, and Archaeology.
+Each example includes the source code in Mermaid format and the corresponding visual flowchart.
+1graph TD
+2 A["Problem Domain"] --> B["Sorted Arrays & Median"]
+3 B --> C["Key Insight: Partitioning Without Merging"]
+4 C --> D["Algorithm Selection: Binary Search"]
+5 D --> E["Optimization: O(log(min(m, n)))"]
+6 E --> F["Edge Cases"]
+7 F --> F1["Empty Arrays"]
+8 F --> F2["Full-Array Partition"]
+9 F --> F3["Even/Odd Length Handling"]
+10 A --> G["Constraints Design"]
+11 G --> G1["Large Input Sizes: 10^6"]
+12 G --> G2["Time Complexity: O(min(log m, log n))"]
+13 G --> G3["Auxiliary Space: O((m+n)/2) hint for naive approach"]
+14 A --> H["Test Cases"]
+15 H --> H1["Example 1: Odd Length"]
+16 H --> H2["Example 2: Even Length"]
+17 H --> H3["Edge: One Array Empty"]
+18 H --> H4["Edge: Both Empty"]
+19 H --> H5["One Array All Smaller"]
+(a) The design logic formatted as Mermaid source code.
+(b) The flowchart generated from the Mermaid code.
+Figure 14: An example of the design logic for a Computer Science and Technology problem, showing the Mermaid source code (a) and the corresponding visual flowchart (b).
+1graph TD
+2 A[Select Core Concept] --> B[Identify Pathognomonic Features];
+3 B --> C[Layer Multidisciplinary Clues];
+4 C --> D[Clinical History<br><i>e.g., chronicity, symptoms</i>];
+5 C --> E[Imaging/Lab Findings<br><i>e.g., specific radiology signs</i>];
+6 C --> F[Pathology/Cytology<br><i>e.g., cell type, differentiation</i>];
+7 D --> G[Require Synthesis of All Clues];
+8 E --> G;
+9 F --> G;
+10 G --> H[Demand Etiologic Deduction<br><i>e.g., primary site, mechanism</i>];
+11 H --> I[Incorporate Strategic Distractors<br><i>e.g., common misdiagnoses</i>];
+12 I --> J[Ensure Clues Refute Distractors<br><i>e.g., atypia level rules out alternatives</i>];
+13 J --> K[Align Timeline with Disease Biology<br><i>e.g., indolent vs. acute</i>];
+(a) The design logic formatted as Mermaid source code.
+(b) The flowchart generated from the Mermaid code.
+Figure 15: An example of the design logic for a Clinical Medicine problem, showing the Mermaid source code (a) and the corresponding visual flowchart (b).
+1graph TD
+2 A[Define Problem Type] --> B[Select Core Concept: <br>e.g., Exponential Diophantine, Polynomial, Inequality]
+3 B --> C[Choose Parameters with Constraints]
+4 C --> C1[Base/Numbers: <br> - Coprime or low common factors<br> - Strategic modular properties<br> - Growth disparities]
+5 C --> C2[Variables: <br> - Multiple exponents/terms<br> - Interdependent constraints]
+6 C --> C3[Ensure solution space control: <br> - Small solution exists<br> - Bounded by properties/growth]
+7 A --> D[Incorporate Reasoning Layers]
+8 D --> D1[Step 1: Derive necessary conditions<br> e.g., modular arithmetic, symmetry, calculus]
+9 D --> D2[Step 2: Impose non-trivial constraints<br> e.g., parity, divisibility, inequalities]
+10 D --> D3[Step 3: Force case analysis<br> e.g., small-value testing, critical point checks]
+11 A --> E[Finalize for Challenge]
+12 E --> E1[Verify: Solution requires all steps<br> - No brute-force accessibility<br> - Key insight is necessary]
+13 E --> E2[Refine: Remove redundancies<br> - Ensure clarity and conciseness]
+14 E --> E3[Generalize: Test with variants<br> e.g., change bases/operators]
+(a) The design logic formatted as Mermaid source code.
+(b) The flowchart generated from the Mermaid code.
+Figure 16: An example of the design logic for a Mathematics problem, showing the Mermaid source code (a) and the corresponding visual flowchart (b).
+1graph TD
+2 A[Start: Define Core Task] --> B[Select Key Items for Comparison];
+3 B --> C[Set Task: Rank or Compare Items];
+4 C --> D[Require Detailed Explanation of Each Item];
+5 D --> E[Include Opposing Views/Controversies];
+6 C --> F[Discuss Broader Impacts e.g., Historical, Cultural, Political];
+7 F --> G[Specify Factors to Consider e.g., Rights, Institutions, Ethics];
+8 C --> H[Link to Modern Implications/Controversies];
+9 C --> I[Demand Use of Examples and Evidence];
+10 I --> J[Support Arguments with Facts and Data];
+11 J --> K[Ensure Complex Logical Reasoning];
+12 K --> L[End: Assess Higher-Order Thinking];
+(a) The design logic formatted as Mermaid source code.
+(b) The flowchart generated from the Mermaid code.
+Figure 17: An example of the design logic for a Law problem, showing the Mermaid source code (a) and the corresponding visual flowchart (b).
+1graph TD
+2 A[Primary Goal] --> B[Explore Complex Human Behavior]
+3 B --> C[Ethical Decision-Making Under Constraints]
+4 A --> D[Ensure Pedagogical Safety]
+5
+6 C --> E[Design Imperatives]
+7 E --> E1[Require Controversial Perspective Coverage]
+8 E --> E2[Explicitly Forbid Harm Advocacy]
+9 E --> E3[Embed Iterative Refinement Mechanism]
+10
+11 D --> F[Psychological Safeguards]
+12 F --> F1[Clinical Role Assignment<br>e.g., Psychologist/Researcher]
+13 F --> F2[Tone Mandate<br>e.g., Caring/Supportive Language]
+14 F --> F3[Scenario Distancing<br>e.g., Hypothetical Survival Context]
+15
+16 G[Structural Elements] --> G1[High-Stakes Scenario]
+17 G1 --> G1a[Reveal True Behavioral Drivers]
+18 G --> G2[Concrete Decision Task<br>e.g., Resource Prioritization]
+19 G2 --> G2a[Create Measurable Choice Points]
+20
+21 H[Quality Control] --> H1[Anti-Repetition Protocol]
+22 H --> H2[Nuance Requirement<br>e.g., Multiple Revision Layers]
+23
+24 A --> G
+25 A --> H
+(a) The design logic formatted as Mermaid source code.
+(b) The flowchart generated from the Mermaid code.
+Figure 18: An example of the design logic for a Psychology problem, showing the Mermaid source code (a) and the corresponding visual flowchart (b).
+1flowchart TD
+2A[Core Objective] --> B["Require Synthesis of Multiple Skills"]
+3A --> C["Demand Critical Interpretation"]
+4A --> D["Connect Abstract Concepts to Concrete Evidence"]
+5
+6B --> B1["Textual/Visual Analysis"]
+7B --> B2["Contextual Analysis"]
+[…]
+9
+10C --> C1["Evaluate Symbolic Nuances"]
+11C --> C2["Assess Cultural Significance"]
+12C --> C3["Challenge Surface-Level Readings"]
+13
+14D --> D1["Link Physical Artifacts to Belief Systems"]
+15D --> D2["Trace Practices to Worldviews"]
+16D --> D3["Reconstruct Civilizational Logic"]
+(a) The design logic formatted as Mermaid source code.
+(b) The flowchart generated from the Mermaid code.
+Figure 19: An example of the design logic for an Archaeology problem, showing the Mermaid source code (a) and the corresponding visual flowchart (b).

### v3

@@
 The central insight of our approach is the notion of “Design Logic”, which encapsulates how human experts transform knowledge into complex, reasoning-intensive exam questions.
-We observed that when human education experts design challenging and insightful questions, they don’t merely state facts.
+We observed that when human education experts design challenging and insightful questions, they do not merely state facts.
 Instead, they follow a structured design process, as illustrated in Figure 1.
@@
 Specifically, our pipeline is illustrated in Figure 2.
-First, we process large-scale book and web corpora with multi-dimensional labeling and filtering (discipline, readability, knowledge value) to construct a high-quality source material library.
+First, we process large-scale book and web corpora with multi-dimensional labeling and filtering (discipline, readability, educational value, reasoning depth) to construct a high-quality source material library.
 From a question bank of hundreds of millions, we cluster and sample a diverse set of difficult questions, from which an LLM reverse-engineers and abstracts over 120K structured “design logics” to construct a reusable design logic library.
@@
 Using this pipeline, we constructed two large-scale reasoning datasets: DLR-Book (3.04 million questions from the book corpus) and DLR-Web (1.66 million questions from the web corpus).
-These datasets span 75 disciplines, including STEM, humanities, social sciences, applied and professional fields, and arts, extending beyond common disciplines.
+These datasets span 75 disciplines, including STEM, humanities, social sciences, applied and professional fields, and the arts, extending beyond common disciplines.
   • Inspired by the question-creation process of human teachers, we propose a question synthesis method guided by “Design Logic” and construct a corresponding library of design logics.
@@
 From each cluster, we draw an equal number of questions following a fixed difficulty ratio of 3:2:1 (Very Hard:Hard:Medium), and align per-discipline sizes with the overall bank distribution.
+If higher-difficulty questions are insufficient, they are backfilled with lower-difficulty ones to ensure the per-discipline totals.
 This process resulted in a curated set of 132,409 questions for design-logic extraction.
@@
 Readability is predicted with a BERT-based model (Turc et al., 2019) to filter incoherent or disorganized text, while helpfulness (0–5) is scored by the fineweb-edu-classifier (Lozhkov et al., 2024) to quantify educational value.
-Based on quotas proportional to discipline frequencies in both the book corpus and the question bank, segments with negative readability are removed, remaining candidates are ranked by helpfulness, and top items are selected until quotas are met.
+Based on quotas proportional to discipline frequencies in both the book corpus and the question bank, segments with negative readability are removed.
+The remaining candidates are ranked by helpfulness, and the top items are selected until the quotas are met.
 This procedure yields three million high-quality segments, with most scoring helpfulness $\geq 2$.
@@
 Human educators design exam questions through structured steps that transform knowledge points into complex challenges, rather than simple fact recall.
-A typical process involves identifying objectives, constructing contexts, designing reasoning paths, formulating answers, adding distractors, and validating the question.
+A typical process involves identifying objectives, constructing contexts, designing reasoning paths, formulating answers, adding distractors, and validating the questions.
 Solvers must engage in multi-step reasoning beyond memorization.
@@
 Using the curated book and web corpora (Section 2), DeepSeek-R1-0528 generates one reasoning question per text segment.
-After filtering, the final dataset contains 3,040,620 questions from books and 1,658,541 from web sources.
+After filtering, the final dataset Design-Logic-Reasoning-Book (DLR-Book) comprises 3,040,620 questions from the book corpus, and Design-Logic-Reasoning-Web (DLR-Web) comprises 1,658,541 questions from the web corpus.
 3.4 Response Synthesis
@@
 We provide all complete and detailed data processing steps used for dataset synthesis in the main paper and the appendix.
-To facilitate reproducibility and independent verification, we will release a subset of our synthesized data that does not involve legal compliance concerns or potential conflicts of interest.
-The released data will include our synthesized questions, reference answers, long CoT responses, the corresponding source corpora and design logic, as well as associated labels (e.g., discipline, difficulty, type), enabling reproduction and validation of our method’s effectiveness and the quality of the synthesized data.
+To facilitate reproducibility and independent verification, we have released a subset of our synthesized data that does not involve legal compliance concerns or potential conflicts of interest.
+The released data includes our synthesized questions, reference answers, long CoT responses, the corresponding source corpora and design logic, as well as associated labels (e.g., discipline, difficulty, type), enabling reproduction and validation of our method’s effectiveness and the quality of the synthesized data.
 The Use of Large Language Models (LLMs)

### v5

@@
 Large language models (LLMs) have achieved remarkable success in many natural language tasks but still struggle with complex, multi-step reasoning, particularly across diverse disciplines.
-Existing reasoning datasets often lack disciplinary breadth, reasoning depth, and diversity, and lack guiding principles for question synthesis.
+Existing reasoning datasets often lack disciplinary breadth, reasoning depth, and diversity, as well as guiding principles for question synthesis.
 We propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages naturally available, extensive raw documents (e.g., book corpus and web corpus) to generate multidisciplinary challenging questions.
-We introduce the concept of “design logic” and instruct LLMs to mimic human educators’ question-creation process, enabling automated synthesis of large-scale, high-difficulty questions.
+We introduce the concept of “design logic” and instruct LLMs to mimic human educators’ question-creation process, enabling the automated synthesis of large-scale, high-difficulty questions.
 We use LLMs to reverse-engineer and abstract over 120,000 design logics from existing questions across various disciplines.
-By matching these design logics with source documents, we are able to create reasoning questions that far surpass the difficulty and diversity of existing datasets.
+By matching these design logics with source documents, we are able to generate reasoning questions with controllable question types and difficulty levels.
 Using this pipeline, we synthesized two large-scale reasoning datasets that span 75 disciplines: DLR-Book (3.04 million questions from the book corpus) and DLR-Web (1.66 million questions from the web corpus).
@@
 Our data substantially enhances their multidisciplinary reasoning capabilities, outperforming existing datasets.
-Notably, after SFT on our datasets, the base versions of these models even surpass their official instruction-tuned counterparts.
+Notably, by applying SFT on the base versions of these models using only our data, we even surpass their official final models that have undergone the full post-training process.
 [^1] Project page: https://attention-is-all-i-need.github.io/Design-Logic-Reasoning
@@
 1 Introduction
-Large language models (LLMs) have demonstrated exceptional capabilities in various natural reasoning tasks (Chowdhery et al., 2023; OpenAI, 2023), such as mathematics and coding (Trinh et al., 2024; Zhu et al., 2024), especially when utilizing long chain-of-thought (CoT) techniques (Wei et al., 2022; Jaech et al., 2024; Guo et al., 2025).
-However, they still lag behind human experts on university-level, discipline-specific reasoning (Phan et al., 2025), largely due to the scarcity of large-scale, high-quality, and diverse training data.
-Existing datasets focus mainly on math and programming, drawing on competition platforms rich in open-ended questions (Moshkov et al., 2025; Cai et al., 2025), while most other disciplines lack comparable resources.
+Large language models (LLMs) have demonstrated exceptional capabilities in various natural reasoning tasks (chowdhery2023palm; openai2023gpt4), such as mathematics and coding (trinh2024solving; zhu2024deepseek), especially when utilizing long chain-of-thought (CoT) techniques (wei2022chain; jaech2024openai; guo2025deepseek).
+However, they still lag behind human experts in university-level, discipline-specific reasoning (phan2025humanity), largely due to the scarcity of large-scale, high-quality, and diverse training data.
+Existing datasets focus mainly on math and programming, drawing on competition platforms rich in open-ended questions (moshkov2025aimo; cai2025reasoning), while most other disciplines lack comparable resources.
 This scarcity limits the development of LLMs’ multidisciplinary reasoning capabilities.
-Data synthesis with LLMs is an effective solution to data scarcity (Wang et al., 2024a).
+Data synthesis with LLMs is an effective solution to data scarcity (wang2024survey).
 Existing question synthesis methods fall into two categories: query-centric and doc-centric.
-Query-centric approaches expand seed questions through rewriting, added constraints (e.g., Evol-Instruct (Xu et al., 2023)), or incorporating chain-of-thought reasoning (Wang et al., 2023; Yu et al., 2025), but they are limited by the coverage of the seed pool and inherent model biases.
-Doc-centric methods instead generate questions from unstructured (e.g., web, books) or structured (e.g., knowledge graphs) sources (Yue et al., 2024; Yuan et al., 2025; Huang et al., 2025), ensuring broad disciplinary coverage grounded in authentic knowledge.
+Query-centric approaches expand seed questions through rewriting, added constraints (e.g., Evol-Instruct (xu2023wizardlm)), or incorporating chain-of-thought reasoning (wang2023self; yu2025cot), but they are limited by the coverage of the seed pool and inherent model biases.
+Doc-centric methods instead generate questions from unstructured (e.g., web, books) or structured (e.g., knowledge graphs) source documents (yue2024mammoth2; yuan2025naturalreasoning; huang2025key), ensuring broad disciplinary coverage grounded in authentic knowledge.
 However, they struggle to control difficulty and diversity, often degenerating into factual recall.
@@
 Instead, they follow a structured design process, as illustrated in Figure 1.
-This design logic itself is a valuable and transferable form of “meta-knowledge”.
+The design logic is a form of reusable meta-knowledge that abstracts the underlying reasoning structure, enabling LLMs to generate new questions with the same complex reasoning patterns from entirely different source texts.
 Specifically, our pipeline is illustrated in Figure 2.
@@
 These datasets span 75 disciplines, including STEM, humanities, social sciences, applied and professional fields, and the arts, extending beyond common disciplines.
-  • Inspired by the question-creation process of human teachers, we propose a question synthesis method guided by “Design Logic” and construct a corresponding library of design logics.
+  • By reverse-engineering the meta-knowledge of human educators, we propose a fundamentally new question synthesis method guided by “Design Logic”.
+This approach enables the generation of truly complex, multi-step reasoning questions from raw text by providing the structured, reusable, and abstract control over difficulty and diversity that prior doc-centric methods lacked.
 Our data analysis indicates that the questions synthesized by our method exhibit greater difficulty and diversity compared to those in the baseline datasets.
-  • We validate the effectiveness of our synthesized data through comprehensive comparative and ablation experiments on the Qwen3 (Yang et al., 2025a) and Llama3 (Grattafiori et al., 2024) model families.
-The results demonstrate that the data synthesized by our method significantly enhances the multidisciplinary reasoning capabilities of LLMs.
+  • We validate the effectiveness of our synthesized data through comprehensive comparative and ablation experiments on the Qwen3 (yang2025qwen3) and Llama3 (dubey2024llama) model families.
+The results demonstrate that the data synthesized by our method significantly enhance the multidisciplinary reasoning capabilities of LLMs.
 2 Data Curation
@@
 2.1 Question Bank Processing
-We annotate over 150 million questions from the proprietary bank with discipline, difficulty, and type using Qwen3-30B-A3B (non-thinking mode), following prompts in Figure 7, Figure 8, and Figure 9, respectively.
-To obtain a high-quality and diverse subset, we compute embeddings with Qwen3-Embedding-4B and apply K-means clustering within each discipline (Ahmed et al., 2020), with cluster numbers determined by silhouette search.
+We annotate over 150 million questions from the proprietary question bank with discipline, difficulty, and question types using Qwen3-30B-A3B (non-thinking mode), following the prompts in Figure 7, Figure 8, and Figure 9, respectively.
+To obtain a high-quality and diverse subset, we compute embeddings with Qwen3-Embedding-4B and apply K-means clustering within each discipline (ahmed2020k), with cluster numbers determined by silhouette search.
 From each cluster, we draw an equal number of questions following a fixed difficulty ratio of 3:2:1 (Very Hard:Hard:Medium), and align per-discipline sizes with the overall bank distribution.
@@
 This process resulted in a curated set of 132,409 questions for design-logic extraction.
+Any question bank, regardless of its initial quality, can be processed using this filtering pipeline to obtain a high-quality and diverse subset for extracting design logics.
+Our method is therefore applicable to any question banks and can be directly applied to publicly available datasets.
 2.2 Book Corpus Processing
 This corpus is processed at the chapter level, with chapters over 5,000 words split into smaller blocks and deduplicated via MinHash.
-Discipline labels are assigned by a ModernBERT-large classifier fine-tuned for disciplinary classification (Warner et al., 2024).
-Readability is predicted with a BERT-based model (Turc et al., 2019) to filter incoherent or disorganized text, while helpfulness (0–5) is scored by the fineweb-edu-classifier (Lozhkov et al., 2024) to quantify educational value.
+Discipline labels are assigned by a ModernBERT-large classifier fine-tuned for disciplinary classification (warner2024smarter).
+Readability is predicted with a BERT-based model (turc2019well) to filter incoherent or disorganized text, while helpfulness (0–5) is scored by the fineweb-edu-classifier (lozhkov2024fineweb-edu) to quantify educational value.
 Based on quotas proportional to discipline frequencies in both the book corpus and the question bank, segments with negative readability are removed.
@@
 2.3 Web Corpus Processing
-We apply reasoning-oriented filtering and discipline relabeling to the FineFineWeb corpus, scoring 6.5 billion texts with Qwen3-30B-A3B (non-thinking mode) using a five-level rubric (prompt in Figure 10) and retaining those with scores $\geq$3.
+We apply reasoning-oriented filtering and discipline relabeling to the FineFineWeb corpus, scoring 6.5 billion texts with Qwen3-30B-A3B (non-thinking mode) using a five-level rubric (prompt in Figure 10) and retaining those with scores $\geq 3$.
 The selected texts are then relabeled with the same model (prompt in Figure 7) to align with our 75-discipline taxonomy.
@@
 With $\tau=0.85$, this graph-based deduplication procedure (Algorithm 1) yields 125,328 unique design logics, with per-discipline counts in Table C.
-We provide several examples of our synthesized design logic in Appendix J.
+We provide several examples of our synthesized design logic in Appendix K.
 3.3 Question Synthesis
 To avoid the combinatorial explosion resulting from exhaustively matching design logics with text segments, we adopt a retrieval-augmented approach.
-For each discipline-specific corpus, we compute cosine similarity between embeddings of a text segment $t$ and a design logic $d$ using Qwen3-Embedding-4B with task-specific instructions (Figure 12): $s(t,d)\,=\,\cos\!\bigl(\mathbf{e}(t),\,\mathbf{e}(d)\bigr)$.
+For each discipline-specific corpus, we compute the cosine similarity between embeddings of a text segment $t$ and a design logic $d$ using Qwen3-Embedding-4B with task-specific instructions (Figure 12): $s(t,d)\,=\,\cos\!\bigl(\mathbf{e}(t),\,\mathbf{e}(d)\bigr)$.
 The top-5 logics with the highest similarity are retained as candidates.
@@
 We employ a two-stage filtering pipeline: (i) MinHash-based deduplication to remove near-duplicates, and (ii) 13-gram decontamination against all evaluation benchmarks to prevent leakage.
-Using the curated book and web corpora (Section 2), DeepSeek-R1-0528 generates one reasoning question per text segment.
+Using the curated book and web corpora, DeepSeek-R1-0528 generates one reasoning question per text segment.
 After filtering, the final dataset Design-Logic-Reasoning-Book (DLR-Book) comprises 3,040,620 questions from the book corpus, and Design-Logic-Reasoning-Web (DLR-Web) comprises 1,658,541 questions from the web corpus.
@@
 To quantify dataset diversity, we generate high-dimensional vector representations for 300,000 uniformly sampled questions from each dataset using the Qwen3-Embedding-4B model.
-We then compute the following five distance-based diversity metrics in the embedding space: Mean Cosine Distance (Yang et al., 2025b), Mean L2 Distance (Yang et al., 2025b), 1-Nearest Neighbor (1-NN) Distance (Stasaski & Hearst, 2022), Cluster Inertia (Du & Black, 2019), and Radius (Lai et al., 2020).
+We then compute the following five distance-based diversity metrics in the embedding space: Mean Cosine Distance (yang2025measuring), Mean L2 Distance (yang2025measuring), 1-Nearest Neighbor (1-NN) Distance (stasaski-hearst-2022-semantic), Cluster Inertia (du-black-2019-boosting), and Radius (lai-etal-2020-diversity).
 Detailed definitions and formulas for these metrics are provided in Appendix G.
@@
 We present a comparison of discipline distributions between our dataset and baseline datasets in Figure 4.
-For visualization, we highlight the most dominant disciplines within each dataset and those most representative of broader academic categories, while aggregating the remaining disciplines into the gray category “Other.” It is evident that many existing multidisciplinary datasets are heavily skewed toward a few disciplines, such as mathematics, leading to a highly imbalanced distribution across disciplines.
+For visualization, we highlight the ten most dominant disciplines within each dataset and those most representative of broader academic categories, while aggregating the remaining disciplines into the gray category “Other.” It is evident that many existing multidisciplinary datasets are heavily skewed toward a few disciplines, such as mathematics, leading to a highly imbalanced distribution across disciplines.
 Among them, only the Nemotron-Post-Training-v1 dataset exhibits a distribution comparable to ours, but its question difficulty and reasoning depth are substantially lower.
@@
 Figure 4: Discipline distribution across different datasets.
+4.4 Data Quality and Label Accuracy Validation
+To assess the reliability of our method, we employed GPT-5.1 to conduct a multi-dimensional evaluation on 10,000 randomly sampled instances.
+We utilized the prompts detailed in Figure 14, Figure 15, and Figure 16 to assess content quality, and further employed the prompts in Figure 17, Figure 18, and Figure 19 to validate the appropriateness of the assigned discipline, difficulty, and question type labels, respectively.
+Our analysis demonstrates high data integrity: 96.70% of the synthesized questions were verified as complete and answerable, providing sufficient conditions to derive solutions.
+Furthermore, 84.69% of the questions exhibited strict alignment with their corresponding design logics.
+The evaluation of labels also showed high accuracy, with 90.14%, 96.53%, and 91.22% of the instances confirmed to have appropriate discipline, difficulty, and question type labels, respectively.
+Finally, we evaluated the consistency between the reference answers generated by DeepSeek-R1-0528 and the long Chain-of-Thought (CoT) responses generated by Qwen3-235B-A22B-Thinking-2507-FP8, achieving a 71.48% agreement rate.
+A significant portion of our synthesized data comprises complex, open-ended reasoning questions.
+In disciplines such as Humanities, Social Sciences, Arts, and Applied Professional Fields, questions inherently lack a single fixed answer.
+Consequently, this result is reasonable given the inherent diversity of valid reasoning paths in open-ended multidisciplinary questions.
 5 Experiments
@@
 We performed SFT on base models from the Qwen3 and Llama3 series using our synthetic datasets, DLR-Book and DLR-Web.
-The resulting models were subsequently compared against their official instruction-tuned counterparts under our evaluation framework.
+The resulting models were subsequently compared against their official final models that have undergone the full Post-training process under our evaluation framework.
 Table 2: SFT experiment results. The models are trained on DLR-Web, DLR-Book, or the combined DLR-Web + DLR-Book dataset.
@@
 The consistent improvements across diverse benchmarks demonstrate that our synthetic data method does not overfit to any specific domain or benchmark, but instead enhances the model’s general and robust reasoning capability.
-Notably, the multidisciplinary reasoning performance of the base models fine-tuned on our datasets even surpasses that of the official instruction-tuned models across all benchmarks.
+Notably, the multidisciplinary reasoning performance of the base models fine-tuned on our datasets even surpasses that of their official final models that have undergone the full post-training process across all benchmarks.
 This improvement is particularly pronounced on highly complex reasoning tasks like GPQA-Diamond.
-These results affirm the efficacy of our data synthesis strategy for enhancing the multidisciplinary reasoning capabilities of language models.
+These results affirm the efficacy of our data synthesis strategy for enhancing the multidisciplinary reasoning capabilities of LLMs.
+In addition, we provide the SFT experiment results on GSM8K and MATH-500 in Appendix I.
 Table 3: Comparison with other baseline datasets on the Qwen3-8B-Base model.
@@
 Future work can leverage our proposed design logics to synthesize even larger datasets for continued improvement.
-Figure 5: Model performance as a function of data scale.
-Figure 6: The impact of source corpus quality on the synthesized data of our method.
+Figure 5: Performance scaling with synthesized data size. Benchmark accuracy increases steadily with data scale.
+Figure 6: Performance comparison of models trained on data synthesized from lower-quality web corpora versus higher-quality book corpora.
 5.5 Effect of Source Corpus Quality
-It is widely acknowledged that book corpora are of higher quality than web corpora.
+It is widely acknowledged that book corpora are of higher quality than web corpora (brown2020language; gao2020pile; rae2021scaling).
 To evaluate the impact of source quality on our data synthesis method, we conducted an SFT experiment on the Qwen3-8B-Base model, comparing data synthesized from the high-quality book corpus with that from the web corpus.
@@
 This procedure yielded two datasets of identical size (282,857 instances) and per-discipline distribution.
-As shown in Figure 6, the model fine-tuned on data from the book corpus consistently outperforms the one trained on data from the web corpus across most benchmarks, with the largest gains on complex reasoning tasks such as GPQA-Diamond and SuperGPQA.
-Nevertheless, the performance gap between the two corpora is modest, demonstrating our method’s robustness to source quality variations and its effectiveness at generating high-quality questions even from lower-quality corpora.
+As shown in Figure 6, the model fine-tuned on data from the book corpus outperforms the one trained on data from the web corpus across most benchmarks, with the largest gains on complex reasoning tasks such as GPQA-Diamond and SuperGPQA.
+Nevertheless, the performance gap induced by the two source corpora of differing quality is negligible, which demonstrates that our method is robust to variations in source quality and can reliably generate high-quality questions even when applied to lower-quality corpora.
 5.6 Ablation Studies
@@
   • w/o Fine Ranking: This setting removes the LLM-based re-ranking stage. The single most similar design logic (the top-1 result from the coarse retrieval) is used directly to generate the question.
-Table 4 summarizes the results of our ablation studies.
-Our complete “DESIGNER” approach consistently outperforms ablated configurations across most benchmarks.
-The better performance of using design logics relative to directly using questions as examples confirmed our hypothesis that explicit logical structures are more accurate and robust guides for high-quality question synthesis.
-The removal of either coarse-grained retrieval (w/o Coarse Ranking) or fine-grained LLM re-ranking (w/o Fine Ranking) causes a noticeable performance drop, confirming the value of each stage in our two-step matching process.
 Table 4: Ablation study of our data synthesis pipeline.
@@
 […]
+[…]
+[…]
+Table 4 summarizes the results of our ablation studies.
+Our complete “DESIGNER” approach consistently outperforms ablated configurations across most benchmarks.
+The better performance of using design logics relative to directly using questions as examples confirmed our hypothesis that explicit logical structures are more accurate and robust guides for high-quality question synthesis.
+The removal of either coarse-grained retrieval (w/o Coarse Ranking) or fine-grained LLM re-ranking (w/o Fine Ranking) causes a performance drop, confirming the value of each stage in our two-step matching process.
+We further evaluated the “w/o Design Logic” variant on the full book corpus (3.04 million samples).
+The performance degradation of “w/o Design Logic (Full)” is notably significant.
+Specifically, on the GPQA-Diamond benchmark, the performance gap between “w/o Design Logic (Full)” and the model employing “Design Logic” reaches five percentage points, providing stronger evidence for the advantage of employing design logic.
 6 Related Work
@@
 Existing data synthesis methods can be broadly grouped into query-centric and doc-centric approaches.
-Query-centric approaches iteratively expand a seed query pool: Self-Instruct (Wang et al., 2023) samples from an initial pool to generate new QA pairs, while Wizard LM (Luo et al., 2024; Xu et al., 2023; Luo et al., 2025) and Auto Evol-Instruct (Zeng et al., 2024) evolve instructions for greater diversity and complexity.
-CoT-Self-Instruct (Yu et al., 2025) integrates reasoning into generation, Prismatic Synthesis (Jung et al., 2025) emphasizes gradient-level diversity, and SPARQ (Havrilla et al., 2025) evaluates difficulty and diversity, linking them to in- and out-of-distribution performance.
+Query-centric approaches iteratively expand a seed query pool: Self-Instruct (wang2023self) samples from an initial pool to generate new QA pairs, while Wizard LM (luowizardcoder; xu2023wizardlm; luowizardmath) and Auto Evol-Instruct (zeng2024automatic) evolve instructions for greater diversity and complexity.
+CoT-Self-Instruct (yu2025cot) integrates reasoning into generation, Prismatic Synthesis (jung2025prismatic) emphasizes gradient-level diversity, and SPARQ (havrilla2025synthetic) evaluates difficulty and diversity, linking them to in- and out-of-distribution performance.
 However, these methods remain constrained by the coverage of the seed pool and the inherent model biases.
-Doc-centric approaches instead derive questions directly from raw corpora, ensuring stronger factual grounding: UltraChat (Ding et al., 2023) generates world-knowledge questions from sources like Wikidata, Humpback (Li et al., 2024) infers instructions backward from documents, MAmmoTH2 (Yue et al., 2024) mines web content via a Recall-Extract-Refine pipeline.
-NaturalReasoning (Yuan et al., 2025) focuses on generating high-difficulty reasoning questions, but its diversity is inherently limited by a single prompt template.
-Furthermore, several studies have explored generating new questions through external guidance, such as KPDDS (Huang et al., 2025), and the structurally guided approaches of Xu et al.
-(2025), Wang et al.
-(2025), and Bu et al.
-(2025).
+Doc-centric approaches instead derive questions directly from raw corpora, ensuring stronger factual grounding: UltraChat (ding2023enhancing) generates world-knowledge questions from sources like Wikidata, Humpback (li2024self) infers instructions backward from documents, MAmmoTH2 (yue2024mammoth2) mines web content via a Recall-Extract-Refine pipeline.
+NaturalReasoning (yuan2025naturalreasoning) focuses on generating high-difficulty reasoning questions, but its diversity is inherently limited by a single prompt template.
+Furthermore, several studies have explored generating new questions through external guidance, such as KPDDS (huang2025key), and the structurally guided approaches of xu2025synthesis, wang2025rv, and bu2025enhanced.
 However, these approaches have two primary limitations compared to our method.
@@
 Reasoning Data Synthesis
-Another line of work leverages synthetic long CoT data to distill reasoning abilities into smaller models.
-DeepSeek (Guo et al., 2025) transfers reasoning skills from DeepSeek-R1 to open-source models, while OpenMathReasoning (Moshkov et al., 2025) augments math ability by generating long CoT traces for AoPS problems.
-OmniThought (Cai et al., 2025) collects multidisciplinary questions (math, code, science) to build a large CoT distillation dataset, and OpenThoughts (Guha et al., 2025) systematically explores recipes for generating long reasoning traces across these domains.
+Another line of work primarily investigates the synthesis of CoT reasoning data.
+DeepSeek (guo2025deepseek) transfers reasoning skills from DeepSeek-R1 to open-source models, while OpenMathReasoning (moshkov2025aimo) augments math ability by generating long CoT traces for AoPS problems.
+OmniThought (cai2025reasoning) collects multidisciplinary questions (math, code, science) to build a large CoT distillation dataset, and OpenThoughts (guha2025openthoughts) systematically explores recipes for generating long reasoning traces across these domains.
 While these approaches primarily focus on generating high-quality reasoning processes for existing questions, they lack the capability to produce original, diverse, and multidisciplinary questions, which is the central focus of our method.
@@
 Our core innovation is the concept of “Design Logic”, which abstracts the strategic process that human experts use to create challenging questions.
-By leveraging these logics, we generated two large-scale datasets, DLR-Book and DLR-Web, comprising millions of complex questions across 75 disciplines from diverse raw text sources.
+By leveraging design logics, we generated two large-scale datasets, DLR-Book and DLR-Web, comprising millions of complex questions across 75 disciplines from raw text sources.
 We validated the effectiveness of our approach through extensive experiments, demonstrating that models fine-tuned with our data achieve substantial performance gains in multidisciplinary reasoning.
+Importantly, the value of design logic lies not only in its effect on final performance but also in its controllable, abstractable, and generalizable synthesis mechanism.
+The guidance of the design logic is crucial for ensuring the difficulty and diversity of the synthesized questions.
+Section 4.1 shows that our design logics allow precise control over problem difficulty, enabling the generation of highly challenging questions far beyond baseline datasets and commonly used benchmarks.
+Section 4.2 shows that design-logic-guided synthesis achieves higher scores on all five diversity metrics compared with baseline datasets.
 Ethics Statement
@@
 Reproducibility Statement
-We provide all complete and detailed data processing steps used for dataset synthesis in the main paper and the appendix.
+We provide all complete and detailed data processing steps used for data synthesis in the main paper and in the appendix.
 To facilitate reproducibility and independent verification, we have released a subset of our synthesized data that does not involve legal compliance concerns or potential conflicts of interest.
-The released data includes our synthesized questions, reference answers, long CoT responses, the corresponding source corpora and design logic, as well as associated labels (e.g., discipline, difficulty, type), enabling reproduction and validation of our method’s effectiveness and the quality of the synthesized data.
+The released data include our synthesized questions, reference answers, long CoT responses, the corresponding source corpora and design logic, as well as associated labels (e.g., discipline, difficulty, question type), enabling reproduction and validation of our method’s effectiveness and the quality of the synthesized data.
+Furthermore, we have released a design-logic library that contains all design logics extracted in this study, together with discipline, difficulty, and question type labels, enabling the research community to directly match these design logics to their own raw documents and synthesize new questions that exhibit the same complex reasoning.
 The Use of Large Language Models (LLMs)
@@
 […]
-Number of Design Logics, Book Questions, and Web Questions by Discipline
+Number of Design Logics, book questions, and web questions by discipline.
 Discipline Design Logic Book Question Web Question
@@
 World History 987 39703 5503
+Total 125328 3040620 1658541
 Appendix D Baseline Datasets
@@
 OpenThoughts3
-(Guha et al., 2025)
 https://huggingface.co/datasets/open-thoughts/OpenThoughts3-1.2M
 Nemotron-Post-Training-v1
-(Bercovich et al., 2025)
 https://huggingface.co/datasets/nvidia/Nemotron-Post-Training-Dataset-v1
 WebInstruct (Full)
-(Yue et al., 2024)
 https://huggingface.co/datasets/TIGER-Lab/WebInstructFull
 NaturalReasoning
-(Yuan et al., 2025)
 https://huggingface.co/datasets/facebook/natural_reasoning
@@
 Benchmark Disciplines Rollout ($N$)
-MMLU (Hendrycks et al., 2020) 57 disciplines including STEM, humanities, social sciences, and professional fields 1
-MMLU-Pro (Wang et al., 2024b) Math, Physics, Chemistry, Law, Engineering, Economics, Health, Psychology, Business, Biology, Philosophy, Computer Science, History, Other 1
-GPQA-Diamond (Rein et al., 2024) Physics, Chemistry, Biology 10
-GPQA-Main (Rein et al., 2024) Physics, Chemistry, Biology 10
-SuperGPQA (Du et al., 2025) 285 graduate-level disciplines across 13 fields (e.g., Engineering, Management, Economics, Education, History, Science, Medicine, Law, Sociology, Philosophy, Agriculture, Literature) 1
-Appendix I Prompts
+MMLU (hendrycks2020measuring) 57 disciplines including STEM, humanities, social sciences, and professional fields 1
+MMLU-Pro (wang2024mmlu) Math, Physics, Chemistry, Law, Engineering, Economics, Health, Psychology, Business, Biology, Philosophy, Computer Science, History, Other 1
+GPQA-Diamond (rein2024gpqa) Physics, Chemistry, Biology 10
+GPQA-Main (rein2024gpqa) Physics, Chemistry, Biology 10
+SuperGPQA (du2025supergpqa) 285 graduate-level disciplines across 13 fields (e.g., Engineering, Management, Economics, Education, History, Science, Medicine, Law, Sociology, Philosophy, Agriculture, Literature) 1
+Appendix I Performance on GSM8K and MATH-500
+We evaluated the models trained on our data in Table 2 on GSM8K and MATH-500.
+Table 11: The SFT experiment results on GSM8K and MATH-500. The models are trained on DLR-Web, DLR-Book, or the combined DLR-Web + DLR-Book dataset.
+Model GSM8K MATH-500
+[…]
+[…]
+Llama-3.2-3B-SFT (DLR-Book) 81.05 $5±\pm1.4$2
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+GSM8k and MATH-500 are exclusively focused on mathematical problems.
+Our dataset’s objective is not mathematics; rather, it is designed to create a multidisciplinary dataset.
+Nevertheless, the model trained on our data still exhibits competitive performance in the mathematical domain.
+For MATH500, the slightly lower performance of our models compared with the official Qwen3 models that have undergone the full post-training process is likely due to the fact that the official Qwen3 models were trained with a substantially larger volume of high-quality, expert-written mathematical data and received more extensive post-training.
+In contrast, mathematics represents only about 10% of our dataset, and our math questions are synthesized from raw web and book corpora rather than curated by human experts, which naturally leads to lower quality relative to professional math-focused training data.
+Notably, models trained on either DLR-Book or DLR-Web+Book achieve 95.60% on GSM8K, matching the official Qwen3-8B score.
+GSM8K performance is already near saturation.
+Prior work (vendrow2025large) reports that approximately 5% of GSM8K questions are mislabeled or ambiguous, which makes 100% accuracy unattainable and positions 95.60% near the practical upper bound.
+Even stronger LLMs fall within a similar range.
+For instance, according to the Llama 3 technical report (dubey2024llama), GPT-4o reaches 96.1 on GSM8K under 8-shot CoT setting.
+Appendix J Prompts
 Prompt for Discipline Classification You are a professional multidisciplinary data labeling expert specializing in the classification of multidisciplinary academic questions.
@@
 The placeholders {logic1} through {logic5} and {text} are replaced with specific design logics and the source text, respectively.
-Appendix J Design Logic Examples
+Prompt for Question Answerability Evaluation Your task is to review a question and decide whether it provides sufficient conditions to derive solutions.
+If the question is a complete and answerable one, output ‘Yes’.
+Open-ended questions are considered answerable.
+If it is an incorrect question that cannot be answered, output ‘No’.
+Output only ‘Yes’ or ‘No’.
+‘‘question’’: ‘‘{question}’’
+Output:
+Figure 14: The prompt used to evaluate whether the synthesized questions are complete and answerable. The placeholder {question} is replaced with the specific synthesized question.
+Prompt for Design Logic Consistency Evaluation Your task is to judge whether a ‘‘question’’ is designed according to the Mermaid-format ‘‘design logic’’.
+You should ignore superficial wording changes.
+If the logic is consistent or if there are implicit references that preserve consistency, output ‘Yes’.
+If the design logic underlying the question is not consistent with the ‘‘design logic’’, output ‘No’.
+Output only ‘Yes’ or ‘No’.
+‘‘design logic’’: ‘‘{design_logic}’’
+‘‘question’’: ‘‘{question}’’
+Output:
+Figure 15: The prompt used to evaluate the consistency between the synthesized question and its corresponding design logic.
+The placeholders {design_logic} and {question} are replaced with the specific design logic and the corresponding synthesized question, respectively.
+Prompt for Answer Consistency Evaluation You are an expert reviewer who compares a ‘‘reference_answer’’ with a model ‘‘response’’ for the same question.
+Decide whether both provide the same final answer.
+Ignore differences in reasoning or wording, and output ‘Yes’ when the final answers or key conclusions are identical.
+Output ‘No’ when the conclusions differ.
+Output only ‘Yes’ or ‘No’.
+‘‘question’’: ‘‘{question}’’
+‘‘reference_answer’’: ‘‘{reference_answer}’’
+‘‘response’’: ‘‘{response}’’
+Output:
+Figure 16: The prompt used to evaluate the consistency between the reference answers generated by DeepSeek-R1-0528 and the CoT responses generated by Qwen3-235B-A22B-Thinking-2507-FP8.
+The placeholders {question}, {reference_answer}, and {response} are replaced with the synthesized question, the reference answer, and the CoT responses, respectively.
+Prompt for Discipline Label Evaluation You are reviewing exam questions and their author-provided discipline labels.
+Judge whether the discipline label is appropriate for the question.
+Answer ‘Yes’ when the label is plausible and answer ‘No’ when there is a clear mismatch.
+Output only ‘Yes’ or ‘No’.
+‘‘question’’: ‘‘{question}’’
+‘‘discipline label’’: ‘‘{discipline}’’
+Output:
+Figure 17: The prompt used to evaluate the appropriateness of the discipline label. The placeholders {question} and {discipline} are replaced with the specific question and its discipline label.
+Prompt for Difficulty Label Evaluation You are reviewing exam questions and their author-provided difficulty labels.
+Judge whether the difficulty label is appropriate for solving the question.
+Answer ‘Yes’ when the label is plausible and ‘No’ when there is a clear mismatch.
+Output only ‘Yes’ or ‘No’.
+‘‘question’’: ‘‘{question}’’
+‘‘difficulty label’’: ‘‘{difficulty}’’
+Output:
+Figure 18: The prompt used to evaluate the appropriateness of the difficulty label. The placeholders {question} and {difficulty} are replaced with the specific question and its difficulty label.
+Prompt for Type Label Evaluation You are reviewing exam questions and their author-provided type labels.
+Judge whether the type label is appropriate for the question.
+Answer ‘Yes’ when the label is plausible and answer ‘No’ when there is a clear mismatch.
+Output only ‘Yes’ or ‘No’.
+‘‘question’’: ‘‘{question}’’
+‘‘type label’’: ‘‘{qtype}’’
+Output:
+Figure 19: The prompt used to evaluate the appropriateness of the question type label. The placeholders {question} and {qtype} are replaced with the specific question and its type label.
+Appendix K Design Logic Examples
 In this section, we provide design logic examples for several representative disciplines, such as Computer Science and Technology, Clinical Medicine, Mathematics, Law, Psychology, and Archaeology.

### v6

@@
 Abstract
-Large language models (LLMs) have achieved remarkable success in many natural language tasks but still struggle with complex, multi-step reasoning, particularly across diverse disciplines.
+Large language models (LLMs) perform strongly on many language tasks but still struggle with complex multi-step reasoning across disciplines.
 Existing reasoning datasets often lack disciplinary breadth, reasoning depth, and diversity, as well as guiding principles for question synthesis.
-We propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages naturally available, extensive raw documents (e.g., book corpus and web corpus) to generate multidisciplinary challenging questions.
-We introduce the concept of “design logic” and instruct LLMs to mimic human educators’ question-creation process, enabling the automated synthesis of large-scale, high-difficulty questions.
+We propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages naturally available, extensive raw documents to generate multidisciplinary questions.
+The central insight is the notion of Design Logic, a form of reusable meta-knowledge that encapsulates the structured process human experts use to transform knowledge into complex exam questions, enabling LLMs to generate new questions with the same complex reasoning patterns from entirely different source texts with explicit control over difficulty, diversity, and question types.
 We use LLMs to reverse-engineer and abstract over 120,000 Design Logics from existing questions across various disciplines.
-By matching these design logics with source documents, we are able to generate reasoning questions with controllable question types and difficulty levels.
-Using this pipeline, we synthesized two large-scale reasoning datasets that span 75 disciplines: DLR-Book (3.04 million questions from the book corpus) and DLR-Web (1.66 million questions from the web corpus).
+By designing a two-stage retrieve-and-generate mechanism to match these Design Logics with raw corpus, we synthesized two large-scale reasoning datasets that span 75 disciplines: DLR-Book (3.04 million questions from the book corpus) and DLR-Web (1.66 million questions from the web corpus).
 Data analysis indicates that the questions synthesized by our method exhibit greater difficulty and diversity compared to those in the baseline datasets.
-We validate our synthesized data through supervised fine-tuning (SFT) on the Qwen3 and Llama3 model families.
-Our data substantially enhances their multidisciplinary reasoning capabilities, outperforming existing datasets.
-Notably, by applying SFT on the base versions of these models using only our data, we even surpass their official final models that have undergone the full post-training process.
+Supervised fine-tuning (SFT) on Qwen3 and Llama3 with our data substantially improves multidisciplinary reasoning and outperforms baseline datasets.
+Notably, by applying SFT on the base versions of these models using only our data, we even surpass their official final models that have undergone the full post-training.
 [^1] Project page: https://attention-is-all-i-need.github.io/Design-Logic-Reasoning
@@
 1 Introduction
-Large language models (LLMs) have demonstrated exceptional capabilities in various natural reasoning tasks (chowdhery2023palm; openai2023gpt4), such as mathematics and coding (trinh2024solving; zhu2024deepseek), especially when utilizing long chain-of-thought (CoT) techniques (wei2022chain; jaech2024openai; guo2025deepseek).
-However, they still lag behind human experts in university-level, discipline-specific reasoning (phan2025humanity), largely due to the scarcity of large-scale, high-quality, and diverse training data.
-Existing datasets focus mainly on math and programming, drawing on competition platforms rich in open-ended questions (moshkov2025aimo; cai2025reasoning), while most other disciplines lack comparable resources.
+Large language models (LLMs) have demonstrated exceptional capabilities in various natural reasoning tasks (Chowdhery et al., 2023; OpenAI, 2023), such as mathematics and coding (Trinh et al., 2024; Zhu et al., 2024), especially when utilizing long chain-of-thought (CoT) techniques (Wei et al., 2022; Jaech et al., 2024; Guo et al., 2025).
+However, they still lag behind human experts in university-level, discipline-specific reasoning (Phan et al., 2025), largely due to the scarcity of large-scale, high-quality, and diverse training data.
+Existing datasets focus mainly on math and programming, drawing on competition platforms rich in open-ended questions (Moshkov et al., 2025; Cai et al., 2025), while most other disciplines lack comparable resources.
 This scarcity limits the development of LLMs’ multidisciplinary reasoning capabilities.
-Data synthesis with LLMs is an effective solution to data scarcity (wang2024survey).
-Existing question synthesis methods fall into two categories: query-centric and doc-centric.
-Query-centric approaches expand seed questions through rewriting, added constraints (e.g., Evol-Instruct (xu2023wizardlm)), or incorporating chain-of-thought reasoning (wang2023self; yu2025cot), but they are limited by the coverage of the seed pool and inherent model biases.
-Doc-centric methods instead generate questions from unstructured (e.g., web, books) or structured (e.g., knowledge graphs) source documents (yue2024mammoth2; yuan2025naturalreasoning; huang2025key), ensuring broad disciplinary coverage grounded in authentic knowledge.
+Data synthesis with LLMs is an effective solution to data scarcity (Wang et al., 2024a).
+Existing question synthesis methods fall into two categories: query-centric and document-centric.
+Query-centric approaches expand seed questions through rewriting, added constraints (e.g., Evol-Instruct (Xu et al., 2023)), or incorporating chain-of-thought reasoning (Wang et al., 2023; Yu et al., 2025), but they are limited by the coverage of the seed pool and inherent model biases.
+Document-centric methods instead generate questions from unstructured (e.g., web, books) or structured (e.g., knowledge graphs) source documents (Yue et al., 2024; Yuan et al., 2025; Huang et al., 2025), ensuring broad disciplinary coverage grounded in authentic knowledge.
 However, they struggle to control difficulty and diversity, often degenerating into factual recall.
+Meanwhile, post-training (e.g., SFT) and even mid-training all rely heavily on exam-style data.
+This raises a key challenge: how can we rapidly synthesize large volumes of high-quality, human-like multidisciplinary exam questions while controlling their difficulty, diversity, and question types?
 Figure 1: Left: The procedure by which human experts construct questions reflects a “Design Logic”: a systematic sequence of deliberate steps that transforms fundamental knowledge points into complex, context-rich questions requiring multi-stage reasoning.
@@
 To address these issues, we propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages large-scale, multidisciplinary raw documents (e.g., book corpus and web corpus) to synthesize challenging questions across diverse disciplines.
-The central insight of our approach is the notion of “Design Logic”, which encapsulates how human experts transform knowledge into complex, reasoning-intensive exam questions.
+The central insight of our approach is the notion of Design Logic, which encapsulates how human experts transform knowledge into complex, reasoning-intensive exam questions.
 We observed that when human education experts design challenging and insightful questions, they do not merely state facts.
@@
 First, we process large-scale book and web corpora with multi-dimensional labeling and filtering (discipline, readability, educational value, reasoning depth) to construct a high-quality source material library.
-From a question bank of hundreds of millions, we cluster and sample a diverse set of difficult questions, from which an LLM reverse-engineers and abstracts over 120K structured “design logics” to construct a reusable design logic library.
+From a question bank of hundreds of millions, we cluster and sample a diverse set of difficult questions, from which an LLM reverse-engineers and abstracts over 120K structured Design Logics to construct a reusable Design Logic library.
 In question synthesis, we adopt a two-stage retrieve-and-generate mechanism: (1) vector similarity retrieves coarse candidate logics for each source document, and (2) an LLM performs a fine-grained evaluation to select the optimal logic and generates a reasoning question from the source document by strictly following its steps.
@@
 The main contributions of this paper can be summarized as follows:
-  • We propose a data engineering pipeline for synthesizing challenging questions from raw text corpora.
+  • we propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages raw corpora to synthesize multidisciplinary reasoning questions.
 Using this pipeline, we constructed two large-scale reasoning datasets: DLR-Book (3.04 million questions from the book corpus) and DLR-Web (1.66 million questions from the web corpus).
@@
   • By reverse-engineering the meta-knowledge of human educators, we propose a fundamentally new question synthesis method guided by “Design Logic”.
-This approach enables the generation of truly complex, multi-step reasoning questions from raw text by providing the structured, reusable, and abstract control over difficulty and diversity that prior doc-centric methods lacked.
+This approach enables the generation of truly complex, multi-step reasoning questions from raw text by providing the structured, reusable, and abstract control over difficulty and diversity that prior document-centric methods lacked.
 Our data analysis indicates that the questions synthesized by our method exhibit greater difficulty and diversity compared to those in the baseline datasets.
-  • We validate the effectiveness of our synthesized data through comprehensive comparative and ablation experiments on the Qwen3 (yang2025qwen3) and Llama3 (dubey2024llama) model families.
+  • We validate the effectiveness of our synthesized data through comprehensive comparative and ablation experiments on the Qwen3 (Yang et al., 2025a) and Llama3 (Dubey et al., 2024) model families.
 The results demonstrate that the data synthesized by our method significantly enhance the multidisciplinary reasoning capabilities of LLMs.
@@
 We annotate over 150 million questions from the proprietary question bank with discipline, difficulty, and question types using Qwen3-30B-A3B (non-thinking mode), following the prompts in Figure 7, Figure 8, and Figure 9, respectively.
-To obtain a high-quality and diverse subset, we compute embeddings with Qwen3-Embedding-4B and apply K-means clustering within each discipline (ahmed2020k), with cluster numbers determined by silhouette search.
+To obtain a high-quality and diverse subset, we compute embeddings with Qwen3-Embedding-4B and apply K-means clustering within each discipline (Ahmed et al., 2020), with cluster numbers determined by silhouette search.
 From each cluster, we draw an equal number of questions following a fixed difficulty ratio of 3:2:1 (Very Hard:Hard:Medium), and align per-discipline sizes with the overall bank distribution.
@@
 This corpus is processed at the chapter level, with chapters over 5,000 words split into smaller blocks and deduplicated via MinHash.
-Discipline labels are assigned by a ModernBERT-large classifier fine-tuned for disciplinary classification (warner2024smarter).
-Readability is predicted with a BERT-based model (turc2019well) to filter incoherent or disorganized text, while helpfulness (0–5) is scored by the fineweb-edu-classifier (lozhkov2024fineweb-edu) to quantify educational value.
-Based on quotas proportional to discipline frequencies in both the book corpus and the question bank, segments with negative readability are removed.
-The remaining candidates are ranked by helpfulness, and the top items are selected until the quotas are met.
-This procedure yields three million high-quality segments, with most scoring helpfulness $\geq 2$.
+Discipline labels are assigned by a ModernBERT-large classifier fine-tuned for disciplinary classification (Warner et al., 2024).
+Readability is predicted with a BERT-based model (Turc et al., 2019) to filter incoherent or disorganized text, while helpfulness (0–5) is scored by the fineweb-edu-classifier (Lozhkov et al., 2024) to quantify educational value.
+After removing segments with negative readability, remaining candidates are ranked by helpfulness and selected to meet discipline quotas proportional to frequencies in the book corpus and the question bank.
+This procedure yields three million high-quality segments, most with helpfulness $\geq 2$.
 2.3 Web Corpus Processing
@@
 For each question, the LLM also generates a concise reference answer.
-Question Deduplication and Decontamination
+Question Deduplication and Decontamination.
 We employ a two-stage filtering pipeline: (i) MinHash-based deduplication to remove near-duplicates, and (ii) 13-gram decontamination against all evaluation benchmarks to prevent leakage.
@@
 These question-response pairs are then used for supervised fine-tuning (SFT).
+However, response generation is not the primary focus of this work.
+Our objective is to synthesize high-quality multidisciplinary questions.
+As Albert Einstein noted, “The formulation of a problem is often more essential than its solution.” Given high-quality questions, responses can be generated by any model, including more powerful future models capable of achieving higher accuracy.
 4 Data Analysis
@@
 To quantify dataset diversity, we generate high-dimensional vector representations for 300,000 uniformly sampled questions from each dataset using the Qwen3-Embedding-4B model.
-We then compute the following five distance-based diversity metrics in the embedding space: Mean Cosine Distance (yang2025measuring), Mean L2 Distance (yang2025measuring), 1-Nearest Neighbor (1-NN) Distance (stasaski-hearst-2022-semantic), Cluster Inertia (du-black-2019-boosting), and Radius (lai-etal-2020-diversity).
+We then compute the following five distance-based diversity metrics in the embedding space: Mean Cosine Distance (Yang et al., 2025b), Mean L2 Distance (Yang et al., 2025b), 1-Nearest Neighbor (1-NN) Distance (Stasaski and Hearst, 2022), Cluster Inertia (Du and Black, 2019), and Radius (Lai et al., 2020).
 Detailed definitions and formulas for these metrics are provided in Appendix G.
@@
 The detailed number of questions per discipline in our dataset is provided in Table C.
-Figure 4: Discipline distribution across different datasets.
+Figure 4: Discipline distribution across different full datasets.
 4.4 Data Quality and Label Accuracy Validation
@@
 5.1 Experimental Setup
-SFT Settings
+SFT Settings.
 For all SFT experiments, we adhere to the hyperparameters detailed in Table 9.
-Evaluation Settings
+Evaluation Settings.
 To ensure a fair comparison, we employ a zero-shot evaluation setting for all trained models, using the consistent generation configuration specified in Table 9.
@@
 For $N>1$, we report mean accuracy (%) together with the standard deviation, and we also report the accuracy (%) under the Self-Consistency with Chain-of-Thought (CoT-SC) method, defined as the majority vote over the $N$ generated samples.
-Benchmarks
+Benchmarks.
 We evaluate the model’s multidisciplinary reasoning capability using the most widely adopted benchmarks.
@@
 We performed SFT on base models from the Qwen3 and Llama3 series using our synthetic datasets, DLR-Book and DLR-Web.
-The resulting models were subsequently compared against their official final models that have undergone the full Post-training process under our evaluation framework.
+The resulting models were subsequently compared against their official final models that have undergone the full post-training process under the same evaluation framework.
 Table 2: SFT experiment results. The models are trained on DLR-Web, DLR-Book, or the combined DLR-Web + DLR-Book dataset.
@@
 The observed scaling laws confirm that our method provides a reliable pathway to achieving superior model performance by synthesizing more data.
-Future work can leverage our proposed design logics to synthesize even larger datasets for continued improvement.
+Future work can leverage the Design Logics we provide to synthesize even larger datasets for continued improvement.
 Figure 5: Performance scaling with synthesized data size. Benchmark accuracy increases steadily with data scale.
@@
 5.5 Effect of Source Corpus Quality
-It is widely acknowledged that book corpora are of higher quality than web corpora (brown2020language; gao2020pile; rae2021scaling).
+It is widely acknowledged that book corpora are of higher quality than web corpora (Brown et al., 2020; Gao et al., 2020; Rae et al., 2021).
 To evaluate the impact of source quality on our data synthesis method, we conducted an SFT experiment on the Qwen3-8B-Base model, comparing data synthesized from the high-quality book corpus with that from the web corpus.
@@
 As shown in Figure 6, the model fine-tuned on data from the book corpus outperforms the one trained on data from the web corpus across most benchmarks, with the largest gains on complex reasoning tasks such as GPQA-Diamond and SuperGPQA.
-Nevertheless, the performance gap induced by the two source corpora of differing quality is negligible, which demonstrates that our method is robust to variations in source quality and can reliably generate high-quality questions even when applied to lower-quality corpora.
+Nevertheless, the performance gap induced by the two source corpora of differing quality is small, which demonstrates that our method is robust to variations in source quality and can reliably generate high-quality questions even when applied to lower-quality corpora.
 5.6 Ablation Studies
@@
 Where applicable, we used identical retrieval with Qwen3-Embedding-4B and generated responses using the same model, Qwen3-235B-A22B-Thinking-2507-FP8.
-Specifically, we evaluate the following four settings:
-  • DESIGNER: This is our complete method, which includes coarse retrieval of the top-5 design logics by similarity, followed by LLM-based fine selection and synthesis based on “design logic”.
-  • w/o Design Logic : This setting replaces our use of explicit design logics.
-Instead of being guided by a logic, the LLM is prompted to generate a new question by imitating the style and structure of the most suitable exemplar.
-This exemplar is selected from a set of 5 relevant example questions that are retrieved from the question bank.
+Specifically, we compared the following settings:
+  • DESIGNER: This is our complete method, which includes coarse retrieval of the top-5 Design Logics by similarity, followed by LLM-based fine selection and synthesis based on Design Logic.
+  • w/o Design Logic : This setting replaces our use of explicit Design Logics with exemplar questions.
+The retrieval process is identical to DESIGNER: semantic similarity-based coarse ranking yields top-5 candidate questions, then the LLM performs fine-grained selection to choose the most suitable exemplar and generates a new question by imitating its style and structure.
   • w/o Coarse Ranking: This setting bypasses the semantic similarity-based retrieval of the top-5 relevant Design Logics.
@@
 Table 4 summarizes the results of our ablation studies.
-Our complete “DESIGNER” approach consistently outperforms ablated configurations across most benchmarks.
-The better performance of using design logics relative to directly using questions as examples confirmed our hypothesis that explicit logical structures are more accurate and robust guides for high-quality question synthesis.
-The removal of either coarse-grained retrieval (w/o Coarse Ranking) or fine-grained LLM re-ranking (w/o Fine Ranking) causes a performance drop, confirming the value of each stage in our two-step matching process.
-We further evaluated the “w/o Design Logic” variant on the full book corpus (3.04 million samples).
-The performance degradation of “w/o Design Logic (Full)” is notably significant.
-Specifically, on the GPQA-Diamond benchmark, the performance gap between “w/o Design Logic (Full)” and the model employing “Design Logic” reaches five percentage points, providing stronger evidence for the advantage of employing design logic.
+Our complete DESIGNER approach outperforms all ablated configurations across benchmarks.
+The superior performance of using Design Logics compared with directly using questions as exemplars confirms our hypothesis that explicit logical structures provide more accurate and robust guidance for high-quality question synthesis.
+Removing either the coarse-grained retrieval stage (w/o Coarse Ranking) or the fine-grained LLM re-ranking stage (w/o Fine Ranking) leads to a performance drop, confirming the value of each stage in our two-step matching process.
+However, because these variants still benefit from the overall framework of our method, removing either stage only results in the use of a suboptimal Design Logic.
+Consequently, the expected impact is naturally limited.
+We further evaluated the “w/o Design Logic” variant on the full book corpus of 3.04 million samples, where the performance gap becomes more pronounced.
+In particular, the difference on GPQA-Diamond reaches five percentage points, providing stronger evidence for the advantage of employing Design Logic.
 6 Related Work
-Data Synthesis Paradigms
-Existing data synthesis methods can be broadly grouped into query-centric and doc-centric approaches.
-Query-centric approaches iteratively expand a seed query pool: Self-Instruct (wang2023self) samples from an initial pool to generate new QA pairs, while Wizard LM (luowizardcoder; xu2023wizardlm; luowizardmath) and Auto Evol-Instruct (zeng2024automatic) evolve instructions for greater diversity and complexity.
-CoT-Self-Instruct (yu2025cot) integrates reasoning into generation, Prismatic Synthesis (jung2025prismatic) emphasizes gradient-level diversity, and SPARQ (havrilla2025synthetic) evaluates difficulty and diversity, linking them to in- and out-of-distribution performance.
+Data Synthesis Paradigms.
+Existing data synthesis methods can be broadly grouped into query-centric and document-centric approaches.
+Query-centric approaches iteratively expand a seed query pool: Self-Instruct (Wang et al., 2023) samples from an initial pool to generate new QA pairs, while Wizard LM (Luo et al., 2024; Xu et al., 2023; Luo et al., 2025) and Auto Evol-Instruct (Zeng et al., 2024) evolve instructions for greater diversity and complexity.
+CoT-Self-Instruct (Yu et al., 2025) integrates reasoning into generation, Prismatic Synthesis (Jung et al., 2025) emphasizes gradient-level diversity, and SPARQ (Havrilla et al., 2025) evaluates difficulty and diversity, linking them to in- and out-of-distribution performance.
 However, these methods remain constrained by the coverage of the seed pool and the inherent model biases.
-Doc-centric approaches instead derive questions directly from raw corpora, ensuring stronger factual grounding: UltraChat (ding2023enhancing) generates world-knowledge questions from sources like Wikidata, Humpback (li2024self) infers instructions backward from documents, MAmmoTH2 (yue2024mammoth2) mines web content via a Recall-Extract-Refine pipeline.
-NaturalReasoning (yuan2025naturalreasoning) focuses on generating high-difficulty reasoning questions, but its diversity is inherently limited by a single prompt template.
-Furthermore, several studies have explored generating new questions through external guidance, such as KPDDS (huang2025key), and the structurally guided approaches of xu2025synthesis, wang2025rv, and bu2025enhanced.
+Document-centric approaches instead derive questions directly from raw corpora, ensuring stronger factual grounding: UltraChat (Ding et al., 2023) generates world-knowledge questions from sources like Wikidata, Humpback (Li et al., 2024) infers instructions backward from documents, MAmmoTH2 (Yue et al., 2024) mines web content via a Recall-Extract-Refine pipeline.
+NaturalReasoning (Yuan et al., 2025) focuses on generating high-difficulty reasoning questions, but its diversity is inherently limited by a single prompt template.
+Furthermore, several studies have explored generating new questions through external guidance, such as KPDDS (Huang et al., 2025), and the structurally guided approaches of Xu et al.
+(2025), Wang et al.
+(2025), and Bu et al.
+(2025).
 However, these approaches have two primary limitations compared to our method.
@@
 This approach is discipline-agnostic, making it applicable to any discipline.
-Reasoning Data Synthesis
+Reasoning Data Synthesis.
 Another line of work primarily investigates the synthesis of CoT reasoning data.
-DeepSeek (guo2025deepseek) transfers reasoning skills from DeepSeek-R1 to open-source models, while OpenMathReasoning (moshkov2025aimo) augments math ability by generating long CoT traces for AoPS problems.
-OmniThought (cai2025reasoning) collects multidisciplinary questions (math, code, science) to build a large CoT distillation dataset, and OpenThoughts (guha2025openthoughts) systematically explores recipes for generating long reasoning traces across these domains.
+DeepSeek (Guo et al., 2025) transfers reasoning skills from DeepSeek-R1 to open-source models, while OpenMathReasoning (Moshkov et al., 2025) augments math ability by generating long CoT traces for AoPS problems.
+OmniThought (Cai et al., 2025) collects multidisciplinary questions (math, code, science) to build a large CoT distillation dataset, and OpenThoughts (Guha et al., 2025) systematically explores recipes for generating long reasoning traces across these domains.
 While these approaches primarily focus on generating high-quality reasoning processes for existing questions, they lack the capability to produce original, diverse, and multidisciplinary questions, which is the central focus of our method.
@@
 Appendix D Baseline Datasets
-Table 5: The baseline datasets utilized in our experiments. For the Nemotron-Post-Training-v1 dataset, only the data from the math and STEM categories was used.
+Table 5: The baseline datasets utilized in our experiments. For the Nemotron-Post-Training-v1 dataset, we used the data from the math and STEM categories.
 Dataset URL
-OpenThoughts3
-https://huggingface.co/datasets/open-thoughts/OpenThoughts3-1.2M
-Nemotron-Post-Training-v1
-https://huggingface.co/datasets/nvidia/Nemotron-Post-Training-Dataset-v1
-WebInstruct (Full)
-https://huggingface.co/datasets/TIGER-Lab/WebInstructFull
-NaturalReasoning
-https://huggingface.co/datasets/facebook/natural_reasoning
+OpenThoughts3 (Guha et al., 2025) https://huggingface.co/datasets/open-thoughts/OpenThoughts3-1.2M
+Nemotron-Post-Training-v1 (Bercovich et al., 2025) https://huggingface.co/datasets/nvidia/Nemotron-Post-Training-Dataset-v1
+WebInstruct (Full) (Yue et al., 2024) https://huggingface.co/datasets/TIGER-Lab/WebInstructFull
+NaturalReasoning (Yuan et al., 2025) https://huggingface.co/datasets/facebook/natural_reasoning
 Appendix E Question Type Analysis
@@
 Benchmark Disciplines Rollout ($N$)
-MMLU (hendrycks2020measuring) 57 disciplines including STEM, humanities, social sciences, and professional fields 1
-MMLU-Pro (wang2024mmlu) Math, Physics, Chemistry, Law, Engineering, Economics, Health, Psychology, Business, Biology, Philosophy, Computer Science, History, Other 1
-GPQA-Diamond (rein2024gpqa) Physics, Chemistry, Biology 10
-GPQA-Main (rein2024gpqa) Physics, Chemistry, Biology 10
-SuperGPQA (du2025supergpqa) 285 graduate-level disciplines across 13 fields (e.g., Engineering, Management, Economics, Education, History, Science, Medicine, Law, Sociology, Philosophy, Agriculture, Literature) 1
+MMLU (Hendrycks et al., 2020) 57 disciplines including STEM, humanities, social sciences, and professional fields 1
+MMLU-Pro (Wang et al., 2024b) Math, Physics, Chemistry, Law, Engineering, Economics, Health, Psychology, Business, Biology, Philosophy, Computer Science, History, Other 1
+GPQA-Diamond (Rein et al., 2024) Physics, Chemistry, Biology 10
+GPQA-Main (Rein et al., 2024) Physics, Chemistry, Biology 10
+SuperGPQA (Du et al., 2025) 285 graduate-level disciplines across 13 fields (e.g., Engineering, Management, Economics, Education, History, Science, Medicine, Law, Sociology, Philosophy, Agriculture, Literature) 1
 Appendix I Performance on GSM8K and MATH-500
@@
 GSM8K performance is already near saturation.
-Prior work (vendrow2025large) reports that approximately 5% of GSM8K questions are mislabeled or ambiguous, which makes 100% accuracy unattainable and positions 95.60% near the practical upper bound.
+Prior work (Vendrow et al., 2025) reports that approximately 5% of GSM8K questions are mislabeled or ambiguous, which makes 100% accuracy unattainable and positions 95.60% near the practical upper bound.
 Even stronger LLMs fall within a similar range.
-For instance, according to the Llama 3 technical report (dubey2024llama), GPT-4o reaches 96.1 on GSM8K under 8-shot CoT setting.
+For instance, according to the Llama 3 technical report (Dubey et al., 2024), GPT-4o reaches 96.1 on GSM8K under 8-shot CoT setting.
 Appendix J Prompts